### v1

# Provably Efficient RLHF Pipeline: A Unified View from Contextual Bandits

Abstract
Reinforcement Learning from Human Feedback (RLHF) is a widely used approach for aligning Large Language Models (LLMs) with human preferences. While recent advancements have provided valuable insights into various stages and settings of RLHF, a comprehensive theoretical understanding of the entire RLHF pipeline remains lacking. Towards this end, we propose a unified framework for the RLHF pipeline from the view of contextual bandits and provide provable efficiency guarantees. In particular, we decompose the RLHF process into two distinct stages: (post-)training and deployment, exploring both passive and active data collection strategies during the training phase. By employing the Bradley-Terry preference model with a linearly parameterized reward function, we reformulate RLHF as a contextual preference bandit problem. We then develop novel algorithms for each stage, demonstrating significant improvements over existing approaches in both statistical and computational efficiency. Finally, we apply our method to train and deploy Llama-3-8B-Instruct on the Ultrafeedback-binarized dataset, and empirical results confirm the effectiveness of our approach.
1 Introduction
Table 1: Comparison between previous works and approach in terms of the statistical and computational efficiency across different stages of the RLHF pipeline. Our pipeline consists of two stages: training and deployment, where we explore both passive and active data collection strategies in the training stage. The statistical efficiency is measured using the sub-optimality gap (for the training stage) and regret (for the deployment stage), while the computational efficiency is evaluated based on time and storage complexity. Here, $d$ is the dimension of the feature space, $T$ is the number of samples, $\kappa$ is the non-linearity coefficient which can be exponentially large, $\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)$ is the concentratability vector, $V_{T}$ and $\mathcal{H}_{T}$ are two kinds of local norm that satisfies $\mathcal{H}_{T}\succeq(1/\kappa)V_{T}$.
Stage Data Reference Sub-optimality Gap/Regret Time Storage
Training Passive Zhu et al. (2023) $\widetilde{\mathcal{O}}\Big{(}\kappa\sqrt{d}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{V_{T}^{-1}}\Big{)}$ $\mathcal{O}(T\log T)$ $\mathcal{O}(T)$
Ours (Theorem 1) $\widetilde{\mathcal{O}}\Big{(}\sqrt{d}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}\Big{)}$ $\mathcal{O}(T)$ $\mathcal{O}(T)$
Active Das et al. (2024) $\widetilde{\mathcal{O}}\big{(}d\sqrt{\kappa/T}\big{)}$ $\mathcal{O}(T^{2}\log T)$ $\mathcal{O}(T)$
Ours (Theorem 2) $\widetilde{\mathcal{O}}\big{(}d\sqrt{\kappa/T}\big{)}$ $\mathcal{O}(T)$ $\mathcal{O}(1)$
Deployment Online Saha (2021) $\widetilde{\mathcal{O}}\big{(}d\kappa\sqrt{T}\big{)}$ $\mathcal{O}(T^{2}\log T)$ $\mathcal{O}(T)$
Ours (Theorem 3) $\widetilde{\mathcal{O}}\big{(}d\sqrt{\kappa T}\big{)}$ $\mathcal{O}(T)$ $\mathcal{O}(1)$
Reinforcement Learning from Human Feedback is a key technique for training large language models using human feedback (Ouyang et al., 2022; Bai et al., 2022). The RLHF process involves collecting data, each consisting of a prompt, a pair of responses, and a human preference label indicating the preferred response. Then, a reward model is trained to predict human preferences, and the LLM is fine-tuned based on the reward model by RL algorithms, such as PPO (Schulman et al., 2017).
Given the notable success of RLHF, recent efforts have been devoted to developing a deeper theoretical understanding of this approach. Zhu et al. (2023) investigated the standard offline setting, where the learner is provided with a fixed dataset and aims to learn a policy that maximizes the expected reward. In this setting, since the learner has no control over the data collection process, the quality of the dataset becomes crucial to the performance of the learned policy and the resulting policy often performs poorly when faced with out-of-distribution data (Burns et al., 2024). In practice, the Claude (Bai et al., 2022) and LLaMA-2 (Touvron et al., 2023) projects have demonstrated that iterative RLHF can significantly enhance model performance. Motivated by these practical advancements, Xiong et al. (2024) investigated the iterative setting, where the learner randomly selects prompts from the candidate set, generates response pairs by sampling from the current policy, and then queries for feedback to update the policy accordingly. They showed that RLHF benefits from this iterative data collection process. While uniform prompt sampling as a simple approach has been proven effective, Das et al. (2024) demonstrated that it can lead to inefficient sampling and the sub-optimality gap may fail to decrease as the number of samples increases. To this end, they studied the active setting where the learner is allowed to select both prompts and responses, query for human feedback, and update the policy accordingly.
Though these works have provided valuable insights into various stages and settings of RLHF, a comprehensive theoretical understanding of the entire RLHF pipeline remains lacking. Towards this end, in this work, we propose a unified framework for the RLHF pipeline from the view of contextual bandits. Specifically, building on recent theoretical advancements (Zhu et al., 2023; Xiong et al., 2024), we leverage the Bradley-Terry preference model with a linearly parameterized reward function to reformulate RLHF as a contextual bandit problem (Abe and Long, 1999). We then decompose the RLHF process into two distinct stages: (post-)training and deployment. In the training stage, the learner’s goal is to develop a high-performing model with minimal sample queries and computational resources, where we explore both passive and active data collection strategies. In the deployment stage, the trained model is used to select good responses to user inputs while concurrently gathering informative feedback from the user for further refinement.
Building on this unified view, we revisit the existing theoretical advancements in RLHF and observe that they primarily focus on specific stages and settings of the overall pipeline. Consequently, an understanding of the entire pipeline remains lacking. Furthermore, while they have proposed various algorithms with theoretical guarantees for different stages of RLHF, the high statistical and computational complexity of their methods presents a significant challenge for practical applications. Thus, a natural question arises:
Can we develop provably efficient algorithms with improved statistical and computational efficiency for the RLHF pipeline?
In this work, we answer this question affirmatively by proposing novel algorithms for each stage of the RLHF pipeline, improving both statistical and computational efficiency over existing methods. Specifically, our contributions are summarized as follows:
  • We propose a unified framework for the RLHF pipeline from the perspective of contextual bandits. We decompose the RLHF process into two stages: (post-)training and deployment, and explore both passive and active data collection strategies during training. This perspective allows us to view several prior works as focusing on specific components of the pipeline.
  • We propose algorithms that are computationally and statistically more efficient than existing methods for each stage of the RLHF pipeline. Specifically, we employ online mirror descent with a tailored local norm to replace the standard maximum likelihood estimation. Additionally, we design distinct query criteria for different stages. Table 1 summarizes the comparison between previous works and ours about statistical and computational efficiency.
  • We train and deploy Llama-3-8B-Instruct on the Ultrafeedback-binarized dataset using our method, testing both passive and active data collection strategies in the training stage. The empirical results validate our theoretical findings and confirm the effectiveness of our method.
Organization.
The remainder of this paper is organized as follows. We review related work in Section 2 and introduce the problem setup in Section 3. Our pipeline is presented in Sections 4 and 5, focusing on the training and deployment stages, respectively. We present experiments in Section 6 and conclude in Section 7. All proofs and experiment details are deferred to the appendix.
Notations.
For a vector $x\in\mathbb{R}^{d}$ and a positive semi-definite matrix $A\in\mathbb{R}^{d\times d}$, we define $\lVert x\rVert_{A}=\sqrt{x^{\top}Ax}$. We use the notation $\mathcal{O}(\cdot)$ to highlight various dependencies depending on the context. For the statistical efficiency, $\mathcal{O}(\cdot)$ omits only constant factors. For computational efficiency, $\mathcal{O}(\cdot)$ solely emphasizes the dependence on the number of samples as this is the primary factor influencing the complexity. Additionally, we employ $\widetilde{\mathcal{O}}(\cdot)$ to hide polylogarithmic factors.
2 Related Work
In this section, we review the works most closely related to ours.
Reinforcement Learning from Human Feedback.
Learning from human preferences with deep learning models dates back to Christiano et al. (2017) and has been recently popularized by the success of large language models (Bai et al., 2022; Touvron et al., 2023; Achiam et al., 2023). Existing results generally fall into two categories: reward-based and reward-free. Reward-based methods typically consist of two steps: reward modeling and reward maximization (Ouyang et al., 2022). Given preference data, the reward model is trained to predict preference labels, and the LLM is fine-tuned based on the reward model using RL algorithms such as PPO (Schulman et al., 2017). In contrast, reward-free methods directly optimize the LLM using human preference feedback, without relying on explicit reward modeling (Rafailov et al., 2023; Azar et al., 2024). Among these works, Direct Preference Optimization (DPO) (Rafailov et al., 2023) is one of the most popular reward-free methods, which treats generative models directly as reward models and optimizes them using human preference feedback. In this work, we focus on the reward-based RLHF methods.
Given the empirical success of RLHF, recent efforts have been devoted to developing a deeper theoretical understanding of this approach. Zhu et al. (2023) investigated the standard offline setting, where the learner is provided with a fixed dataset and aims to learn a policy that maximizes the expected reward. Xiong et al. (2024) investigated the iterative setting, where the learner randomly selects prompts from the candidate set, generates response pairs by sampling from the current policy, and then queries for feedback to update the policy accordingly. Das et al. (2024) studied the active setting where the learner is allowed to select both prompts and responses, query for feedback, and update the policy accordingly. Liu et al. (2024) focused on the dual active setting where the learner is allowed to select conversations and teachers simultaneously for feedback. Essentially, these works focus on specific stages and settings within the RLHF pipeline, while we offer a unified view on the entire process. Furthermore, we propose algorithms that are both computationally and statistically more efficient than existing methods for each stage of the RLHF pipeline.
Contextual Dueling Bandits and RL.
Dueling bandits is a variant of the multi-armed bandit problem that incorporates pairwise comparison feedback (Yue et al., 2012). Contextual dueling bandits further extend this framework by integrating context information (Dudík et al., 2015). In this setting, Saha (2021) studied the $K$-arm contextual dueling bandit problem and proposed an algorithm with a near-optimal regret guarantee. Saha et al. (2023) extended the contextual dueling bandit problem to the reinforcement learning setting, presenting an algorithm with a sublinear regret bound. Sekhari et al. (2024) further studied the contextual dueling bandit problem with active learning, where the learner is allowed to query for feedback adaptively and the goal is to minimize the regret while minimizing the number of queries simultaneously. However, these works primarily focus on the online setting, where the context is provided in an arbitrary and online manner and goal is to minimize cumulative regret. In this work, we consider the entire RLHF pipeline, including both training and deployment stages, and explore both passive and active data collection strategies in the training stage.
Active Learning.
Active learning is a well-established field that aims to reduce the labeling cost by selecting the most informative samples for annotation (Settles, 2009). The first line of work focuses on the pool-based setting (Seung et al., 1992; Freund et al., 1997; Huang et al., 2010). In this setting, the learner strategically selects a batch of the most informative data at each step and queries labels only for this selected batch. The learner then uses this labeled batch to update the model. Once the model is enhanced, the learner selects another batch of data and queries labels for it, continuing the iterative training process. The other line of work focuses on the online active setting (Cesa-Bianchi et al., 2004, 2006). In this setting, the learner sequentially observes data points and decides whether to query labels for them. Ji et al. (2024) explored the online active RLHF setting where prompts are provided in an arbitrary and online manner, with the goal of minimizing cumulative regret while minimizing the number of queries simultaneously. In this work, we leverage the pool-based active learning strategy in the training stage of the RLHF pipeline, where the learner is allowed to select the dataset actively.
3 Problem Setup
In this section, we introduce the problem setup for RLHF.
Following recent advancements in RLHF (Zhu et al., 2023; Xiong et al., 2024), we formulate RLHF as a contextual bandit problem. Specifically, we have a set of contexts $\mathcal{X}$ and a set of possible actions $\mathcal{A}$ per context. To learn with human preference feedback, the learner selects a tuple $(x,a,a^{\prime})$ to present to the human, where $x\in\mathcal{X}$ is the context, $a,a^{\prime}\in\mathcal{A}$ are the actions. The human then provides a binary preference feedback $y\in\{0,1\}$, where $y=1$ indicates the human prefers action $a$ over action $a^{\prime}$, and $y=0$ otherwise. We study the commonly used Bradley-Terry (BT) model in preference learning (Bradley and Terry, 1952), which assumes that the human’s preference is generated by a logistic function of the difference in the rewards of the two actions.
Definition 1 (Bradley-Terry Model).
Given a context $x\in\mathcal{X}$ and two actions $a,a^{\prime}\in\mathcal{A}$, the probability of the human preferring action $a$ over action $a^{\prime}$ is given by
$\mathbb{P}\left[y=1\mid x,a,a^{\prime}\right]=\frac{\exp\left(r(x,a)\right)}{\exp\left(r(x,a)\right)+\exp\left(r\left(x,a^{\prime}\right)\right)},$
where $r$ is a latent reward function.
We consider the linear reward model which is widely used in the RLHF literature (Zhu et al., 2023; Ji et al., 2024; Das et al., 2024).
Assumption 1.
It holds that $r(x,a)=\phi(x,a)^{\top}\theta^{*}$ where $\phi(x,a):\mathcal{X}\times\mathcal{A}\to\mathbb{R}^{d}$ is the known and fixed feature map, and $\theta^{*}\in\mathbb{R}^{d}$ is the unknown parameter vector. Furthermore, we assume $\lVert\phi(x,a)\rVert_{2}\leq L$ for all $x\in\mathcal{X}$ and $a\in\mathcal{A}$ and $\theta^{*}\in\Theta$ where $\Theta=\{\theta\in\mathbb{R}^{d}\mid\langle\bm{1},\theta\rangle=0,\|\theta\|_{2}\leq B\}$.
Remark 1.
The feature mapping $\phi$ can be constructed by removing the last layer of a pre-trained large language model, and $\theta^{*}$ corresponds to the weights of the last layer.
Based on the linear model, we can rewrite the probability as
$\mathbb{P}\left[y=1\mid x,a,a^{\prime}\right]=\sigma\left(\phi(x,a)^{\top}\theta^{*}-\phi(x,a^{\prime})^{\top}\theta^{*}\right),$ (1)
where $\sigma(w)=\frac{1}{1+\exp(-w)}$ is the sigmoid function. Next, we introduce a key quantity that captures learning complexity under the BT model.
Definition 2.
The non-linearity coefficient $\kappa$ is defined as
$\kappa=\max_{x\in\mathcal{X},a,a^{\prime}\in\mathcal{A}}\max_{\theta\in\Theta}\frac{1}{\dot{\sigma}\left(\phi(x,a)^{\top}\theta-\phi(x,a^{\prime})^{\top}\theta\right)},$ (2)
where $\dot{\sigma}(w)=\sigma(w)(1-\sigma(w))$ is the derivative function.
Remark 2.
Through direct calculation, we obtain $\kappa\leq 2+\exp(2BL)+\exp(-2BL)$. Thus, $\kappa$ may be exceedingly large with an exponential dependence on the magnitude of features and parameters, as justified in Appendix G.2.
The RLHF pipeline can generally be divided into two distinct stages: the (post-)training stage and the deployment stage, each with different objectives. In the training stage, the learner aims to develop a high-performing model with minimal samples and computational resources. In the deployment stage, the trained model is used to generate good responses to user inputs, while simultaneously gathering informative feedback for further refinement. We provide a detailed discussion of both stages in the following sections.
4 Training Stage
In the training stage, the learner aims to acquire a high-performing model with minimal number of samples and computational resources. To effectively evaluate the model’s performance, we define the following performance measure.
Performance Measure.
For any policy $\pi$, we define the sub-optimality gap compared to the optimal policy as follows:
$\operatorname*{\mathsf{SubOpt}}(\pi)=\mathbb{E}_{x\sim\rho}\left[r\left(x,\pi^{*}(x)\right)-r(x,\pi(x)\right],$ (3)
where $\rho$ is the context distribution, $\pi^{*}=\arg\max_{a}r(x,a)$ is the optimal policy under the true reward $r$. The goal of the learner is to minimize the performance gap $\operatorname*{\mathsf{SubOpt}}(\pi)$.
In the training stage, two data collection strategies are commonly used: passive and active. In passive setting, the learner is given a fixed set of samples. Conversely, in active setting, the learner can actively select queries for human feedback. We discuss both strategies below.
4.1 Passive Data Collection
In the passive data collection setting, we are given a fixed dataset $\left\{\left(x_{t},a_{t},a_{t}^{\prime},y_{t}\right)\right\}_{t=1}^{T}$ where $x_{t}\in\mathcal{X}$ is the context, $a_{t},a_{t}^{\prime}\in\mathcal{A}$ are the actions, and $y_{t}\in\{0,1\}$ is the feedback. Since the parameter $\theta^{*}$ is unknown, we need to estimate it from the available data. The most common method for this estimation is maximum likelihood estimation (MLE). First, let $z=\phi(x,a)-\phi(x,a^{\prime})$ be the feature difference, then we define the negative log-likelihood $\ell_{t}(\theta)$ for sample $t$ as
$\ell_{t}(\theta)=-y_{t}\log(\sigma(z_{t}^{\top}\theta))-\left(1-y_{t}\right)\log(1-\sigma(z_{t}^{\top}\theta)).$ (4)
The MLE estimator $\widehat{\theta}_{T+1}$ is then defined as
$\widehat{\theta}_{T+1}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\mathcal{L}_{T+1}(\theta)\triangleq\sum_{t=1}^{T}\ell_{t}(\theta).$ (5)
Several works (Zhu et al., 2023; Ji et al., 2024) have analyzed the statistical properties of MLE and showed that for any $\delta\in(0,1)$, with probability at least $1-\delta$, the MLE estimator $\widehat{\theta}_{T+1}$ converges to the true parameter $\theta^{*}$ by
$\lVert\widehat{\theta}_{T+1}-\theta^{*}\rVert_{V_{T+1}}\leq\widehat{\beta}\triangleq\widetilde{\mathcal{O}}\big{(}\kappa\sqrt{d}\big{)},$ (6)
where $V_{t}\triangleq\sum_{i=1}^{t-1}z_{i}z_{i}^{\top}+\lambda I$ is the feature covariance matrix, $\lambda$ is the regularization parameter.
For computational complexity, the MLE estimator does not have a closed-form solution, requiring iterative optimization techniques, such as gradient descent, to obtain an $\varepsilon$-accurate solution. As discussed in Faury et al. (2022), optimizing the MLE typically requires $\mathcal{O}(\log(1/\varepsilon))$ iterations to reach the desired accuracy. Since the loss function is defined over all data samples, each iteration involves $\mathcal{O}(T)$ gradient computations. Given that $\varepsilon$ is typically set to $1/T$ to prevent it from dominating the overall estimation error, the total computational complexity is $\mathcal{O}(T\log T)$.
Although the MLE estimator is a commonly used approach, we propose a novel algorithm that enhances both computational efficiency and sample efficiency compared to MLE, drawing inspiration from recent advancements in logistic bandits (Faury et al., 2022; Zhang and Sugiyama, 2023) and multinomial logit MDPs (Li et al., 2024). First, we define the gradient $g_{t}(\theta)$ and Hessian $H_{t}(\theta)$ of the loss function $\ell_{t}(\theta)$ as
$g_{t}(\theta)=\left(\sigma\left(z_{t}^{\top}\theta\right)-y_{t}\right)z_{t},~{}~{}H_{t}(\theta)=\dot{\sigma}\left(z_{t}^{\top}\theta\right)z_{t}z_{t}^{\top}.$ (7)
Implicit OMD.
To improve the computational efficiency, Faury et al. (2022) observed that the cumulative past log-loss is strongly convex and can therefore be well approximated by a quadratic function. Building on this observation, they proposed the following update rule:
$\bar{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\ell_{t}(\theta)+\frac{1}{2\eta}\left\|\theta-\bar{\theta}_{t}\right\|_{\bar{\mathcal{H}}_{t}}^{2}\Big{\}},$ (8)
where ¯$\bar{\mathcal{H}}_{t}=\sum_{i=1}^{t-1}H_{i}\left(\bar{\theta}_{i+1}\right)+\lambda I$ is the local norm, and $\eta$ is the step size. The optimization problem can be decomposed in two terms. The first term is the instantaneous log-loss $\ell_{t}(\theta)$, which accounts for the information of the current sample. The second consists of a quadratic proxy for the past losses constructed through the sequence $\{\bar{\theta}_{i}\}_{i\leq t}$. A key component is the design of the local norm ¯$\bar{\mathcal{H}}_{t}$, which approximates the Hessian matrix by $H_{i}\left(\bar{\theta}_{i+1}\right)$ at a lookahead point $\bar{\theta}_{i+1}$. Such a Hessian matrix effectively captures local information and is crucial for ensuring statistical efficiency.
Although this formulation may appear computationally more efficient than MLE due to its one-pass data processing, the optimization problem in Eq. (8) still lacks a closed-form solution. Consequently, it still requires $\mathcal{O}(\log(1/\varepsilon))$ iterations per sample to achieve an $\varepsilon$-accurate solution, leading to the same $\mathcal{O}(T\log T)$ computational complexity as MLE.
Standard OMD.
To enhance computational efficiency, we revisit the formulation in Eq. (8) and observe that it essentially corresponds to the implicit OMD framework (Kivinen and Warmuth, 1997). Consequently, a natural alternative is to replace this formulation with the standard OMD framework, which permits a closed-form solution and thus eliminates the need for iterative optimization. However, the standard OMD minimizes a first-order approximation of the loss function, which sacrifices several key properties compared to its implicit counterpart, as demonstrated by Campolongo and Orabona (2020). Specifically, the standard OMD formulation updates using $\partial\ell_{t}(\theta_{t})$, whereas the implicit OMD updates the algorithm approximately with the subsequent sub-gradient, $\partial\ell_{t}(\theta_{t+1})$. This distinction results in a notable gap in the convergence rates of the two methods. To this end, we propose to approximate the current loss $\ell_{t}(\theta)$ using a second-order Taylor expansion, drawing inspiration from Zhang and Sugiyama (2023). We define the second-order approximation of the original loss function $\ell_{t}(\theta)$ at the current estimator $\widetilde{\theta}_{t}$ as ~$\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+g_{t}(\widetilde{\theta}_{t})^{\top}(\theta-\widetilde{\theta}_{t})+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}$. Then, we replace the current loss $\ell_{t}(\theta)$ in Eq. (8) with the approximation ~$\widetilde{\ell}_{t}(\theta)$, leading to the following update rule:
$\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\big{\langle}g_{t}(\widetilde{\theta}_{t}),\theta\big{\rangle}+\frac{1}{2\eta}\big{\|}\theta-\widetilde{\theta}_{t}\big{\|}_{\widetilde{\mathcal{H}}_{t}}^{2}\Big{\}},$ (9)
where $\eta$ is the step size and ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$ is the local norm with $\mathcal{H}_{t}\triangleq\mathcal{H}_{t}(\widetilde{\theta}_{i+1})=\sum_{i=1}^{t-1}H_{i}(\widetilde{\theta}_{i+1})+\lambda I$.
The optimization problem in Eq. (9) can be solved with a projected gradient step with the following equivalent formulation:
$\widetilde{\theta}_{t+1}^{\prime}=\widetilde{\theta}_{t}-\eta\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t}),\quad\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\widetilde{\theta}_{t+1}^{\prime}\rVert_{\widetilde{\mathcal{H}}_{t}}^{2}.$
Thus, the estimator $\widetilde{\theta}_{t+1}$ can be computed in an $\mathcal{O}(1)$ time complexity, leading to a total time complexity of $\mathcal{O}(T)$ for the entire procedure. This represents a significant improvement over the $\mathcal{O}(T\log T)$ complexity of the MLE estimator in Eq. (5), particularly considering that the number of samples $T$ is typically large in the training of LLMs. Furthermore, since the estimator processes the samples in a one-pass manner, it mitigates the memory burden associated with computing the gradient of the full dataset at each iteration. This makes it more suitable for large-scale applications, where memory and time efficiency are crucial.
It is important to note that the update rule in Eq. (9) represents a special case of online mirror descent, specifically, it can be derived from the following general form:
$\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\big{\langle}g_{t}(\widetilde{\theta}_{t}),\theta\big{\rangle}+\frac{1}{\eta}\mathcal{D}_{\psi_{t}}(\theta,\widetilde{\theta}_{t})\Big{\}},$
where $\psi_{t}(\theta)=\frac{1}{2}\lVert\theta\rVert_{\widetilde{\mathcal{H}}_{t}}^{2}$ is the regularizer and $\mathcal{D}_{\psi_{t}}(\theta,\widetilde{\theta}_{t})=\psi_{t}(\theta)-\psi_{t}(\widetilde{\theta}_{t})-\langle\nabla\psi_{t}(\widetilde{\theta}_{t}),\theta-\widetilde{\theta}_{t}\rangle$ is the Bregman divergence. Therefore, we can analyze the estimation error using the online mirror descent framework (Orabona, 2019). Specifically, we have the following lemma.
Algorithm 1 Training Stage: Passive Data Collection
0:  Dataset $\left\{\left(x_{t},a_{t},a_{t}^{\prime},y_{t}\right)\right\}_{t=1}^{T}$, regularization $\lambda>0$
1:  Initialize $\widetilde{\theta}_{1}=0$, $\mathcal{H}_{1}=\lambda I_{d}$.
2:  for $t=1,2,\ldots,T$ do
3:     Define $g_{t}(\theta)$ and $H_{t}(\theta)$ as in Eq. (7)
4:     Update ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$
5:     Compute $\widetilde{\theta}_{t+1}$ by Eq. (9)
6:     Update $\mathcal{H}_{t+1}=\mathcal{H}_{t}+H_{t}(\widetilde{\theta}_{t+1})$
7:  end for
8:  Compute $\widetilde{J}_{T}(\pi)$ by Eq. (11)
8:  Compute final policy $\pi_{T}=\operatorname*{arg\,max}_{\pi}\widetilde{J}_{T}(\pi)$
Lemma 1.
Let $\delta\in(0,1]$, set $\eta=(1/2)\log 2+\left(BL+1\right)$ and $\lambda=84\sqrt{2}\eta(dL^{2}+BL^{3})$, define the confidence set as
$\mathcal{C}_{t}=\left\{\theta\in\Theta\mid\big{\|}\theta-\widetilde{\theta}_{t}\big{\|}_{\mathcal{H}_{t}}\leq\widetilde{\beta}_{t}\triangleq\mathcal{O}\big{(}\sqrt{d}(\log(t/\delta))^{2}\big{)}\right\}.$
Then, we have $\operatorname{Pr}\left[\forall t\geqslant 1,\theta^{*}\in\mathcal{C}_{t}(\delta)\right]\geqslant 1-\delta$.
Intuitively, the confidence set shrinks as the number of samples $t$ increases, owing to the growing local norm $\mathcal{H}_{t}$. The local norm $\mathcal{H}_{t}$ captures the directions explored by the samples. The more we explore in a particular direction, the more accurate our estimation becomes along that direction.
By the definition of $\kappa$ in Definition 2, it holds that $\mathcal{H}_{t}\succeq(1/\kappa)V_{t}$, Lemma 1 implies $\forall t\in[T+1]$,
$\lVert\widetilde{\theta}_{t}-\theta^{*}\rVert_{V_{t}}\leq\sqrt{\kappa}\lVert\widetilde{\theta}_{t}-\theta^{*}\rVert_{\mathcal{H}_{t}}\leq\widetilde{\mathcal{O}}\big{(}\sqrt{\kappa d}\big{)}.$ (10)
Compared with the confidence set in Eq. (6), our confidence set strictly improves the estimation error by a factor of $\sqrt{\kappa}$. As highlighted in Remark 2, the non-linearity coefficient $\kappa$ may be exceedingly large during the training of LLMs, which makes our estimator statistically more efficient.
Pessimistic Policy.
Based on the estimator in Eq. (9), we adopt the “pessimism in the face of uncertainty” principle and define the pessimistic value function as
$\widetilde{J}_{T}(\pi)=\min_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[(\phi(x,\pi(x))-v)^{\top}\theta\right]=(\bar{\phi}(\pi)-v)^{\top}\widetilde{\theta}_{T}-\widetilde{\beta}_{T}\lVert\bar{\phi}(\pi)-v\rVert_{\mathcal{H}_{T}^{-1}},$ (11)
where $\bar{\phi}(\pi)=\mathbb{E}_{x\sim\rho}[\phi(x,\pi(x))]$ is the feature expectation of policy $\pi$, and $v$ is the baseline feature vector. While $v$ does not affect the induced policy, it may help reduce the sub-optimality gap (Zhu et al., 2023). The final policy is given by $\pi=\operatorname*{arg\,min}_{\pi}\widetilde{J}(\pi)$. The procedure is present in Algorithm 1. We show it enjoys the following guarantee.
Theorem 1.
Let $\delta\in(0,1]$, set the parameters as in Lemma 1, with probability at least $1-\delta$, Algorithm 1 ensures the following sub-optimality gap:
${\operatorname*{\mathsf{SubOpt}}}(\pi_{T})\leq\widetilde{\mathcal{O}}\left(\sqrt{d}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}\right).$
Remark 3.
The term $\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}$ is referred to as the “concentratability coefficient”, a widely used concept in offline learning (Jin et al., 2021; Zhu et al., 2023). It measures the distribution shift between the optimal policy and the dataset distribution. A smaller concentratability coefficient value indicates that the dataset better covers the data distribution under the optimal policy, resulting in a smaller sub-optimality gap.
Comparison with Zhu et al. (2023).
Our result improves upon their work in both statistical and computational efficiency. First, similar to Eq. (10), Theorem 1 implies an $\widetilde{\mathcal{O}}\big{(}\sqrt{\kappa d}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{V_{T}^{-1}}\big{)}$ sub-optimality gap, improving upon theirs by a $\sqrt{\kappa}$ factor. Moreover, our estimator has an $\mathcal{O}(T)$ time complexity, more efficient than their $\mathcal{O}(T\log T)$ complexity.
4.2 Active Data Collection
In this part, we consider the active data collection setting, where the learner can actively select tuple $\{x_{t},a_{t},a^{\prime}_{t}\}_{t=1}^{T}$ to present to the human for feedback $\{y_{t}\}_{t=1}^{T}$.
Estimator Construction.
For each round $t\in[T]$, given $\left\{\left(x_{i},a_{i},a_{i}^{\prime},y_{i}\right)\right\}_{i=1}^{t-1}$, we define the loss function $\ell_{t}(\theta)$ as in Eq. (4), then the most standard approach is to estimate the parameter $\theta^{*}$ by MLE, which is given by $\widehat{\theta}_{t}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum_{i=1}^{t-1}\ell_{i}(\theta)$. However, as noted before, the MLE estimator leads to a time complexity of $\mathcal{O}(t\log t)$ per round. Moreover, unlike the passive setting where the estimator is constructed only once, in the active setting, we need to construct the estimator in each round. This is prohibitively expensive for large-scale applications.
To improve the computational efficiency, we can leverage the OMD framework as in the passive data collection setting. Specifically, we construct the estimator $\widetilde{\theta}_{t}$ as in Eq. (9) by
$\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\big{\langle}g_{t}(\widetilde{\theta}_{t}),\theta\big{\rangle}+\frac{1}{2\eta}\big{\|}\theta-\widetilde{\theta}_{t}\big{\|}_{\widetilde{\mathcal{H}}_{t}}^{2}\Big{\}}.$ (12)
Algorithm 2 Training Stage: Active Data Collection
0:  Regularization $\lambda>0$, failure probability $\delta\in(0,1]$
1:  Initialize $\widetilde{\theta}_{1}=0$, $\mathcal{H}_{1}=\lambda I_{d}$.
2:  for $t=1,2,\ldots,T$ do
3:     Choose $(x_{t},a_{t},a_{t}^{\prime})$ by Eq. (13), observe feedback $y_{t}$
4:     Define $g_{t}(\theta)$ and $H_{t}(\theta)$ as in Eq. (7)
5:     Update ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$
6:     Compute $\widetilde{\theta}_{t+1}$ by Eq. (12)
7:     Update $\mathcal{H}_{t+1}=\mathcal{H}_{t}+H_{t}(\widetilde{\theta}_{t+1})$
8:  end for
9:  Construct $\widetilde{r}_{T}(x,a)=\frac{1}{T}\sum_{t=1}^{T}\phi(x,a)^{\top}\widetilde{\theta}_{t},\forall x,a$
9:  Compute policy $\pi_{T}(x)=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{r}_{T}(x,a)$.
Query Selection.
Given the estimator $\widetilde{\theta}_{t}$, we choose the query with the maximum uncertainty inspired by (Das et al., 2024). Specifically, we choose the tuple by
$(x_{t},a_{t},a_{t}^{\prime})=\operatorname*{arg\,max}_{x,a,a^{\prime}\in\mathcal{X}\times\mathcal{A}\times\mathcal{A}}\left\{\big{\lVert}\phi(x,a)-\phi(x,a^{\prime})\big{\rVert}_{\mathcal{H}_{t}^{-1}}\right\}.$ (13)
Intuitively, $\mathcal{H}_{t}$ captures the directions explored so far, the selection criterion essentially seeks the tuple whose feature difference is poorly covered by the past observations.
Policy Output.
After $T$ rounds, we define the reward function as $\widetilde{r}_{T}(x,a)=\frac{1}{T}\sum_{t=1}^{T}\phi(x,a)^{\top}\widetilde{\theta}_{t}$ as the average of all the past estimations. Then, the final policy is given by
$\pi_{T}(x)=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{r}_{T}(x,a).$ (14)
The detailed procedure is present in Algorithm 2. We show it enjoys the following guarantee.
Theorem 2.
Let $\delta\in(0,1]$, set the parameters as in Lemma 1, with probability at least $1-\delta$, Algorithm 2 ensures the following sub-optimality gap:
$\operatorname*{\mathsf{SubOpt}}(\pi_{T})\leq\widetilde{\mathcal{O}}\big{(}d\sqrt{{\kappa}/{T}}\big{)}.$
Remark 4.
In contrast to the passive data collection setting, the sub-optimality gap in Theorem 2 does not depend on the concentratability coefficient term. Instead, it decays at a rate of $T^{-1/2}$. This improvement arises because the active setting enables our algorithm to actively select queries with the maximum uncertainty, ensuring the uncertainty across all directions is effectively controlled.
Remark 5.
It is important to note that our result has a dependence on $\sqrt{\kappa}$. Though recent advances in the logistic bandits (Abeille et al., 2021; Zhang and Sugiyama, 2023) have shown that this dependence can be eliminated in the dominant term and shifted to lower-order terms, we believe this term remains necessary in our setting. This is due to the fact that the sub-optimality gap in our setting is defined as the difference between the reward $z^{\top}\theta^{*}$ while they focus the probability difference, i.e., $\sigma(z^{\top}\theta^{*})$. This difference prevents us from removing the $\sqrt{\kappa}$ term in our setting.
Comparison with Das et al. (2024).
Our algorithm attains the same sub-optimality gap as the work of Saha (2021), but with a significantly improved computational efficiency. Specifically, our algorithm has a time complexity of $\mathcal{O}(T)$ for the total $T$ round, which is more efficient than the $\mathcal{O}(T^{2}\log T)$ complexity of their MLE estimator. Moreover, our algorithm does not require storing the historical samples, resulting in only an $\mathcal{O}(1)$ storage complexity. In contrast, their MLE estimator necessitates storing all historical samples, leading to an $\mathcal{O}(T)$ storage complexity.
5 Deployment Stage
In the deployment stage, the learner uses the trained model to generate responses to user inputs and simultaneously gathering informative feedback to improve the model. This contrasts with the active data collection phase during training, which is focused solely on optimizing the final policy. In deployment, however, the learner is faced with a dual objective: selecting actions that maximize rewards to ensure a positive user experience, while also choosing actions that provide informative feedback for further model improvement.
Performance Measure.
To this end, we consider the cumulative regret compared to the optimal policy, which is defined as
${\operatorname*{Reg}}_{T}=\sum_{t=1}^{T}\left(r\left(x_{t},\pi^{*}(x_{t})\right)-\frac{r\left(x_{t},a_{t}\right)+r\left(x_{t},a^{\prime}_{t}\right)}{2}\right),$
where $\pi^{*}$ is the optimal policy, $a_{t}$ and $a^{\prime}_{t}$ are the actions selected by the learner at round $t$. This measure is stronger than the measure defined in Ji et al. (2024), which considers the gap between the optimal action and a single selected action. In the deployment phase, our proposed measure is more appropriate, as both actions must be sufficiently effective to ensure a positive user experience.
While our measure is more suitable for deployment, it also introduces significant challenges. To optimize the measure of Ji et al. (2024), the learner can exploit one action with high reward for exploitation while using the other action for exploration. In contrast, our setting requires both actions to be good enough. To address this, we propose a novel algorithm that balances exploration and exploitation for deployment. At a high level, our algorithm consists of two key components: estimator construction and query selection. We describe each component in detail below.
Estimator Construction.
The estimator construction is the same as the process used in the training stage. For each round $t\in[T]$, given $\left\{\left(x_{i},a_{i},a_{i}^{\prime},y_{i}\right)\right\}_{i=1}^{t-1}$, we construct the estimator $\widetilde{\theta}_{t}$ by OMD in an one-pass manner as in Eq. (9). Specifically, the estimator is updated by
$\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\big{\langle}g_{t}(\widetilde{\theta}_{t}),\theta\big{\rangle}+\frac{1}{2\eta}\big{\|}\theta-\widetilde{\theta}_{t}\big{\|}_{\widetilde{\mathcal{H}}_{t}}^{2}\Big{\}}.$ (15)
Query Selection.
Unlike in the training stage, during deployment, the learner must balance exploration and exploitation to minimize regret. Specifically, the learner must select queries that are both informative and likely to yield high rewards. Focusing solely on high-reward queries may cause the learner to miss valuable feedback for model improvement, while prioritizing only informative queries could prevent reward maximization. To address this, we propose a novel strategy inspired by Saha (2021), which accounts for both reward maximization and uncertainty maximization.
Based on the estimator $\theta_{t}$, we first define the promising set $\mathcal{A}_{t}$ that only contains the actions with potential to be the best action. Specifically, we define the promising set as
$\mathcal{A}_{t}=\big{\{}a\in\mathcal{A}\mid(\phi(x_{t},a)-\phi(x_{t},a^{\prime}))^{\top}\widetilde{\theta}_{t}+\widetilde{\beta}_{t}\lVert\phi(x_{t},a)-\phi(x_{t},a^{\prime})\rVert_{\mathcal{H}_{t}^{-1}}\geq 0,\forall a^{\prime}\in\mathcal{A}\big{\}}.$ (16)
Then, we select the action pairs with the maximum uncertainty in the promising set, i.e.,
$(a_{t},a_{t}^{\prime})=\operatorname*{arg\,max}_{(a,a^{\prime})\in\mathcal{A}_{t}\times\mathcal{A}_{t}}\left\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\right\|_{\mathcal{H}_{t}^{-1}}.$ (17)
The overall algorithm is summarized in Algorithm 3. We show it enjoys the following regret bound.
Algorithm 3 Deployment Stage
0:  Regularization $\lambda>0$, failure probability $\delta\in(0,1]$
1:  Initialize $\widetilde{\theta}_{1}=0$, $\mathcal{H}_{1}=\lambda I_{d}$.
2:  for $t=1,2,\ldots,T$ do
3:     Observes context $x_{t}$.
4:     Compute the promising set by Eq. (16)
5:     Choose duel actions by Eq. (17).
6:     Observe preference feedback $y_{t}$
7:     Update ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$
8:     Compute $\widetilde{\theta}_{t+1}$ by Eq. (15)
9:     Update $\mathcal{H}_{t+1}=\mathcal{H}_{t}+H_{t}(\widetilde{\theta}_{t+1})$
10:  end for
Theorem 3.
Let $\delta\in(0,1]$, set parameters as in Lemma 1, with probability at least $1-\delta$, Algorithm 3 enjoys the following regret guarantee:
${\operatorname*{Reg}}_{T}\leq\widetilde{\mathcal{O}}\big{(}d\sqrt{{\kappa}{T}}\big{)}.$
Comparison with Saha (2021).
Our result improves upon theirs in both computational and statistical efficiency. For statistical efficiency, our algorithm achieves a regret of $\widetilde{\mathcal{O}}\big{(}d\sqrt{{\kappa}{T}}\big{)}$, which is tighter by a factor of $\sqrt{\kappa}$ compared to their $\widetilde{\mathcal{O}}\big{(}d{\kappa}\sqrt{{T}}\big{)}$ result. For computational efficiency, our algorithm updates in an online manner, with $\mathcal{O}(1)$ storage and $\mathcal{O}(T)$ time complexity for the total $T$ rounds. In contrast, their algorithm requires storing all the data and solving an optimization problem each round, leading to $\mathcal{O}(T)$ storage and $\mathcal{O}(T^{2}\log T)$ time complexity. As $T$ is large or even infinite during deployment, our algorithm has a significant advantage in computational efficiency.
6 Experiment
In this section, we empirically evaluate the performance of our proposed method in terms of both effectiveness and efficiency. Specifically, we empirically validate our method with LLMs, and construct an online RLHF pipeline to evaluate different methods in both the training and deployment stage. [^1] Equal contribution. Next, we first describe the experimental setup, and then present the empirical results.
Experimental Setup.
In our experiments, we employ the Llama-3-8B-Instruct  [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct as the foundation model for reward model. We extract features $\phi(x,a)$ using last layer of the model, and the feature space dimension is $d=4096$. We use the Ultrafeedback-binarized dataset, a pro-processed version of the original Ultrafeedback dataset (Cui et al., 2023) for evaluation, which is a large-scale, fine-grained, diverse preference dataset, widely used as a benchmark for RLHF. It collects about 64, 000 prompts from diverse resources, including question answering, summarization, and dialogue generation. Each data consists of a context $x$, two responses $a$ and $a^{\prime}$, and a human preference label $y$.
11footnotetext: Our online RLHF pipeline code is available at https://github.com/ZinYY/RLHF_Pipeline
Implementation Details.
All experiments in this section were performed using four Nvidia A800 (80 GB) GPUs with two Intel(R) Xeon(R) Gold 6430 CPUs. The learning rate is set to $1\times 10^{-3}$ for all methods. We use DeepSpeed (Rasley et al., 2020) with mixed-precision training (BF16). The maximum sequence length is set to 1024 tokens, and we use a global batch size of 32. For practical efficiency, we implement our method using Hessian-vector product (HVP) and conjugate gradient methods, which efficiently approximate the Hessian matrix without the need for computing $d^{2}$ matrix. We use the default partition of the Ultrafeedback-binarized dataset with the standard train_prefs split for training and test_prefs split for evaluation. Furthermore, we also enable flash attention for the improved training efficiency.
6.1 Training Stage
(a) training loss
(b) training accuracy
(c) evaluation loss
(d) evaluation accuracy
Figure 1: For the training stage with passive data collection, we compare our method with MLE (Zhu et al., 2023). We report the average accuracy and loss of the reward model during training.
In this part, we conduct experiments to evaluate the performance of our method in the training stage, including both passive and active data collection strategies.
Contenders.
In the training stage, we compare with:
  • Random: A baseline that randomly selects queries.
  • Passive-MLE (Zhu et al., 2023): A method that uses maximum likelihood estimation to learn the reward function and compute the pessimistic policy based on the learned reward model.
  • Active-MLE (Das et al., 2024): A method using MLE to optimize the reward model and selects queries based on uncertainty, i.e., the maximum difference of two actions.
We compared our method with other methods in terms of the accuracy of the reward model, as the method for training the final policy based on the learned reward model is the same as contenders.
Passive Data Collection Results.
We first evaluate our method in the passive data collection setting. We randomly sample $T=30,000$ data points from the Ultrafeedback dataset for training. Figure 1 shows the loss and accuracy vs. the number of training samples. We observe that our method converges faster to a lower loss and achieves a higher evaluation accuracy compared to baselines. The improvement is particularly pronounced in the small-sample regime ($T<10,000$), where our method achieves a higher evaluation accuracy with the same amount of samples compared to Passive-MLE which employs conventional stochastic gradient descent (SGD) updates. This demonstrates the superior statistical efficiency of our approach, therefore achieving a better performance with fewer training samples.
Active Data Collection Results.
We then evaluate our method in the active data collection setting. In this scenario, the global batch size is set to 8, and we only allow the method to select 6, 400 samples out of the whole training datasets for training according to different selection strategies. Table 2 and Figure 2 demonstrate that our active learning strategy achieves comparable performance to passive learning while using only 21% of the labeled data. Specifically, our method achieves an average accuracy of 70.43% after 6,400 actively selected samples, which is comparable with the previous state-of-the-art method, Active-MLE, while our method is 3 times faster because we do not need to recompute all the selected data in the buffer, but we update the model in an online manner. The training loss curves in Figure 2(a) also show that our method converges faster compared to other contenders. This improved sample efficiency can be attributed to our theoretically-motivated query selection criterion that effectively identifies the most informative samples, which further validates the two key factors in our method: (1) the use of local norm $\mathcal{H}_{t}$ that better captures the uncertainty in different directions, and (2) the computationally efficient one-pass update rule that enables processing large-scale datasets.
(a) training loss
(b) evaluation accuracy
Figure 2: For the training stage with active data collection, we report (a) the training loss vs. the training iterations, and (b) the evaluation accuracy vs. the training iterations.
Method ACC (%) Time (s)
Rand-MLE 69.51$\pm$0.5 4876$\pm$47
Active-MLE 69.82$\pm$0.4 4982$\pm$52
Rand-OMD 68.97$\pm$0.6 1456$\pm$31
Ours 70.43$\pm$0.3 1489$\pm$36
Table 2: Comparison of different methods in terms of evaluation accuracy and training time in seconds. Results are averaged over five runs.
6.2 Deployment Stage
In this part, we evaluate our method on deployment stage.
Contenders.
In the deployment stage, we compare with:
  • Random: A baseline strategy that randomly select responses from responses generated by the current policy model.
  • Best-Two: A baseline strategy that selects the best two responses from responses generated by the current policy model.
  • Iterative-RLHF (Dong et al., 2024): A method that takes the best and the worst responses from the current policy model to construct a new preference pair for RLHF iteratively.
Deployment Stage Analysis.
For the scenario of deployment stage, we evaluate on a pre-processed online-variant of the Ultrafeedback-binarized dataset, where we divide the dataset into 20 chunks and process them sequentially to simulate the online deployment scenario. As shown in Figure 3, our method consistently outperforms the baselines in terms of both average cumulative rewards and win rates. The average cumulative rewards in Figure 3(a) show that our method learns more effectively from interactions, achieving higher rewards over time. The win rates in Figure 3(b) demonstrate that our method makes better decisions when selecting between response pairs, outperforming both the baseline methods of Random and Best-Two, and also better than Iterative-RLHF which selects the best and worst response pairs for training. The superior performance of our method further validates the effectiveness of our theoretically grounded action selection strategy. This success stems from an efficient online update mechanism that enables rapid adaptation to new preference feedback, as well as the design of the promising set, which balances high-reward responses for immediate performance with sufficient exploration for future model improvements. These results align with our theoretical analysis in Section 5, validating the practical effectiveness of our proposed algorithm in the deployment stage.
(a) average cumulative rewards
(b) win rates
Figure 3: For the deployment stage, we divide the Ultrafeedback dataset into 20 chunks, and report (a) the average cumulative rewards and the (b) win rates of different methods.
7 Conclusion and Future Work
In this work, we present a unified framework of the RLHF pipeline from the view of contextual bandits, leveraging the Bradley-Terry preference model with a linearly parameterized reward function. We decompose the RLHF process into two stages: (post-)training and deployment, exploring both passive and active data collection strategies during the training phase. We then develop novel algorithms for each stage, applying online mirror descent with a tailored local norm to replace standard maximum likelihood estimation, and designing distinct query criteria for different stages. Our theoretical analysis shows that our method improves upon existing methods in both statistical and computational efficiency. Finally, we train and deploy Llama-3-8B-Instruct on the Ultrafeedback-binarized dataset using our approach. The empirical results validate our theoretical findings and confirm the effectiveness of our proposed method.
While our work advances the theoretical understanding of RLHF, several important directions remain for future work. First, we considered the fixed feature mapping $\phi$ for the reward model, but in practice, this mapping may evolve during the training process. Thus, analyzing this dynamic feature mapping is an important future work. Second, though we focused on the Bradley-Terry model, there are other preference models for human feedback, such as the Plackett-Luce model (Luce, 1959; Plackett, 1975). Extending our results to these preference models is another promising direction.
Appendix A Useful Lemmas
Lemma 2.
For any $t\in[T]$, define the second-order approximation of the loss function $\ell_{t}(\theta)$ at the estimator $\widetilde{\theta}_{t}$ as
~$\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+\langle\nabla\ell_{t}(\widetilde{\theta}_{t}),\theta-\widetilde{\theta}_{t}\rangle+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}.$
Then, for the following update rule
$\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\left\{\widetilde{\ell}_{t}(\theta)+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{t}\rVert_{\mathcal{H}_{t}}^{2}\right\},$
it holds that
$\lVert\widetilde{\theta}_{t+1}-\theta^{*}\rVert_{\mathcal{H}_{t+1}}^{2}$
[…]
Proof.
Based on the analysis of (implicit) OMD update (see Lemma 5), for any $i\in[T]$, we have
~$\big{\langle}\nabla\widetilde{\ell}_{i}(\widetilde{\theta}_{i+1}),\widetilde{\theta}_{i+1}-\theta^{*}\big{\rangle}\leqslant\frac{1}{2\eta}\left(\|\widetilde{\theta}_{i}-\theta^{*}\|_{\mathcal{H}_{i}}^{2}-\|\widetilde{\theta}_{i+1}-\theta^{*}\|_{\mathcal{H}_{i}}^{2}-\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\|_{\mathcal{H}_{i}}^{2}\right)$
According to Lemma 6, we have
$\ell_{i}(\widetilde{\theta}_{i+1})-\ell_{i}\left(\theta^{*}\right)\leqslant\big{\langle}\nabla\ell_{i}(\widetilde{\theta}_{i+1}),\widetilde{\theta}_{i+1}-\theta^{*}\big{\rangle}-\frac{1}{\zeta}\big{\|}\widetilde{\theta}_{i+1}-\theta^{*}\big{\|}_{\nabla^{2}\ell_{i}(\widetilde{\theta}_{i+1})}^{2},$
where $\zeta=\log 2+2(LB+1)$. Then, by combining the above two inequalities, we have
$\ell_{i}(\widetilde{\theta}_{i+1})-\ell_{i}(\theta^{*})$ $\leqslant\langle\nabla\ell_{i}(\widetilde{\theta}_{i+1})-\nabla\widetilde{\ell}_{i}(\widetilde{\theta}_{i+1}),\widetilde{\theta}_{i+1}-\theta^{*}\rangle$
$\qquad+\frac{1}{\zeta}\Big{(}\|\widetilde{\theta}_{i}-\theta^{*}\|_{\mathcal{H}_{i}}^{2}-\|\widetilde{\theta}_{i+1}-\theta^{*}\|_{\mathcal{H}_{i+1}}^{2}-\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\|_{\mathcal{H}_{i}}^{2}\Big{)}.$
We can further bound the first term of the right-hand side as:
$\big{\langle}\nabla\ell_{i}(\widetilde{\theta}_{i+1})-\nabla\widetilde{\ell}_{i}(\widetilde{\theta}_{i+1}),\widetilde{\theta}_{i+1}-\theta^{*}\big{\rangle}=$ $\big{\langle}\nabla\ell_{i}(\widetilde{\theta}_{i+1})-\nabla\ell_{i}(\widetilde{\theta}_{i})-\nabla^{2}\ell_{i}(\widetilde{\theta}_{i})(\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}),\widetilde{\theta}_{i+1}-\theta^{*}\big{\rangle}$
=\ = $\big{\langle}D^{3}\ell_{i}({\xi}_{i+1})[\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}](\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}),\widetilde{\theta}_{i+1}-\theta^{*}\big{\rangle}$
[…]
where the second equality holds by the mean value theorem, the first inequality holds by the self-concordant-like property of $\ell_{i}(\cdot)$ in Lemma 3, and the last inequality holds by $\widetilde{\theta}_{i+1}$ and $\theta^{*}$ belong to $\Theta=\{\theta\in\mathbb{R}^{d},\lVert\theta\rVert_{2}\leq B\}$, and ${\nabla^{2}\ell_{i}({\xi}_{i+1})}\preceq L^{2}I_{d}$.
Then, by taking the summation over $i$ and rearranging the terms, we obtain
$\big{\|}\widetilde{\theta}_{t+1}-\theta^{*}\big{\|}_{\mathcal{H}_{t+1}}^{2}$
[…]
where the last inequality is by $\|\widetilde{\theta}_{1}-\theta^{*}\|_{\mathcal{H}_{1}}^{2}\leq\lambda\|\widetilde{\theta}_{1}-\theta^{*}\|_{2}^{2}\leq 4\lambda B^{2}$. Set $\zeta=2\eta$ finishes the proof. ∎
Appendix B Proof of Lemma 1
Proof.
Based on Lemma 2, we have
$\lVert\widetilde{\theta}_{t+1}-\theta^{*}\rVert_{\mathcal{H}_{t+1}}^{2}$
[…]
It remains to bound the right-hand side of the above inequality in the following. The most challenging part is to bound the term $\sum_{i=1}^{t}\ell_{i}(\theta^{*})-\sum_{i=1}^{t}\ell_{i}(\widetilde{\theta}_{i+1})$. This term might seem straightforward to control, as it can be observed that $\theta^{*}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\bar{\ell}(\theta)\triangleq\mathbb{E}_{y_{i}}[\ell_{i}(\theta)]$, where $\ell_{i}(\theta)$ serves as an empirical observation of ¯$\bar{\ell}(\theta)$. Consequently, the loss gap term seemingly can be bounded using appropriate concentration results. However, a caveat lies in the fact that the update of the estimator $\widetilde{\theta}_{i+1}$ depends on $\ell_{i}$, or more precisely $y_{i}$, making it difficult to directly apply such concentrations.
To address this issue, following the analysis in Zhang and Sugiyama [2023], we decompose the loss gap into two components by introducing an intermediate term. Specifically, we define the softmax function as
$[\sigma_{i}(q)]_{1}=\frac{\exp(q)}{1+\exp(q)}\quad\text{and}\quad[\sigma_{i}(q)]_{0}=\frac{1}{1+\exp(q)},$
where $[\cdot]_{i}$ denotes the $i$-th element of the vector. Then, the loss function $\ell_{i}(\theta)$ can be rewritten as
$\ell(q_{t},y_{t})=-\mathds{1}_{\{y_{t}=1\}}\cdot\log\left({[\sigma(q_{t})]_{1}}\right)-\mathds{1}_{\{y_{t}=0\}}\cdot\log\left({[\sigma(q_{t})]_{0}}\right).$
Then, we define the pseudo-inverse function of $\sigma^{-1}(p)$ with
$[\sigma^{-1}(p)]=\log({q}/({1-q}))\quad\text{and}\quad[\sigma^{-1}(p)]_{0}=\log((1-q)/{q}).$
Then, we decompose the regret into two terms by introducing an intermediate term.
$\sum_{i=1}^{t}\ell_{s}\left(\theta^{*}\right)-\sum_{i=1}^{t}\ell_{s}(\widetilde{\theta}_{i+1})=$ $\underbrace{\sum_{i=1}^{t}\ell_{s}\left(\theta^{*}\right)-\sum_{i=1}^{t}\ell_{s}(q_{i},y_{i})}_{\mathtt{term~{}(a)}}+\underbrace{\sum_{i=1}^{t}\ell_{s}(q_{i},y_{i})-\sum_{i=1}^{t}\ell_{s}(\widetilde{\theta}_{i+1})}_{\mathtt{term~{}(b)}}$
where $q_{i}$ is an aggregating forecaster for logistic loss defined by $q_{i}=\sigma^{-1}(\mathbb{E}_{\theta\sim P_{i}}[\sigma(\theta^{\top}z_{i})])$ and $P_{i}=\mathcal{N}(\widetilde{\theta}_{i},(1+c\mathcal{H}_{i}^{-1}))$ is the Gaussian distribution with mean $\widetilde{\theta}_{i}$ and covariance $(1+c\mathcal{H}_{i}^{-1})$, where $c>0$ is a constant to be specified later. It remains to bound the terms $\mathtt{term~{}(a)}$ and $\mathtt{term~{}(b)}$, which were initially analyzed in Zhang and Sugiyama [2023] and further refined by Lee and Oh [2024]. Specifically, using Lemmas F.2 and F.3 in Lee and Oh [2024], we can bound them as follows.
For $\mathtt{term~{}(a)}$, let $\delta\in(0,1)$ and $\lambda\geq 1$. With probability at least $1-\delta$, for all $t\in[T]$, we have
$\mathtt{term~{}(a)}\leq\left(3\log(1+2t)+2+LB\right)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right)+2.$
For $\mathtt{term~{}(b)}$, let $\lambda\geq\max\{2,72cd\}$. Then, for all $t\in[T]$, we have
$\mathtt{term~{}(b)}\leq\frac{1}{2c}\sum_{i=1}^{t}\Big{\|}\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\Big{\|}_{\mathcal{H}_{i}}^{2}+\sqrt{6}cd\log\left(1+\frac{2tB^{2}}{d\lambda}\right)$
Combing the above two bounds, we have
$\big{\|}\widetilde{\theta}_{t+1}-\theta^{*}\big{\|}_{\mathcal{H}_{t+1}}^{2}$ $\leq 2\eta\left(3\log(1+2t)+2+BL\right)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right)$
$+4\eta+2\eta\sqrt{6}cd\log\left(1+\frac{2tL^{2}}{d\lambda}\right)+4\lambda L_{\theta}^{2}+12\sqrt{2}BL^{3}\eta\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{2}^{2}+\left(\frac{\eta}{c}-1\right)\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{\mathcal{H}_{i}}^{2}.$
Setting $c=7\eta/6$ and $\lambda\geq 84\sqrt{2}BL^{3}\eta$, we have
$12\sqrt{2}BL^{3}\eta\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{2}^{2}+\left(\frac{\eta}{c}-1\right)\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{\mathcal{H}_{i}}^{2}\leq\left(12\sqrt{2}BL^{3}\eta-\frac{\lambda}{7}\right)\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{2}^{2}\leq 0$
Note that $84\sqrt{2}\left(BL^{3}+dL^{2}\right)\eta\geq\max\left\{2L^{2},72cdL^{2},84\sqrt{2}BL^{3}\eta\right\}$, so we set $\lambda\geq 84\sqrt{2}\left(BL^{3}+\right.$ $\left.dL^{2}\right)\eta$. As we have $\eta=(1/2)\log 2+\left(BL+1\right)$, we have
$\big{\|}\widetilde{\theta}_{t+1}-\theta^{*}\big{\|}_{\mathcal{H}_{t+1}}\leq\mathcal{O}\Big{(}\sqrt{d}(\log(t/\delta))^{2}\Big{)}.$
This finishes the proof. ∎
Appendix C Proof of Theorem 1
Proof.
Define $J(\pi)=\mathbb{E}_{x\sim\rho}[r\left(x,\pi(x)\right)]$, $J^{\prime}(\pi)=J(\pi)-\langle\theta^{*},v\rangle$, we have
$\operatorname*{\mathsf{SubOpt}}\left(\pi_{T}\right)=\left(J^{\prime}\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)\right)+\left(\widetilde{J}\left(\pi^{*}\right)-\widetilde{J}\left(\pi_{T}\right)\right)+\left(\widetilde{J}\left(\pi_{T}\right)-J^{\prime}\left(\pi_{T}\right)\right).$
Since $\pi_{T}$ is the optimal policy under expected value $\widetilde{J}(\pi)$, i.e., $\widetilde{J}(\pi_{T})=\max_{\pi\in\Pi}\widetilde{J}(\pi)$, we have
$\widetilde{J}\left(\pi^{*}\right)-\widetilde{J}\left(\pi_{T}\right)\leq 0$ (18)
For the third term, we have with probability at least $1-\delta$, it holds that
$\widetilde{J}\left(\pi_{T}\right)-J^{\prime}\left(\pi_{T}\right)=\min_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\theta^{\top}(\phi(s,\pi_{T}(s))-v)\right]-\mathbb{E}_{x\sim\rho}\left[\theta^{*\top}(\phi(s,\pi_{T}(s))-v)\right]\leq 0,$ (19)
where the last inequality holds by $\theta^{*}\in\mathcal{C}_{T}$ with probability at least $1-\delta$.
For the first term, we have with probability at least $1-\delta$, it holds that
$J^{\prime}\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)$
=\ = $\mathbb{E}_{x\sim\rho}\left[(\theta^{*})^{\top}(\phi(s,\pi^{*}(s))-v)\right]-\min_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\theta^{\top}(\phi(s,\pi^{*}(s))-v)\right]$
=\ = $\sup_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\left(\theta^{*}-\widetilde{\theta}_{T}+\widetilde{\theta}_{T}-\theta\right)^{\top}\left(\phi(x,\pi^{*}(x))-v\right)\right]$
=\ = $\mathbb{E}_{x\sim\rho}\left[\left(\theta^{*}-\widetilde{\theta}_{T}\right)^{\top}\left(\phi(x,\pi^{*}(x))-v\right)\right]+\sup_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\left(\widetilde{\theta}_{T}-\theta\right)^{\top}\left(\phi(x,\pi^{*}(x))-v\right)\right]$
[…]
where the first inequality holds by the Cauchy-Schwarz inequality.
Since it holds $\widetilde{\theta}_{T}\in\mathcal{C}_{T}$ with probability at least $1-\delta$ by Lemma 1, we have $\lVert\theta^{*}-\widetilde{\theta}_{T}\rVert_{\mathcal{H}_{T}}\leq\widetilde{\beta}_{T}$ and $\sup_{\theta\in\mathcal{C}_{T}}\lVert\theta-\widetilde{\theta}_{T}\rVert_{\mathcal{H}_{T}}\leq\widetilde{\beta}_{T}$. Thus, we obtain
$J^{\prime}\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)\leq$ $2\widetilde{\beta}_{T}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}.$ (20)
Combining Eq. (18), Eq. (19), and Eq. (20) and substituting $\widetilde{\beta}_{T}=\mathcal{O}(\sqrt{d}(\log(T/\delta))^{2})$, we have with probability at least $1-\delta$, it holds that
$\operatorname*{\mathsf{SubOpt}}\left(\pi_{T}\right)$ $=\left(J^{\prime}\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)\right)+\left(\widetilde{J}\left(\pi^{*}\right)-\widetilde{J}\left(\pi_{T}\right)\right)+\left(\widetilde{J}\left(\pi_{T}\right)-J^{\prime}\left(\pi_{T}\right)\right)$
$\leq 2\widetilde{\beta}_{T}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}$
$\leq\mathcal{O}\left(\sqrt{d}\left(\log\frac{T}{\delta}\right)^{2}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}\right).$
This completes the proof. ∎
Appendix D Proof of Theorem 2
Proof.
Let the sub-optimality gap for a context $x\in\mathcal{X}$ be denoted as $\operatorname*{\mathsf{SubOpt}}(x)$. Thus, for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
$\operatorname*{\mathsf{SubOpt}}(x)$ $=\left(\phi\left(x,\pi^{*}(x)\right)-\phi\left(x,\pi_{T}(x)\right)\right)^{\top}\theta^{*}$
$\leq\left(\phi\left(x,\pi^{*}(x)\right)-\phi\left(x,\pi_{T}(x)\right)\right)^{\top}\theta^{*}+\left(\phi\left(x,\pi_{T}(x)\right)-\phi\left(x,\pi^{*}(x)\right)\right)^{\top}\left(\frac{1}{T}\sum_{t=1}^{T}\widetilde{\theta}_{t}\right)$
$=\left(\phi\left(x,\pi^{*}(x)\right)-\phi\left(x,\pi_{T}(x)\right)\right)^{\top}\left(\theta^{*}-\frac{1}{T}\sum_{t=1}^{T}\widetilde{\theta}_{t}\right)$
$=\frac{1}{T}\sum_{t=1}^{T}\left(\phi\left(x,\pi^{*}(x)\right)-\phi\left(x,\pi_{T}(x)\right)\right)^{\top}\left(\theta^{*}-\widetilde{\theta}_{t}\right)$
$\leq\frac{1}{T}\sum_{t=1}^{T}\left\|\phi\left(x,\pi^{*}(x)\right)-\phi\left(x,\pi_{T}(x)\right)\right\|_{\mathcal{H}_{t}^{-1}}\left\|\theta^{*}-\widetilde{\theta}_{t}\right\|_{\mathcal{H}_{t}}$
$\leq\frac{\widetilde{\beta}_{T}}{T}\sum_{t=1}^{T}\left\|\phi\left(x,\pi^{*}(x)\right)-\phi\left(x,\pi_{T}(x)\right)\right\|_{\mathcal{H}_{t}^{-1}},$
where the first inequality is due to the fact that $\left(\phi\left(x,\pi_{T}(x)\right)-\phi\left(x,\pi^{*}(x)\right)\right)^{\top}\left(\frac{1}{T}\sum_{t=1}^{T}\widetilde{\theta}_{t}\right)\geq 0$ by the design of $\pi_{T}(x)$, the second is due to the Cauchy-Schwarz inequality, and the last inequality is due to $\|\theta^{*}-\widetilde{\theta}_{t}\|_{\mathcal{H}_{t}}\leq\beta_{T}$ with probability at least $1-\delta$ by Lemma 1.
By our algorithm’s choice $(x_{t},a_{t},a_{t}^{\prime})=\operatorname*{arg\,max}_{x\in\mathcal{X},a,a^{\prime}\in\mathcal{A}}\left\|\phi(x,a)-\phi\left(x,a^{\prime}\right)\right\|_{\mathcal{H}_{t}^{-1}}$, we have
$\sum_{t=1}^{T}\left\|\phi\left(x,\pi^{*}(x)\right)-\phi\left(x,\pi_{T}(x)\right)\right\|_{\mathcal{H}_{t}^{-1}}\leq\sum_{t=1}^{T}\left\|\phi\left(x_{t},a_{t}\right)-\phi\left(x_{t},a_{t}^{\prime}\right)\right\|_{\mathcal{H}_{t}^{-1}}=\sum_{t=1}^{T}\left\|z_{t}\right\|_{\mathcal{H}_{t}^{-1}}.$
Furthermore, by the definition of $\mathcal{H}_{t}$, we have
$\mathcal{H}_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}\dot{\sigma}\left(z_{s}^{\top}\widetilde{\theta}_{s+1}\right)z_{s}z_{s}^{\top}\geq\lambda I_{d}+\frac{1}{\kappa}\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}=\frac{1}{\kappa}\left(\kappa\lambda I_{d}+\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}\right)\triangleq\frac{1}{\kappa}V_{t}.$
Thus, we have
$\sum_{t=1}^{T}\left\|z_{t}\right\|_{\mathcal{H}_{t}^{-1}}\leq\sqrt{\kappa}\sum_{t=1}^{T}\left\|z_{t}\right\|_{V_{t}^{-1}}\leq\sqrt{\kappa}\sqrt{T\sum_{t=1}^{T}\left\|z_{t}\right\|^{2}_{V_{t}^{-1}}}\leq\sqrt{2\kappa dT\log\left(1+\frac{4TL^{2}}{\lambda\kappa d}\right)},$
where the first inequality holds by the fact that $\mathcal{H}_{t}\succeq\frac{1}{\kappa}V_{t}$, the second inequality holds by the Cauchy-Schwarz inequality, and the last inequality holds by the elliptic potential lemma in Lemma 4. Thus, we have for any context $x\in\mathcal{X}$,
$\operatorname*{\mathsf{SubOpt}}(x)\leq\frac{\widetilde{\beta}_{T}}{T}\sqrt{2\kappa dT\log\left(1+\frac{4TL^{2}}{\lambda\kappa d}\right)}.$
By the definition of $\operatorname*{\mathsf{SubOpt}}(\pi_{T})$, we have with probability at least $1-\delta$,
$\operatorname*{\mathsf{SubOpt}}\left(\pi_{T}\right)=\mathbb{E}_{x\sim\rho}\left[\operatorname*{\mathsf{SubOpt}}(x)\right]\leq\frac{\widetilde{\beta}_{T}}{T}\sqrt{2\kappa dT\log\left(1+\frac{T}{\lambda\kappa d}\right)}\leq\widetilde{\mathcal{O}}\left(d\sqrt{\frac{\kappa}{T}}\right).$
This finishes the proof. ∎
Appendix E Proof of Theorem 3
Proof.
We first analyze the instantaneous regret at round $t$. For any $\delta\in(0,1)$, with probability at least $1-\delta$, it holds that
$\big{(}r(x_{t},\pi^{*}(x_{t}))-r(x_{t},a_{t})\big{)}+\big{(}r(x_{t},\pi^{*}(x_{t}))-r(x_{t},a_{t}^{\prime})\big{)}$
=\ = $\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\theta^{*}+\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}\theta^{*}$
=\ = $\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\widetilde{\theta}_{t}+\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}(\theta^{*}-\widetilde{\theta}_{t})$
$+\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}\widetilde{\theta}_{t}+\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}(\theta^{*}-\widetilde{\theta}_{t})$
[…]
$+\widetilde{\beta}_{t}\big{\lVert}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a^{\prime}_{t})\big{\rVert}_{\mathcal{H}_{t}^{-1}}+\big{\lVert}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a^{\prime}_{t})\big{\rVert}_{\mathcal{H}_{t}^{-1}}\big{\lVert}\theta^{*}-\widetilde{\theta}_{t}\big{\rVert}_{\mathcal{H}_{t}}$
[…]
$+\widetilde{\beta}_{t}\lVert\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\rVert_{\mathcal{H}_{t}^{-1}}+\widetilde{\beta}_{t}\lVert\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\rVert_{\mathcal{H}_{t}^{-1}}$
[…]
Here, the first inequality holds since both $a_{t},a^{\prime}_{t}\in\mathcal{A}_{t}$ and by the definition of $\mathcal{A}_{t}$, it implies that
$\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\widetilde{\theta}_{t}$ $\leq\widetilde{\beta}_{t}\big{\lVert}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{\rVert}_{\mathcal{H}_{t}^{-1}},$
$\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a^{\prime}_{t})\big{)}^{\top}\widetilde{\theta}_{t}$ $\leq\widetilde{\beta}_{t}\big{\lVert}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big{\rVert}_{\mathcal{H}_{t}^{-1}}.$
The second inequality holds by selection strategy of $(a_{t},a_{t}^{\prime})$, it holds that
$\lVert\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\rVert_{\mathcal{H}_{t}^{-1}}$ $\leq\lVert\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\rVert_{\mathcal{H}_{t}^{-1}},$
$\lVert\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a^{\prime}_{t})\rVert_{\mathcal{H}_{t}^{-1}}$ $\leq\lVert\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\rVert_{\mathcal{H}_{t}^{-1}}.$
Thus, we have
${\operatorname*{Reg}}_{T}=\frac{1}{2}\Big{(}\big{(}r(x_{t},\pi^{*}(x_{t}))-r(x_{t},a_{t})\big{)}+\big{(}r(x_{t},\pi^{*}(x_{t}))-r(x_{t},a_{t}^{\prime})\big{)}\Big{)}\leq 2\widetilde{\beta}_{T}\sum_{t=1}^{T}\lVert\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\rVert_{\mathcal{H}_{t}^{-1}}.$
Furthermore, by the definition of $\mathcal{H}_{t}$, we have
$\mathcal{H}_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}\dot{\sigma}\left(z_{s}^{\top}\widetilde{\theta}_{s+1}\right)z_{s}z_{s}^{\top}\geq\lambda I_{d}+\frac{1}{\kappa}\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}=\frac{1}{\kappa}\left(\kappa\lambda I_{d}+\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}\right)\triangleq\frac{1}{\kappa}V_{t}$
Thus, we obtain
$\sum_{t=1}^{T}\left\|z_{t}\right\|_{\mathcal{H}_{t}^{-1}}\leq\sqrt{\kappa}\sum_{t=1}^{T}\left\|z_{t}\right\|_{V_{t}^{-1}}\leq\sqrt{\kappa}\sqrt{T\sum_{t=1}^{T}\left\|z_{t}\right\|^{2}_{V_{t}^{-1}}}\leq\sqrt{2\kappa dT\log\left(1+\frac{4TL^{2}}{\lambda\kappa d}\right)},$
where the first inequality holds by the fact that $\mathcal{H}_{t}\succeq\frac{1}{\kappa}V_{t}$, the second inequality holds by the Cauchy-Schwarz inequality, and the last inequality holds by the elliptic potential lemma in Lemma 4. Thus, we have with probability at least $1-\delta$,
${\operatorname*{Reg}}_{T}\leq 2{\widetilde{\beta}_{T}}\sqrt{2\kappa dT\log\left(1+\frac{4TL^{2}}{\lambda\kappa d}\right)}\leq\widetilde{\mathcal{O}}\left(d\sqrt{{\kappa}{T}}\right).$
This finishes the proof. ∎
Appendix F Supporting Lemmas
Definition 3 (Tran-Dinh et al. [2015]).
A convex function $f\in\mathcal{C}^{3}\left(\mathbb{R}^{m}\right)$ is $M$-self-concordant-like function if
$\left|\psi^{\prime\prime\prime}(s)\right|\leqslant M\|\mathbf{b}\|_{2}\psi^{\prime\prime}(s),$
for $s\in\mathbb{R}$ and $M>0$, where $\psi(s):=f(\mathbf{a}+s\mathbf{b})$ for any $\mathbf{a},\mathbf{b}\in\mathbb{R}^{m}$.
Lemma 3 (Lee and Oh [2024, Proposition C.1]).
The loss $\ell_{t}(\theta)$ defined in Eq. (4) is $3\sqrt{2}L$-self-concordant-like for $\forall t\in[T]$.
Lemma 4 (Abbasi-Yadkori et al. [2011, Lemma 11]).
Suppose $x_{1},\ldots,x_{t}\in\mathbb{R}^{d}$ and for any $1\leq s\leq t$, $\lVert x_{s}\rVert_{2}\leq L$. Let $V_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}x_{s}x_{s}^{\top}$ for $\lambda\geq 0$. Then, we have
$\sum_{s=1}^{t}\left\|z_{s}\right\|_{V_{s}^{-1}}^{2}\leq 2d\log\left(1+\frac{tL^{2}}{\lambda d}\right).$
Lemma 5 (Campolongo and Orabona [2020, Proposition 4.1]).
Define $\mathbf{w}_{t+1}$ as the solution of
$\mathbf{w}_{t+1}=\operatorname*{arg\,min}_{\mathbf{w}\in\mathcal{V}}\big{\{}\eta\ell_{t}(\mathbf{w})+\mathcal{D}_{\psi}\left(\mathbf{w},\mathbf{w}_{t}\right)\big{\}},$
[…]
$\left\langle\eta_{t}\mathbf{g}_{t}^{\prime},\mathbf{w}_{t+1}-\mathbf{u}\right\rangle\leq\left\langle\nabla\psi\left(\mathbf{w}_{t}\right)-\nabla\psi\left(\mathbf{w}_{t+1}\right),\mathbf{w}_{t+1}-\mathbf{u}\right\rangle$
for any $\mathbf{u}\in\mathcal{W}$.
Lemma 6 (Zhang and Sugiyama [2023, Lemma 1]).
Let $\ell(\mathbf{z},y)=\sum_{k=0}^{K}\mathbf{1}\{y=k\}\cdot\log\left(\frac{1}{[\sigma(\mathbf{z})]_{k}}\right)$ where $\sigma(\mathbf{z})_{k}=\frac{e^{z_{k}}}{\sum_{j=0}^{K}e^{z_{j}}}$, $\mathbf{a}\in[-C,C]^{K}$, $y\in\{0\}\cup[K]$ and $\mathbf{b}\in\mathbb{R}^{K}$ where $C>0$. Then, we have
$\ell(\mathbf{a},y)\geq\ell(\mathbf{b},y)+\nabla\ell(\mathbf{b},y)^{\top}(\mathbf{a}-\mathbf{b})+\frac{1}{\log(K+1)+2(C+1)}(\mathbf{a}-\mathbf{b})^{\top}\nabla^{2}\ell(\mathbf{b},y)(\mathbf{a}-\mathbf{b}).$
Appendix G Details of Experiments
In this section, we provide the omitted details of the experiment details and additional results.
G.1 Implementation Details
Datasets.   We use the UltraFeedback-binarized dataset [Rafailov et al., 2023] for the experiments. This dataset is derived from the original UltraFeedback dataset, which comprises 64, 000 prompts sourced from diverse datasets including UltraChat, ShareGPT, Evol-Instruct, TruthfulQA, FalseQA, and FLAN. For each prompt, four model completions were generated using various open-source and proprietary language models, with GPT-4 providing comprehensive evaluations across multiple criteria including helpfulness, honesty, and truthfulness. The binarized version was constructed by selecting the completion with the highest overall score as the ”chosen” response and randomly selecting one of the remaining completions as the ”rejected” response, creating clear preference pairs suitable for reward modeling and direct preference optimization. This dataset structure aligns well with our experimental setup, providing a robust foundation for evaluating different preference learning approaches. The dataset’s diverse prompt sources and evaluation criteria make it particularly valuable for training and evaluating reward models in a real-world context. To further tailor the dataset to our experimental setup, we organize the dataset as follows for the training and deployment stages:
  • Training stage with passive data collection: We randomly sample $T=30,000$ data points from the UltraFeedback-binarized dataset’s train_prefs split for training. Each data point consists of a prompt and two responses with a binary preference label indicating the preferred response. We use the test_prefs split for evaluation.
  • Training stage with active data collection: We allow the method to actively select 6,400 samples from the train_prefs split according to different selection strategies. The global batch size is set to 8 for training. The selection is performed iteratively, where in each iteration, the method selects the most informative samples based on its selection criterion.
  • Deployment stage: We use a pre-processed online variant of the UltraFeedback-binarized dataset from the test_gen split. The dataset is divided into 20 sequential chunks to simulate an online deployment scenario. For each chunk, we generate responses using the current policy (the foundation model of policy model is chosen to be meta-llama / Llama-3.2-1B), evaluate them using both the learned reward model and an oracle reward model. We choose NCSOFT/Llama-3-OffsetBias-RM-8B [Park et al., 2024] as the oracle reward model. After each chunk, we use the policy model to randomly generate 64 responses using different seeds. We then apply various strategies (Random, Best-Two, etc.) to select responses and construct new preference pairs, which are then used to update the reward model and the policy model.
Algorithm 4 Efficient Update using Hessian-Vector Product with Conjugate Gradient
0:  Current parameter $\widetilde{\theta}_{t}$, gradient $g_{t}(\widetilde{\theta}_{t})$, learning rate $\eta$, max CG steps $K$, base damping $\lambda_{0}$, error tolerance $\epsilon$
1:  Initialize $v_{0}=0$, $r_{0}=g_{t}(\widetilde{\theta}_{t})$, $p_{0}=r_{0}$
2:  Compute damping $\lambda_{t}=\lambda_{0}\cdot\min\{1,f(t/T)\}$
3:  for $k=0,1,\ldots,K-1$ do
4:     Compute HVP: ~$\widetilde{\mathcal{H}}_{t}p_{k}=\nabla_{\theta}(\nabla_{\theta}\mathcal{L}(\theta)^{\top}p_{k})|_{\theta=\widetilde{\theta}_{t}}+\lambda_{t}p_{k}$
5:     $\alpha_{k}=\frac{r_{k}^{\top}r_{k}}{p_{k}^{\top}\widetilde{\mathcal{H}}_{t}p_{k}}$, $v_{k+1}=v_{k}+\alpha_{k}p_{k}$, $r_{k+1}=r_{k}-\alpha_{k}\widetilde{\mathcal{H}}_{t}p_{k}$,
6:     $\beta_{k+1}=\frac{r_{k+1}^{\top}r_{k+1}}{r_{k}^{\top}r_{k}}$, $p_{k+1}=r_{k+1}+\beta_{k+1}p_{k}$
7:     if $\|r_{k+1}\|\leq\epsilon$ then
8:        break
9:     end if
10:  end for
11:  Update parameter: $\widetilde{\theta}_{t+1}=\widetilde{\theta}_{t}-\eta v_{K}$
11:  Updated parameter $\widetilde{\theta}_{t+1}$
Practical Implementation.   To efficiently implement the OMD update in Eq. (9) without the costly computing and storing the full Hessian matrix, we utilize the Hessian-vector product (HVP) method combined with conjugate gradient descent. The key insight is that the update can be reformulated as solving a linear system. For the OMD update:
$\widetilde{\theta}_{t+1}$ $=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\big{\langle}g_{t}(\widetilde{\theta}_{t}),\theta\big{\rangle}+\frac{1}{2\eta}\big{\|}\theta-\widetilde{\theta}_{t}\big{\|}_{\widetilde{\mathcal{H}}_{t}}^{2}\Big{\}}\approx\widetilde{\theta}_{t}-\eta\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t}),$
where we ignore the projection operation. The key challenge is computing ~$\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t})$ efficiently. Instead of explicitly computing ~$\widetilde{\mathcal{H}}_{t}^{-1}$, we solve the linear system ~$\widetilde{\mathcal{H}}_{t}v=g_{t}(\widetilde{\theta}_{t})$
~$\widetilde{\mathcal{H}}_{t}v=g_{t}(\widetilde{\theta}_{t}),$
where $v$ is the solution we seek. We use the conjugate gradient method with a damped Hessian to solve this system iteratively. The HVP operation ~$\widetilde{\mathcal{H}}_{t}v$ for any vector $v$ can be computed efficiently using automatic differentiation ~$\widetilde{\mathcal{H}}_{t}v=\nabla_{\theta}\big{(}\nabla_{\theta}\mathcal{L}(\theta)^{\top}v\big{)}\big{|}_{\theta=\widetilde{\theta}_{t}}+\lambda v$, where $\lambda$ is a damping coefficient that ensures numerical stability and $\mathcal{L}(\theta)$ is the loss function. The full algorithm is summarized in Algorithm 4. To ensure stability and convergence, we employ an adaptive damping strategy as $\lambda_{t}=\lambda_{0}\cdot\min\{1,f(t/T)\}$, where $f(\cdot)$ is a damping growth function that can be linear, logarithmic, or cosine. In practice, we set $K=3$ and $\lambda_{0}=0.8$ for all experiments.
G.2 Validating the Magnitude of $\kappa$
We validate the magnitude of $\kappa$ in Eq. (2) by computing its value during the training process. The results show that $\kappa=171.62\pm 85.49$ during our training process, which is relatively large.
G.3 Combined with Adam Optimizer
(a) training loss
(b) training accuracy
(c) evaluation loss
(d) evaluation accuracy
Figure 4: For the training stage with passive data collection, we compare our proposed method and MLE [Zhu et al., 2023] in with passive data collection combined with Adam. We report the average accuracy and loss of the reward model during the training process.
In previous experiments, we used SGD to update model parameters. In this section, we integrate the methods with the Adam optimizer [Kingma and Ba, 2015], i.e., adding the first and second momentum terms to the model updates. The results, shown in Figure 4, indicate that the Adam optimizer further enhances the performance of our method by leveraging the momentum term to accelerate convergence. With the momentum term, our method remains superior to the MLE-based method; however, the performance gap is reduced. We think that this is because the Adam optimizer incorporates second-order information for optimization, diminishing the advantage of our method compared to the SGD cases.
G.4 Comparison with DPO
Figure 5: Comparison with DPO in the deployment stage.
We also compare with DPO (Direct Preference Optimization) [Rafailov et al., 2023] in the deployment stage, as a reward-free method, DPO optimizes the policy directly using preference feedback without explicit reward modeling. To ensure a fair comparison, we initialize the policy with 400 samples and use the same dataset settings as PPO to iteratively update the policy model using the DPO algorithm. The results are illustrated in Figure 5. While DPO outperforms the random baseline (Rand-MLE), it achieves lower cumulative rewards than the PromisingSet-based method. This result suggests that DPO’s online learning capability remains a challenge. In contrast, the reward model learned by our selection strategy effectively learned streaming data and continuously updates the policy as new data arrive, indicating that in our deployment stage, a reward model with PPO may be a more suitable choice for sequentially learning from new data.

### v2

@@
-Provably Efficient RLHF Pipeline: A Unified View from Contextual Bandits
+Provably Efficient Online RLHF with One-Pass Reward Modeling
 Abstract
-Reinforcement Learning from Human Feedback (RLHF) is a widely used approach for aligning Large Language Models (LLMs) with human preferences.
-While recent advancements have provided valuable insights into various stages and settings of RLHF, a comprehensive theoretical understanding of the entire RLHF pipeline remains lacking.
-Towards this end, we propose a unified framework for the RLHF pipeline from the view of contextual bandits and provide provable efficiency guarantees.
-In particular, we decompose the RLHF process into two distinct stages: (post-)training and deployment, exploring both passive and active data collection strategies during the training phase.
-By employing the Bradley-Terry preference model with a linearly parameterized reward function, we reformulate RLHF as a contextual preference bandit problem.
-We then develop novel algorithms for each stage, demonstrating significant improvements over existing approaches in both statistical and computational efficiency.
-Finally, we apply our method to train and deploy Llama-3-8B-Instruct on the Ultrafeedback-binarized dataset, and empirical results confirm the effectiveness of our approach.
+Reinforcement Learning from Human Feedback (RLHF) has shown remarkable success in aligning Large Language Models (LLMs) with human preferences.
+Traditional RLHF approaches rely on a fixed dataset, which often suffers from limited coverage.
+To this end, online RLHF has emerged as a promising direction, enabling iterative data collection and model improvement.
+Despite its potential, this paradigm faces a key bottleneck: the requirement to continuously integrate new data into the historical dataset and re-optimize the model from scratch at each iteration, resulting in computational and storage costs that grow linearly with the number of iterations.
+In this work, we address this challenge by proposing a one-pass reward modeling method that does not require storing the historical data and can be computed in constant time.
+Specifically, we first formalize RLHF as a contextual preference bandit problem and design an online mirror descent algorithm with a tailored local norm to replace the standard maximum likelihood estimation for reward modeling.
+We then apply our method to various online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation.
+We provide theoretical guarantees showing that our method improves both statistical and computational efficiency.
+Finally, we provide practical algorithms and conduct experiments using Llama-3-8B-Instruct and Qwen2.5-7B-Instruct models on the Ultrafeedback-binarized and Mixture2 datasets, validating the effectiveness of our proposed method.
 1 Introduction
-Table 1: Comparison between previous works and approach in terms of the statistical and computational efficiency across different stages of the RLHF pipeline.
-Our pipeline consists of two stages: training and deployment, where we explore both passive and active data collection strategies in the training stage.
-The statistical efficiency is measured using the sub-optimality gap (for the training stage) and regret (for the deployment stage), while the computational efficiency is evaluated based on time and storage complexity.
-Here, $d$ is the dimension of the feature space, $T$ is the number of samples, $\kappa$ is the non-linearity coefficient which can be exponentially large, $\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)$ is the concentratability vector, $V_{T}$ and $\mathcal{H}_{T}$ are two kinds of local norm that satisfies $\mathcal{H}_{T}\succeq(1/\kappa)V_{T}$.
-Stage Data Reference Sub-optimality Gap/Regret Time Storage
-Training Passive Zhu et al.
-(2023) $\widetilde{\mathcal{O}}\Big{(}\kappa\sqrt{d}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{V_{T}^{-1}}\Big{)}$ $\mathcal{O}(T\log T)$ $\mathcal{O}(T)$
-Ours (Theorem 1) $\widetilde{\mathcal{O}}\Big{(}\sqrt{d}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}\Big{)}$ $\mathcal{O}(T)$ $\mathcal{O}(T)$
-Active Das et al. (2024) $\widetilde{\mathcal{O}}\big{(}d\sqrt{\kappa/T}\big{)}$ $\mathcal{O}(T^{2}\log T)$ $\mathcal{O}(T)$
-Ours (Theorem 2) $\widetilde{\mathcal{O}}\big{(}d\sqrt{\kappa/T}\big{)}$ $\mathcal{O}(T)$ $\mathcal{O}(1)$
-Deployment Online Saha (2021) $\widetilde{\mathcal{O}}\big{(}d\kappa\sqrt{T}\big{)}$ $\mathcal{O}(T^{2}\log T)$ $\mathcal{O}(T)$
-Ours (Theorem 3) $\widetilde{\mathcal{O}}\big{(}d\sqrt{\kappa T}\big{)}$ $\mathcal{O}(T)$ $\mathcal{O}(1)$
-Reinforcement Learning from Human Feedback is a key technique for training large language models using human feedback (Ouyang et al., 2022; Bai et al., 2022).
-The RLHF process involves collecting data, each consisting of a prompt, a pair of responses, and a human preference label indicating the preferred response.
-Then, a reward model is trained to predict human preferences, and the LLM is fine-tuned based on the reward model by RL algorithms, such as PPO (Schulman et al., 2017).
-Given the notable success of RLHF, recent efforts have been devoted to developing a deeper theoretical understanding of this approach.
-Zhu et al.
-(2023) investigated the standard offline setting, where the learner is provided with a fixed dataset and aims to learn a policy that maximizes the expected reward.
-In this setting, since the learner has no control over the data collection process, the quality of the dataset becomes crucial to the performance of the learned policy and the resulting policy often performs poorly when faced with out-of-distribution data (Burns et al., 2024).
-In practice, the Claude (Bai et al., 2022) and LLaMA-2 (Touvron et al., 2023) projects have demonstrated that iterative RLHF can significantly enhance model performance.
-Motivated by these practical advancements, Xiong et al.
-(2024) investigated the iterative setting, where the learner randomly selects prompts from the candidate set, generates response pairs by sampling from the current policy, and then queries for feedback to update the policy accordingly.
-They showed that RLHF benefits from this iterative data collection process.
-While uniform prompt sampling as a simple approach has been proven effective, Das et al.
-(2024) demonstrated that it can lead to inefficient sampling and the sub-optimality gap may fail to decrease as the number of samples increases.
-To this end, they studied the active setting where the learner is allowed to select both prompts and responses, query for human feedback, and update the policy accordingly.
-Though these works have provided valuable insights into various stages and settings of RLHF, a comprehensive theoretical understanding of the entire RLHF pipeline remains lacking.
-Towards this end, in this work, we propose a unified framework for the RLHF pipeline from the view of contextual bandits.
-Specifically, building on recent theoretical advancements (Zhu et al., 2023; Xiong et al., 2024), we leverage the Bradley-Terry preference model with a linearly parameterized reward function to reformulate RLHF as a contextual bandit problem (Abe and Long, 1999).
-We then decompose the RLHF process into two distinct stages: (post-)training and deployment.
-In the training stage, the learner’s goal is to develop a high-performing model with minimal sample queries and computational resources, where we explore both passive and active data collection strategies.
-In the deployment stage, the trained model is used to select good responses to user inputs while concurrently gathering informative feedback from the user for further refinement.
-Building on this unified view, we revisit the existing theoretical advancements in RLHF and observe that they primarily focus on specific stages and settings of the overall pipeline.
-Consequently, an understanding of the entire pipeline remains lacking.
-Furthermore, while they have proposed various algorithms with theoretical guarantees for different stages of RLHF, the high statistical and computational complexity of their methods presents a significant challenge for practical applications.
-Thus, a natural question arises:
-Can we develop provably efficient algorithms with improved statistical and computational efficiency for the RLHF pipeline?
-In this work, we answer this question affirmatively by proposing novel algorithms for each stage of the RLHF pipeline, improving both statistical and computational efficiency over existing methods.
-Specifically, our contributions are summarized as follows:
-  • We propose a unified framework for the RLHF pipeline from the perspective of contextual bandits.
-We decompose the RLHF process into two stages: (post-)training and deployment, and explore both passive and active data collection strategies during training.
-This perspective allows us to view several prior works as focusing on specific components of the pipeline.
-  • We propose algorithms that are computationally and statistically more efficient than existing methods for each stage of the RLHF pipeline.
-Specifically, we employ online mirror descent with a tailored local norm to replace the standard maximum likelihood estimation.
-Additionally, we design distinct query criteria for different stages.
-Table 1 summarizes the comparison between previous works and ours about statistical and computational efficiency.
-  • We train and deploy Llama-3-8B-Instruct on the Ultrafeedback-binarized dataset using our method, testing both passive and active data collection strategies in the training stage.
-The empirical results validate our theoretical findings and confirm the effectiveness of our method.
+Reinforcement Learning from Human Feedback is a critical technique for training large language models using human preference feedback (Ouyang et al., 2022; Bai et al., 2022).
+Typical RLHF methods involve collecting extensive data, each consisting of a prompt, a pair of responses, and a preference label indicating which response is preferred.
+Then, a reward model is trained to predict the human preference, and the LLM is fine-tuned based on the reward model by the RL algorithms.
+Traditional RLHF methods primarily rely on fixed preference datasets, which typically suffer from limited coverage.
+As a result, the learned reward models struggle to generalize to out-of-distribution (OOD) samples, constraining the effectiveness of the aligned models.
+To address this, online RLHF has emerged as a promising paradigm, enabling iterative data collection and model improvement.
+The general process can be described as (i) collect the preference data; (ii) update the model using the collected preference data.
+The above two steps are repeated for several iterations to boost model performance.
+In practice, the Claude (Bai et al., 2022) and LLaMA-2 (Touvron et al., 2023) projects have demonstrated that online RLHF can significantly enhance model performance.
+Despite its empirical success, online RLHF introduces significant computational challenges.
+Specifically, the typical process of online RLHF involves continuously integrating newly collected data into the dataset and re-optimizing the model from scratch over the expanded dataset.
+While this strategy is statistically efficient, its computational and storage costs scale linearly with the number of iterations, which becomes prohibitive in long-term iterations, especially on edge devices where computation and memory resources are inherently limited.
+This raises a pressing question:
+Can we design online RLHF algorithms that are both statistically and computationally efficient?
+[…]
+Setting Context Action Gap/Regret Time Storage Reference
+[…]
+$\widetilde{\mathcal{O}}\big{(}\sqrt{d}\cdot\lVert\Phi\rVert_{\mathcal{H}_{T}^{-1}}\big{)}$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ Ours (Theorem 1)
+[…]
+$\widetilde{\mathcal{O}}\big{(}d\sqrt{\kappa/T}\big{)}$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ Ours (Theorem 2)
+[…]
+$\widetilde{\mathcal{O}}\big{(}d\sqrt{\kappa T}\big{)}$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ Ours (Theorem 3)
+In this work, we provide an affirmative answer to this question in the setting of contextual preference bandits with linearly parameterized reward functions.
+Specifically, building on recent theoretical advancements in RLHF (Zhu et al., 2023; Das et al., 2024; Ji et al., 2024), we formulate the RLHF problem as a contextual dueling bandit problem (Yue et al., 2012; Saha, 2021).
+While prior work has explored this formulation, most existing methods focus on statistical efficiency and overlook the growing computational burden.
+To bridge this gap, we introduce a novel one-pass reward modeling algorithm which does not require storing the historical data and can be computed in constant time.
+Our method is based on the online mirror descent framework with a tailored local norm that captures second-order information.
+We then apply our method to several online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation.
+We establish theoretical guarantees showing that our method improves both statistical and computational efficiency.
+Table 1 summarizes the comparison of our method with the existing methods.
+To enable usage in LLMs, we develop practical variants of our method.
+Direct computation and storage of the Hessian matrix is prohibitively expensive; thus, we propose an efficient approximation using Hessian-Vector Products (HVP) combined with conjugate gradient descent, avoiding explicit second-order information and relying only on first-order computation.
+Additionally, we employ rejection sampling to approximate model uncertainty in a computationally efficient manner.
+With the above techniques, we conduct experiments using the LLaMA-3-8B-Instruct (Llama Team, 2023) and Qwen2.5-7B-Instruct (Qwen Team, 2024) models on the Ultrafeedback-binarized (Cui et al., 2023) and Mixture2 (Dong et al., 2024) datasets, validating the effectiveness of our method.
+To summarize, our contributions are as follows:
+  • By formulating the RLHF problem as a contextual dueling bandit, we propose a novel one-pass reward modeling algorithm and establish the corresponding estimation error bound.
+Our method is built upon the online mirror descent framework and incorporates a carefully designed local norm that captures second-order information for improved learning efficiency.
+  • We apply our method to a broad range of online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation.
+For each setting, we design tailored algorithms and establish corresponding theoretical guarantees, demonstrating that our approach achieves improved statistical and computational efficiency than existing methods.
+  • We develop practical algorithms by approximating the update using Hessian-Vector Products combined with conjugate gradient descent, and estimating uncertainty via rejection sampling.
+Based on the above techniques, we conduct empirical evaluations using LLaMA-3-8B-Instruct and Qwen2.5-7B-Instruct models on Ultrafeedback-binarized and Mixture2 datasets, showing our method achieves better performance in both statistical and computational efficiency.
 Organization.
-The remainder of this paper is organized as follows.
-We review related work in Section 2 and introduce the problem setup in Section 3.
-Our pipeline is presented in Sections 4 and 5, focusing on the training and deployment stages, respectively.
-We present experiments in Section 6 and conclude in Section 7.
-All proofs and experiment details are deferred to the appendix.
-Notations.
-For a vector $x\in\mathbb{R}^{d}$ and a positive semi-definite matrix $A\in\mathbb{R}^{d\times d}$, we define $\lVert x\rVert_{A}=\sqrt{x^{\top}Ax}$.
-We use the notation $\mathcal{O}(\cdot)$ to highlight various dependencies depending on the context.
-For the statistical efficiency, $\mathcal{O}(\cdot)$ omits only constant factors.
-For computational efficiency, $\mathcal{O}(\cdot)$ solely emphasizes the dependence on the number of samples as this is the primary factor influencing the complexity.
-Additionally, we employ $\widetilde{\mathcal{O}}(\cdot)$ to hide polylogarithmic factors.
+The rest of the paper is organized as follows: Section 3 introduces the problem setup.
+Section 4 presents our proposed one-pass reward modeling method.
+Section 5 applies our method to different online RLHF settings.
+Section 6 provides the practical implementations.
+Section 7 presents experimental results.
+Section 8 concludes the paper.
+We defer the proofs, experiment details, and more experimental results to the appendix.
 2 Related Work
-In this section, we review the works most closely related to ours.
-Reinforcement Learning from Human Feedback.
+In this section, we review the works most closely related to ours, including offline RLHF, online RLHF, contextual dueling bandits, and active learning.
+Offline RLHF.
 Learning from human preferences with deep learning models dates back to Christiano et al.
@@
 Among these works, Direct Preference Optimization (DPO) (Rafailov et al., 2023) is one of the most popular reward-free methods, which treats generative models directly as reward models and optimizes them using human preference feedback.
-In this work, we focus on the reward-based RLHF methods.
 Given the empirical success of RLHF, recent efforts have been devoted to developing a deeper theoretical understanding of this approach.
 Zhu et al.
-(2023) investigated the standard offline setting, where the learner is provided with a fixed dataset and aims to learn a policy that maximizes the expected reward.
+(2023) proposed to formulate RLHF as a contextual preference bandit problem and proved the maximum likelihood estimator converges under both the Bradley-Terry (BT) model and Plackett-Luce (PL) model with linear reward functions.
+Online RLHF.
+Offline RLHF methods rely on fixed preference datasets, which often suffer from limited data coverage.
+Consequently, the resulting reward models struggle to generalize to out-of-distribution samples, thereby limiting the effectiveness of the aligned models.
+To overcome this limitation,online RLHF has emerged as a promising alternative, enabling iterative data collection and continuous model refinement.
+The works (Dong et al., 2023; Guo et al., 2024; Yuan et al., 2024; Wu et al., 2025) have demonstrated that online iterative variants of direct preference learning algorithms significantly outperform their offline counterparts.
 Xiong et al.
-(2024) investigated the iterative setting, where the learner randomly selects prompts from the candidate set, generates response pairs by sampling from the current policy, and then queries for feedback to update the policy accordingly.
-Das et al.
-(2024) studied the active setting where the learner is allowed to select both prompts and responses, query for feedback, and update the policy accordingly.
-Liu et al.
-(2024) focused on the dual active setting where the learner is allowed to select conversations and teachers simultaneously for feedback.
-Essentially, these works focus on specific stages and settings within the RLHF pipeline, while we offer a unified view on the entire process.
-Furthermore, we propose algorithms that are both computationally and statistically more efficient than existing methods for each stage of the RLHF pipeline.
+(2024) identified key challenges in offline RLHF and theoretically demonstrated the potential benefits of online exploration.
+Recent work has incorporated optimism-driven bonus terms into the DPO loss to encourage exploration in online RLHF (Xie et al., 2025; Cen et al., 2025; Zhang et al., 2025).
+These studies primarily focus on the sample efficiency, but do not consider the accompanying increase in computational complexity.
+In this paper, we propose algorithms that achieve both statistical and computational efficiency.
 Contextual Dueling Bandits and RL.
-Dueling bandits is a variant of the multi-armed bandit problem that incorporates pairwise comparison feedback (Yue et al., 2012).
-Contextual dueling bandits further extend this framework by integrating context information (Dudík et al., 2015).
-In this setting, Saha (2021) studied the $K$-arm contextual dueling bandit problem and proposed an algorithm with a near-optimal regret guarantee.
-Saha et al.
-(2023) extended the contextual dueling bandit problem to the reinforcement learning setting, presenting an algorithm with a sublinear regret bound.
-Sekhari et al.
-(2024) further studied the contextual dueling bandit problem with active learning, where the learner is allowed to query for feedback adaptively and the goal is to minimize the regret while minimizing the number of queries simultaneously.
-However, these works primarily focus on the online setting, where the context is provided in an arbitrary and online manner and goal is to minimize cumulative regret.
-In this work, we consider the entire RLHF pipeline, including both training and deployment stages, and explore both passive and active data collection strategies in the training stage.
+Dueling bandits are a variant of the multi-armed bandit problem in which the learner sequentially selects a pair of arms and receives noisy binary feedback (Yue et al., 2012).
+The contextual dueling bandit framework extends this setting by incorporating contextual information (Dudík et al., 2015; Saha, 2021; Bengs et al., 2022).
+Within this framework, Saha (2021) studied the $K$-armed contextual dueling bandit problem, and Saha et al.
+(2023) further extended it to the reinforcement learning setting.
+Additionally, Sekhari et al.
+(2024) investigated the contextual dueling bandit problem under an active learning paradigm, where the learner adaptively queries for feedback, aiming to minimize both regret and the number of queries.
+To move beyond linear reward functions, Verma et al.
+(2025) introduced the neural dueling bandit problem, modeling the reward function using neural networks.
+These prior works commonly rely on maximum likelihood estimation to learn the reward function, leading to computational complexity that grows linearly with the number of iterations.
+In contrast, we propose algorithms that maintain constant per-iteration computational complexity, while preserving statistical efficiency.
 Active Learning.
-Active learning is a well-established field that aims to reduce the labeling cost by selecting the most informative samples for annotation (Settles, 2009).
-The first line of work focuses on the pool-based setting (Seung et al., 1992; Freund et al., 1997; Huang et al., 2010).
-In this setting, the learner strategically selects a batch of the most informative data at each step and queries labels only for this selected batch.
-The learner then uses this labeled batch to update the model.
-Once the model is enhanced, the learner selects another batch of data and queries labels for it, continuing the iterative training process.
-The other line of work focuses on the online active setting (Cesa-Bianchi et al., 2004, 2006).
-In this setting, the learner sequentially observes data points and decides whether to query labels for them.
-Ji et al.
-(2024) explored the online active RLHF setting where prompts are provided in an arbitrary and online manner, with the goal of minimizing cumulative regret while minimizing the number of queries simultaneously.
-In this work, we leverage the pool-based active learning strategy in the training stage of the RLHF pipeline, where the learner is allowed to select the dataset actively.
+Active learning is a paradigm that aims to reduce the labeling cost by selecting the most informative samples for annotation (Settles, 2009).
+In general, existing work can be categorized into two settings: pool-based and stream-based.
+The pool-based setting (Seung et al., 1992; Freund et al., 1997; Huang et al., 2010) involves the learner iteratively selecting a batch of informative samples from a large unlabeled pool, querying their labels, updating the model, and repeating this process.
+In contrast, the stream-based setting (Cesa-Bianchi et al., 2004, 2006; Cacciarelli and Kulahci, 2024) requires the learner to sequentially observe data points and decide in real time whether to query their labels.
+Within the context of RLHF, Das et al.
+(2024) studied pool-based active learning, while Ji et al.
+(2024) focused on the stream-based setting.
+In this work, we focus on the pool-based strategy, which can be naturally extended to the stream-based scenario.
 3 Problem Setup
-In this section, we introduce the problem setup for RLHF.
-Following recent advancements in RLHF (Zhu et al., 2023; Xiong et al., 2024), we formulate RLHF as a contextual bandit problem.
+Following recent advancements in RLHF (Zhu et al., 2023; Das et al., 2024; Xiong et al., 2024), we formulate RLHF as a contextual bandit problem.
 Specifically, we have a set of contexts $\mathcal{X}$ and a set of possible actions $\mathcal{A}$ per context.
 To learn with human preference feedback, the learner selects a tuple $(x,a,a^{\prime})$ to present to the human, where $x\in\mathcal{X}$ is the context, $a,a^{\prime}\in\mathcal{A}$ are the actions.
-The human then provides a binary preference feedback $y\in\{0,1\}$, where $y=1$ indicates the human prefers action $a$ over action $a^{\prime}$, and $y=0$ otherwise.
+The human then provides a binary preference feedback $y\in\{0,1\}$, where $y=1$ indicates that the human prefers action $a$ over action $a^{\prime}$, and $y=0$ otherwise.
 We study the commonly used Bradley-Terry (BT) model in preference learning (Bradley and Terry, 1952), which assumes that the human’s preference is generated by a logistic function of the difference in the rewards of the two actions.
@@
 Given a context $x\in\mathcal{X}$ and two actions $a,a^{\prime}\in\mathcal{A}$, the probability of the human preferring action $a$ over action $a^{\prime}$ is given by
-$\mathbb{P}\left[y=1\mid x,a,a^{\prime}\right]=\frac{\exp\left(r(x,a)\right)}{\exp\left(r(x,a)\right)+\exp\left(r\left(x,a^{\prime}\right)\right)},$
-where $r$ is a latent reward function.
-We consider the linear reward model which is widely used in the RLHF literature (Zhu et al., 2023; Ji et al., 2024; Das et al., 2024).
+$\mathbb{P}\left[y=1\mid x,a,a^{\prime}\right]=\frac{\exp\left(r(x,a)\right)}{\exp\left(r(x,a)\right)+\exp\left(r\left(x,a^{\prime}\right)\right)},$ (1)
+where $r:\mathcal{X}\times\mathcal{A}\to\mathbb{R}$ is a latent reward function.
+We consider the widely used linear reward in the literature (Zhu et al., 2023; Cen et al., 2025).
 Assumption 1.
 It holds that $r(x,a)=\phi(x,a)^{\top}\theta^{*}$ where $\phi(x,a):\mathcal{X}\times\mathcal{A}\to\mathbb{R}^{d}$ is the known and fixed feature map, and $\theta^{*}\in\mathbb{R}^{d}$ is the unknown parameter vector.
-Furthermore, we assume $\lVert\phi(x,a)\rVert_{2}\leq L$ for all $x\in\mathcal{X}$ and $a\in\mathcal{A}$ and $\theta^{*}\in\Theta$ where $\Theta=\{\theta\in\mathbb{R}^{d}\mid\langle\bm{1},\theta\rangle=0,\|\theta\|_{2}\leq B\}$.
+Furthermore, we assume $\lVert\phi(x,a)\rVert_{2}\leq L$ for all $x\in\mathcal{X}$ and $a\in\mathcal{A}$ and $\theta^{*}\in\Theta$ where $\Theta=\{\theta\in\mathbb{R}^{d}\mid\|\theta\|_{2}\leq B\}$.
 Remark 1.
 The feature mapping $\phi$ can be constructed by removing the last layer of a pre-trained large language model, and $\theta^{*}$ corresponds to the weights of the last layer.
-Based on the linear model, we can rewrite the probability as
-$\mathbb{P}\left[y=1\mid x,a,a^{\prime}\right]=\sigma\left(\phi(x,a)^{\top}\theta^{*}-\phi(x,a^{\prime})^{\top}\theta^{*}\right),$ (1)
-where $\sigma(w)=\frac{1}{1+\exp(-w)}$ is the sigmoid function. Next, we introduce a key quantity that captures learning complexity under the BT model.
+Then, we can rewrite the probability as $\mathbb{P}\left[y=1\mid x,a,a^{\prime}\right]=\sigma(\phi(x,a)^{\top}\theta^{*}-\phi(x,a^{\prime})^{\top}\theta^{*})$, where $\sigma(w)=\frac{1}{1+\exp(-w)}$.
+Next, we introduce a key quantity that captures learning complexity.
 Definition 2.
@@
 where $\dot{\sigma}(w)=\sigma(w)(1-\sigma(w))$ is the derivative function.
-Remark 2.
-Through direct calculation, we obtain $\kappa\leq 2+\exp(2BL)+\exp(-2BL)$.
-Thus, $\kappa$ may be exceedingly large with an exponential dependence on the magnitude of features and parameters, as justified in Appendix G.2.
-The RLHF pipeline can generally be divided into two distinct stages: the (post-)training stage and the deployment stage, each with different objectives.
-In the training stage, the learner aims to develop a high-performing model with minimal samples and computational resources.
-In the deployment stage, the trained model is used to generate good responses to user inputs, while simultaneously gathering informative feedback for further refinement.
-We provide a detailed discussion of both stages in the following sections.
-4 Training Stage
-In the training stage, the learner aims to acquire a high-performing model with minimal number of samples and computational resources.
-To effectively evaluate the model’s performance, we define the following performance measure.
-Performance Measure.
-For any policy $\pi$, we define the sub-optimality gap compared to the optimal policy as follows:
-$\operatorname*{\mathsf{SubOpt}}(\pi)=\mathbb{E}_{x\sim\rho}\left[r\left(x,\pi^{*}(x)\right)-r(x,\pi(x)\right],$ (3)
-where $\rho$ is the context distribution, $\pi^{*}=\arg\max_{a}r(x,a)$ is the optimal policy under the true reward $r$.
-The goal of the learner is to minimize the performance gap $\operatorname*{\mathsf{SubOpt}}(\pi)$.
-In the training stage, two data collection strategies are commonly used: passive and active.
-In passive setting, the learner is given a fixed set of samples.
-Conversely, in active setting, the learner can actively select queries for human feedback.
-We discuss both strategies below.
-4.1 Passive Data Collection
-In the passive data collection setting, we are given a fixed dataset $\left\{\left(x_{t},a_{t},a_{t}^{\prime},y_{t}\right)\right\}_{t=1}^{T}$ where $x_{t}\in\mathcal{X}$ is the context, $a_{t},a_{t}^{\prime}\in\mathcal{A}$ are the actions, and $y_{t}\in\{0,1\}$ is the feedback.
-Since the parameter $\theta^{*}$ is unknown, we need to estimate it from the available data.
-The most common method for this estimation is maximum likelihood estimation (MLE).
-First, let $z=\phi(x,a)-\phi(x,a^{\prime})$ be the feature difference, then we define the negative log-likelihood $\ell_{t}(\theta)$ for sample $t$ as
-$\ell_{t}(\theta)=-y_{t}\log(\sigma(z_{t}^{\top}\theta))-\left(1-y_{t}\right)\log(1-\sigma(z_{t}^{\top}\theta)).$ (4)
-The MLE estimator $\widehat{\theta}_{T+1}$ is then defined as
-$\widehat{\theta}_{T+1}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\mathcal{L}_{T+1}(\theta)\triangleq\sum_{t=1}^{T}\ell_{t}(\theta).$ (5)
-Several works (Zhu et al., 2023; Ji et al., 2024) have analyzed the statistical properties of MLE and showed that for any $\delta\in(0,1)$, with probability at least $1-\delta$, the MLE estimator $\widehat{\theta}_{T+1}$ converges to the true parameter $\theta^{*}$ by
-$\lVert\widehat{\theta}_{T+1}-\theta^{*}\rVert_{V_{T+1}}\leq\widehat{\beta}\triangleq\widetilde{\mathcal{O}}\big{(}\kappa\sqrt{d}\big{)},$ (6)
-where $V_{t}\triangleq\sum_{i=1}^{t-1}z_{i}z_{i}^{\top}+\lambda I$ is the feature covariance matrix, $\lambda$ is the regularization parameter.
-For computational complexity, the MLE estimator does not have a closed-form solution, requiring iterative optimization techniques, such as gradient descent, to obtain an $\varepsilon$-accurate solution.
-As discussed in Faury et al.
-(2022), optimizing the MLE typically requires $\mathcal{O}(\log(1/\varepsilon))$ iterations to reach the desired accuracy.
-Since the loss function is defined over all data samples, each iteration involves $\mathcal{O}(T)$ gradient computations.
-Given that $\varepsilon$ is typically set to $1/T$ to prevent it from dominating the overall estimation error, the total computational complexity is $\mathcal{O}(T\log T)$.
-Although the MLE estimator is a commonly used approach, we propose a novel algorithm that enhances both computational efficiency and sample efficiency compared to MLE, drawing inspiration from recent advancements in logistic bandits (Faury et al., 2022; Zhang and Sugiyama, 2023) and multinomial logit MDPs (Li et al., 2024).
-First, we define the gradient $g_{t}(\theta)$ and Hessian $H_{t}(\theta)$ of the loss function $\ell_{t}(\theta)$ as
-$g_{t}(\theta)=\left(\sigma\left(z_{t}^{\top}\theta\right)-y_{t}\right)z_{t},~{}~{}H_{t}(\theta)=\dot{\sigma}\left(z_{t}^{\top}\theta\right)z_{t}z_{t}^{\top}.$ (7)
+The quantity $\kappa$ captures the learning complexity that satisfies $\kappa\leq 2+\exp(2BL)+\exp(-2BL)$.
+It may be exceedingly large with exponential dependence on the scale of features and parameters.
+4 Our Framework
+In this section, we first introduce the general framework for online RLHF. We then present our one-pass reward modeling method. Finally, we show the theoretical guarantee of our method.
+4.1 General Framework for Online RLHF
+The general process of online RLHF involves iteratively collecting data and updating the model based on the collected data. At iteration $t$, the process can be formulated as:
+  1. (i)
+    New data collection: Sample a prompt $x_{t}$ and two responses $a_{t}$ and $a_{t}^{\prime}$, query the oracle to obtain the preference label $y_{t}\in\{0,1\}$, expand the dataset $\mathcal{D}_{t+1}=\mathcal{D}_{t}\cup\{(x_{t},a_{t},a_{t}^{\prime},y_{t})\}$.
+  2. (ii)
+    Reward modeling: Train a reward model $r_{t+1}$ using the historical dataset $\mathcal{D}_{t+1}$.
+  3. (iii)
+    Policy optimization (Optional): Update the policy $\pi_{t+1}$ using the reward model $r_{t+1}$.
+A key challenge in online RLHF is that the reward model needs to be trained on the entire historical dataset at each iteration, which is computationally expensive.
+Specifically, let $z_{t}=\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})$ be the feature difference, given the historical dataset $\mathcal{D}_{t+1}=\left\{(x_{i},a_{i},a_{i}^{\prime},y_{i})\right\}_{i=1}^{t}$, the reward model is estimated via maximum likelihood estimation as
+$\widehat{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum_{i=1}^{t}\ell_{i}(\theta),\text{where }\ell_{t}(\theta)=-y_{t}\log(\sigma(z_{t}^{\top}\theta))-\left(1-y_{t}\right)\log(1-\sigma(z_{t}^{\top}\theta)).$ (3)
+However, Eq.
+(3) does not admit a closed-form solution, requiring iterative optimization techniques, such as gradient descent, to achieve an $\varepsilon$-accurate estimate.
+As discussed by Faury et al.
+(2022), obtaining such accuracy with MLE typically requires $\mathcal{O}(\log(1/\varepsilon))$ optimization steps.
+Since the loss function is defined over the entire historical dataset, each iteration incurs a computational cost of $\mathcal{O}(t)$ gradient evaluations.
+In practice, $\varepsilon$ is often set to $1/t$ to ensure that the optimization error does not dominate the overall estimation error.
+As a result, the total computational complexity at iteration $t$ becomes $\mathcal{O}(t\log t)$, a cost that is prohibitive for long-term online RLHF applications.
+4.2 One-pass Reward Modeling
+Drawing inspiration from recent advancements in logistic bandits (Faury et al., 2022; Zhang and Sugiyama, 2023) and multinomial logit MDPs (Li et al., 2024), we propose a novel one-pass reward modeling method that reduces the complexity to constant time per iteration.
+First, define the gradient $g_{t}(\theta)$ and Hessian $H_{t}(\theta)$ of loss $\ell_{t}(\theta)$ as $g_{t}(\theta)=(\sigma(z_{t}^{\top}\theta)-y_{t})z_{t}$ and $H_{t}(\theta)=\dot{\sigma}(z_{t}^{\top}\theta)z_{t}z_{t}^{\top}$.
 Implicit OMD.
@@
 where ¯$\bar{\mathcal{H}}_{t}=\sum_{i=1}^{t-1}H_{i}\left(\bar{\theta}_{i+1}\right)+\lambda I$ is the local norm, and $\eta$ is the step size.
-The optimization problem can be decomposed in two terms.
+The optimization problem can be decomposed into two terms.
 The first term is the instantaneous log-loss $\ell_{t}(\theta)$, which accounts for the information of the current sample.
@@
 Such a Hessian matrix effectively captures local information and is crucial for ensuring statistical efficiency.
-Although this formulation may appear computationally more efficient than MLE due to its one-pass data processing, the optimization problem in Eq.
-(8) still lacks a closed-form solution.
-Consequently, it still requires $\mathcal{O}(\log(1/\varepsilon))$ iterations per sample to achieve an $\varepsilon$-accurate solution, leading to the same $\mathcal{O}(T\log T)$ computational complexity as MLE.
+The update rule in Eq.
+(4) benefits from a one-pass data processing property, which eliminates the need to store the entire historical dataset.
+However, the optimization problem in Eq.
+(4) still does not have a closed-form solution.
+But since the loss is defined only on the current sample, it requires only $\mathcal{O}(1)$ gradient computations per step, leading to a total computational complexity of $\mathcal{O}(\log t)$ at iteration $t$.
+This represents a significant improvement over the $\mathcal{O}(t\log t)$ complexity of the MLE estimator in Eq.
+(3).
+Nevertheless, the computational complexity of the implicit OMD is still increasing with the number of iterations, which motivates us to design a constant-time method.
 Standard OMD.
-To enhance computational efficiency, we revisit the formulation in Eq.
-(8) and observe that it essentially corresponds to the implicit OMD framework (Kivinen and Warmuth, 1997).
-Consequently, a natural alternative is to replace this formulation with the standard OMD framework, which permits a closed-form solution and thus eliminates the need for iterative optimization.
+To enhance computational efficiency, a natural alternative is to replace this formulation with the standard OMD framework, which permits a closed-form solution and thus eliminates the need for iterative optimization.
 However, the standard OMD minimizes a first-order approximation of the loss function, which sacrifices several key properties compared to its implicit counterpart, as demonstrated by Campolongo and Orabona (2020).
-Specifically, the standard OMD formulation updates using $\partial\ell_{t}(\theta_{t})$, whereas the implicit OMD updates the algorithm approximately with the subsequent sub-gradient, $\partial\ell_{t}(\theta_{t+1})$.
+Specifically, the standard OMD formulation updates using $g_{t}(\theta_{t})$, whereas the implicit OMD updates the algorithm approximately with the subsequent sub-gradient, $g_{t}(\theta_{t+1})$.
 This distinction results in a notable gap in the convergence rates of the two methods.
 To this end, we propose to approximate the current loss $\ell_{t}(\theta)$ using a second-order Taylor expansion, drawing inspiration from Zhang and Sugiyama (2023).
-We define the second-order approximation of the original loss function $\ell_{t}(\theta)$ at the current estimator $\widetilde{\theta}_{t}$ as ~$\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+g_{t}(\widetilde{\theta}_{t})^{\top}(\theta-\widetilde{\theta}_{t})+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}$.
-Then, we replace the current loss $\ell_{t}(\theta)$ in Eq.
-(8) with the approximation ~$\widetilde{\ell}_{t}(\theta)$, leading to the following update rule:
+Define the second-order approximation of $\ell_{t}(\theta)$ as ~$\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+g_{t}(\widetilde{\theta}_{t})^{\top}(\theta-\widetilde{\theta}_{t})+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}$.
+Then, we replace the loss $\ell_{t}(\theta)$ in Eq.
+(4) with the approximation ~$\widetilde{\ell}_{t}(\theta)$, leading to the update rule:
 $\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\big{\langle}g_{t}(\widetilde{\theta}_{t}),\theta\big{\rangle}+\frac{1}{2\eta}\big{\|}\theta-\widetilde{\theta}_{t}\big{\|}_{\widetilde{\mathcal{H}}_{t}}^{2}\Big{\}},$ (5)
-where $\eta$ is the step size and ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$ is the local norm with $\mathcal{H}_{t}\triangleq\mathcal{H}_{t}(\widetilde{\theta}_{i+1})=\sum_{i=1}^{t-1}H_{i}(\widetilde{\theta}_{i+1})+\lambda I$.
-The optimization problem in Eq. (9) can be solved with a projected gradient step with the following equivalent formulation:
-$\widetilde{\theta}_{t+1}^{\prime}=\widetilde{\theta}_{t}-\eta\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t}),\quad\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\widetilde{\theta}_{t+1}^{\prime}\rVert_{\widetilde{\mathcal{H}}_{t}}^{2}.$
-Thus, the estimator $\widetilde{\theta}_{t+1}$ can be computed in an $\mathcal{O}(1)$ time complexity, leading to a total time complexity of $\mathcal{O}(T)$ for the entire procedure.
-This represents a significant improvement over the $\mathcal{O}(T\log T)$ complexity of the MLE estimator in Eq.
-(5), particularly considering that the number of samples $T$ is typically large in the training of LLMs.
-Furthermore, since the estimator processes the samples in a one-pass manner, it mitigates the memory burden associated with computing the gradient of the full dataset at each iteration.
-This makes it more suitable for large-scale applications, where memory and time efficiency are crucial.
-It is important to note that the update rule in Eq. (9) represents a special case of online mirror descent, specifically, it can be derived from the following general form:
+where $\eta$ is the step size and ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$ is the local norm with $\mathcal{H}_{t}\triangleq\sum_{i=1}^{t-1}H_{i}(\widetilde{\theta}_{i+1})+\lambda I$.
+Eq.
+(5) can be solved with a projected gradient step with the following equivalent form:
+$\widetilde{\theta}_{t+1}^{\prime}=\widetilde{\theta}_{t}-\eta\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t}),~{}~{}~{}\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\widetilde{\theta}_{t+1}^{\prime}\rVert_{\widetilde{\mathcal{H}}_{t}}^{2}.$
+Thus, the estimator $\widetilde{\theta}_{t+1}$ provides a closed-form solution, leading to a $\mathcal{O}(1)$ computational complexity per iteration.
+Since the estimator processes the samples in a one-pass manner, it mitigates the memory burden associated with computing the gradient of the full dataset.
+These properties make the method particularly suitable for edge devices, where both memory and computational resources are severely constrained.
+The detailed process of our proposed method is presented in Algorithm 1.
+Algorithm 1 One-Pass Reward Modeling
+0:  Preference data $(x_{t},a_{t},a_{t}^{\prime},y_{t})$
+1:  Define the loss function $\ell_{t}(\theta)$ as Eq. (3)
+2:  Update ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$
+3:  Compute $\widetilde{\theta}_{t+1}^{\prime}=\widetilde{\theta}_{t}-\eta\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t})$
+4:  Compute $\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\widetilde{\theta}_{t+1}^{\prime}\rVert_{\widetilde{\mathcal{H}}_{t}}^{2}$
+5:  Update $\mathcal{H}_{t+1}=\mathcal{H}_{t}+H_{t}(\widetilde{\theta}_{t+1})$
+5:  $\widetilde{\theta}_{t+1}$
+Algorithm 2 Passive Data Collection
+0:  Regularization $\lambda$, step size $\eta$
+1:  Initialize $\widetilde{\theta}_{1}=\bm{0}$ and ~$\widetilde{\mathcal{H}}_{1}=\lambda I$
+2:  for $t=1,2,\ldots,T$ do
+3:     Observe preference data $(x_{t},a_{t},a_{t}^{\prime},y_{t})$
+[…]
+5:  end for
+6:  Construct $\widetilde{J}_{T+1}(\pi)$ as in Eq. (6)
+6:  $\pi_{T+1}=\operatorname*{arg\,max}_{\pi\in\Pi}\widetilde{J}_{T+1}(\pi)$
+4.3 Theoretical Guarantee
+Note that the update rule in Eq. (5) is a special case of online mirror descent, specifically:
 $\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\big{\langle}g_{t}(\widetilde{\theta}_{t}),\theta\big{\rangle}+\frac{1}{\eta}\mathcal{D}_{\psi_{t}}(\theta,\widetilde{\theta}_{t})\Big{\}},$
-where $\psi_{t}(\theta)=\frac{1}{2}\lVert\theta\rVert_{\widetilde{\mathcal{H}}_{t}}^{2}$ is the regularizer and $\mathcal{D}_{\psi_{t}}(\theta,\widetilde{\theta}_{t})=\psi_{t}(\theta)-\psi_{t}(\widetilde{\theta}_{t})-\langle\nabla\psi_{t}(\widetilde{\theta}_{t}),\theta-\widetilde{\theta}_{t}\rangle$ is the Bregman divergence.
-Therefore, we can analyze the estimation error using the online mirror descent framework (Orabona, 2019).
-Specifically, we have the following lemma.
-Algorithm 1 Training Stage: Passive Data Collection
-0:  Dataset $\left\{\left(x_{t},a_{t},a_{t}^{\prime},y_{t}\right)\right\}_{t=1}^{T}$, regularization $\lambda>0$
-1:  Initialize $\widetilde{\theta}_{1}=0$, $\mathcal{H}_{1}=\lambda I_{d}$.
+where $\psi_{t}(\theta)=\frac{1}{2}\lVert\theta\rVert_{\widetilde{\mathcal{H}}_{t}}^{2}$ is the regularizer and $\mathcal{D}_{\psi_{t}}(\theta,\widetilde{\theta}_{t})=\psi_{t}(\theta)-\psi_{t}(\widetilde{\theta}_{t})-\langle\nabla\psi_{t}(\widetilde{\theta}_{t}),\theta-\widetilde{\theta}_{t}\rangle$ is Bregman divergence.
+Based on the analysis of online mirror descent, we have the following lemma.
+Lemma 1.
+Let $\delta\in(0,1]$, set $\eta=(1/2)\log 2+\left(BL+1\right)$ and $\lambda=84\sqrt{2}\eta(dL^{2}+BL^{3})$, define $\mathcal{C}_{t}=\{\theta\in\Theta\mid\|\theta-\widetilde{\theta}_{t}\|_{\mathcal{H}_{t}}\leq\widetilde{\beta}_{t}\triangleq\mathcal{O}\big{(}\sqrt{d}(\log(t/\delta))^{2}\big{)}\}$.
+Then, we have $\operatorname{Pr}\left[\forall t\geqslant 1,\theta^{*}\in\mathcal{C}_{t}\right]\geqslant 1-\delta$.
+Comparison with MLE.
+For the MLE estimator in Eq.
+(3), prior works (Zhu et al., 2023; Das et al., 2024; Ji et al., 2024) have shown $\|\theta-\widetilde{\theta}_{t}\|_{V_{t}}\leq\widetilde{\mathcal{O}}(\kappa\sqrt{d})$, where $V_{t}=\sum_{i=1}^{t-1}z_{i}z_{i}^{\top}+\lambda I$.
+By the definition of $\mathcal{H}_{t}$, it holds that $\mathcal{H}_{t}\succeq\kappa^{-1}V_{t}$, Lemma 1 implies $\|\theta-\widetilde{\theta}_{t}\|_{V_{t}}\leq\sqrt{\kappa}\|\theta-\widetilde{\theta}_{t}\|_{\mathcal{H}_{t}}\leq\widetilde{\mathcal{O}}(\sqrt{\kappa d})$.
+This result shows that Lemma 1 improves upon previous bounds by at least a factor of $\sqrt{\kappa}$.
+5 Applications in Three Online RLHF Scenarios
+In this section, we apply our framework to three distinct RLHF scenarios, including online RLHF with passive data collection, active data collection, and deployment-time adaptation.
+(a) Passive Data Collection
+(b) Active Data Collection
+(c) Deployment-Time Adaptation
+[…]
+5.1 Online RLHF with Passive Data Collection
+We first consider the passive data collection setting, where the algorithm can not control the data collection process.
+At each iteration, the learner obtains $(x_{t},a_{t},a_{t}^{\prime},y_{t})$ and updates by Eq.
+(5).
+We adopt the “pessimism in the face of uncertainty” principle and define the value function $\widetilde{J}_{t+1}(\pi)$ as
+$\widetilde{J}_{T+1}(\pi)=(\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi(x))\right])^{\top}\widetilde{\theta}_{T+1}-\widetilde{\beta}_{T+1}\lVert\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi(x))\right]\rVert_{\mathcal{H}_{T+1}^{-1}}.$ (6)
+where $\rho$ is the context distribution.
+The policy $\pi_{T+1}$ is selected as $\pi_{T+1}=\operatorname*{arg\,max}_{\pi\in\Pi}\widetilde{J}_{T+1}(\pi)$.
+The detailed procedure is present in Algorithm 2, and we show it enjoys the following guarantee:
+Theorem 1.
+Set parameters as in Lemma 1, with probability at least $1-\delta$, Algorithm 2 ensures
+${\operatorname*{\mathsf{SubOpt}}}(\pi_{T+1})=\mathbb{E}_{x\sim\rho}\left[r(x,\pi^{*}(x))-r(x,\pi_{T+1}(x))\right]\leq\widetilde{\mathcal{O}}\left(\sqrt{d}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\big{\rVert}_{\mathcal{H}_{T+1}^{-1}}\right),$
+where $\rho$ is the context distribution and $\pi^{*}$ is the optimal policy.
+Remark 2.
+The term $\big{\lVert}\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\big{\rVert}_{\mathcal{H}_{T+1}^{-1}}$ is referred to “concentrability coefficient” in the literature.
+It measures the distribution shift between the optimal policy and the collected data.
+Remark 3.
+Our result improves upon Zhu et al.
+(2023) in both statistical and computational efficiency.
+Specifically, since $\mathcal{H}_{t}\succeq\kappa^{-1}V_{t}$, Theorem 2 directly implies a $\widetilde{\mathcal{O}}(\sqrt{d\kappa}\cdot\lVert\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\rVert_{V_{T+1}^{-1}})$ guarantee, improving their $\widetilde{\mathcal{O}}(\sqrt{d}\kappa\cdot\lVert\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\rVert_{V_{T+1}^{-1}})$ result by a factor of $\sqrt{\kappa}$.
+Regarding computational efficiency, their algorithm has a total storage complexity of $\mathcal{O}(T)$ and a time complexity of $\mathcal{O}(T\log T)$, leading to an amortized per-iteration cost of $\mathcal{O}(\log T)$.
+In contrast, our algorithm maintains a strict $\mathcal{O}(1)$ complexity per iteration, offering a substantial computational advantage.
+Algorithm 3 Active Data Collection
+0:  Regularization $\lambda$, step size $\eta$
+1:  Initialize $\widetilde{\theta}_{1}=\bm{0}$ and $\mathcal{H}_{1}=\lambda I$
 2:  for $t=1,2,\ldots,T$ do
-3:     Define $g_{t}(\theta)$ and $H_{t}(\theta)$ as in Eq. (7)
-4:     Update ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$
-5:     Compute $\widetilde{\theta}_{t+1}$ by Eq. (9)
-6:     Update $\mathcal{H}_{t+1}=\mathcal{H}_{t}+H_{t}(\widetilde{\theta}_{t+1})$
+3:     Choose $(x_{t},a_{t},a_{t}^{\prime})$ as Eq. (7), observe $y_{t}$
+[…]
+5:  end for
+6:  Set $\widetilde{r}_{T+1}(x,a)=\frac{1}{T+1}\sum_{t=1}^{T+1}\phi(x,a)^{\top}\widetilde{\theta}_{t}$
+6:  $\pi_{T+1}(x)=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{r}_{T+1}(x,a)$
+Algorithm 4 Deployment-Time Adaptation
+0:  Regularization $\lambda$, step size $\eta$
+1:  Initialize $\widetilde{\theta}_{1}=0$ and $\mathcal{H}_{1}=\lambda I$.
+2:  for $t=1,2,\ldots,T$ do
+3:     Observes the context $x_{t}$.
+4:     Selects $a_{t}$ and $a_{t}^{\prime}$ as Eq. (9) and Eq. (10)
+5:     Observe the preference feedback $y_{t}$
+[…]
 7:  end for
-8:  Compute $\widetilde{J}_{T}(\pi)$ by Eq. (11)
-8:  Compute final policy $\pi_{T}=\operatorname*{arg\,max}_{\pi}\widetilde{J}_{T}(\pi)$
-Lemma 1.
-Let $\delta\in(0,1]$, set $\eta=(1/2)\log 2+\left(BL+1\right)$ and $\lambda=84\sqrt{2}\eta(dL^{2}+BL^{3})$, define the confidence set as
-$\mathcal{C}_{t}=\left\{\theta\in\Theta\mid\big{\|}\theta-\widetilde{\theta}_{t}\big{\|}_{\mathcal{H}_{t}}\leq\widetilde{\beta}_{t}\triangleq\mathcal{O}\big{(}\sqrt{d}(\log(t/\delta))^{2}\big{)}\right\}.$
-Then, we have $\operatorname{Pr}\left[\forall t\geqslant 1,\theta^{*}\in\mathcal{C}_{t}(\delta)\right]\geqslant 1-\delta$.
-Intuitively, the confidence set shrinks as the number of samples $t$ increases, owing to the growing local norm $\mathcal{H}_{t}$.
-The local norm $\mathcal{H}_{t}$ captures the directions explored by the samples.
-The more we explore in a particular direction, the more accurate our estimation becomes along that direction.
-By the definition of $\kappa$ in Definition 2, it holds that $\mathcal{H}_{t}\succeq(1/\kappa)V_{t}$, Lemma 1 implies $\forall t\in[T+1]$,
-$\lVert\widetilde{\theta}_{t}-\theta^{*}\rVert_{V_{t}}\leq\sqrt{\kappa}\lVert\widetilde{\theta}_{t}-\theta^{*}\rVert_{\mathcal{H}_{t}}\leq\widetilde{\mathcal{O}}\big{(}\sqrt{\kappa d}\big{)}.$ (10)
-Compared with the confidence set in Eq.
-(6), our confidence set strictly improves the estimation error by a factor of $\sqrt{\kappa}$.
-As highlighted in Remark 2, the non-linearity coefficient $\kappa$ may be exceedingly large during the training of LLMs, which makes our estimator statistically more efficient.
-Pessimistic Policy.
-Based on the estimator in Eq. (9), we adopt the “pessimism in the face of uncertainty” principle and define the pessimistic value function as
-$\widetilde{J}_{T}(\pi)=\min_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[(\phi(x,\pi(x))-v)^{\top}\theta\right]=(\bar{\phi}(\pi)-v)^{\top}\widetilde{\theta}_{T}-\widetilde{\beta}_{T}\lVert\bar{\phi}(\pi)-v\rVert_{\mathcal{H}_{T}^{-1}},$ (11)
-where $\bar{\phi}(\pi)=\mathbb{E}_{x\sim\rho}[\phi(x,\pi(x))]$ is the feature expectation of policy $\pi$, and $v$ is the baseline feature vector.
-While $v$ does not affect the induced policy, it may help reduce the sub-optimality gap (Zhu et al., 2023).
-The final policy is given by $\pi=\operatorname*{arg\,min}_{\pi}\widetilde{J}(\pi)$.
-The procedure is present in Algorithm 1.
-We show it enjoys the following guarantee.
-Theorem 1.
-Let $\delta\in(0,1]$, set the parameters as in Lemma 1, with probability at least $1-\delta$, Algorithm 1 ensures the following sub-optimality gap:
-${\operatorname*{\mathsf{SubOpt}}}(\pi_{T})\leq\widetilde{\mathcal{O}}\left(\sqrt{d}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}\right).$
-Remark 3.
-The term $\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}$ is referred to as the “concentratability coefficient”, a widely used concept in offline learning (Jin et al., 2021; Zhu et al., 2023).
-It measures the distribution shift between the optimal policy and the dataset distribution.
-A smaller concentratability coefficient value indicates that the dataset better covers the data distribution under the optimal policy, resulting in a smaller sub-optimality gap.
-Comparison with Zhu et al. (2023).
-Our result improves upon their work in both statistical and computational efficiency.
-First, similar to Eq.
-(10), Theorem 1 implies an $\widetilde{\mathcal{O}}\big{(}\sqrt{\kappa d}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{V_{T}^{-1}}\big{)}$ sub-optimality gap, improving upon theirs by a $\sqrt{\kappa}$ factor.
-Moreover, our estimator has an $\mathcal{O}(T)$ time complexity, more efficient than their $\mathcal{O}(T\log T)$ complexity.
-4.2 Active Data Collection
-In this part, we consider the active data collection setting, where the learner can actively select tuple $\{x_{t},a_{t},a^{\prime}_{t}\}_{t=1}^{T}$ to present to the human for feedback $\{y_{t}\}_{t=1}^{T}$.
-Estimator Construction.
-For each round $t\in[T]$, given $\left\{\left(x_{i},a_{i},a_{i}^{\prime},y_{i}\right)\right\}_{i=1}^{t-1}$, we define the loss function $\ell_{t}(\theta)$ as in Eq.
-(4), then the most standard approach is to estimate the parameter $\theta^{*}$ by MLE, which is given by $\widehat{\theta}_{t}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum_{i=1}^{t-1}\ell_{i}(\theta)$.
-However, as noted before, the MLE estimator leads to a time complexity of $\mathcal{O}(t\log t)$ per round.
-Moreover, unlike the passive setting where the estimator is constructed only once, in the active setting, we need to construct the estimator in each round.
-This is prohibitively expensive for large-scale applications.
-To improve the computational efficiency, we can leverage the OMD framework as in the passive data collection setting. Specifically, we construct the estimator $\widetilde{\theta}_{t}$ as in Eq. (9) by
-$\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\big{\langle}g_{t}(\widetilde{\theta}_{t}),\theta\big{\rangle}+\frac{1}{2\eta}\big{\|}\theta-\widetilde{\theta}_{t}\big{\|}_{\widetilde{\mathcal{H}}_{t}}^{2}\Big{\}}.$ (12)
-Algorithm 2 Training Stage: Active Data Collection
-0:  Regularization $\lambda>0$, failure probability $\delta\in(0,1]$
-1:  Initialize $\widetilde{\theta}_{1}=0$, $\mathcal{H}_{1}=\lambda I_{d}$.
-2:  for $t=1,2,\ldots,T$ do
-3:     Choose $(x_{t},a_{t},a_{t}^{\prime})$ by Eq. (13), observe feedback $y_{t}$
-4:     Define $g_{t}(\theta)$ and $H_{t}(\theta)$ as in Eq. (7)
-5:     Update ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$
-6:     Compute $\widetilde{\theta}_{t+1}$ by Eq. (12)
-7:     Update $\mathcal{H}_{t+1}=\mathcal{H}_{t}+H_{t}(\widetilde{\theta}_{t+1})$
-8:  end for
-9:  Construct $\widetilde{r}_{T}(x,a)=\frac{1}{T}\sum_{t=1}^{T}\phi(x,a)^{\top}\widetilde{\theta}_{t},\forall x,a$
-9:  Compute policy $\pi_{T}(x)=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{r}_{T}(x,a)$.
-Query Selection.
-Given the estimator $\widetilde{\theta}_{t}$, we choose the query with the maximum uncertainty inspired by (Das et al., 2024). Specifically, we choose the tuple by
-$(x_{t},a_{t},a_{t}^{\prime})=\operatorname*{arg\,max}_{x,a,a^{\prime}\in\mathcal{X}\times\mathcal{A}\times\mathcal{A}}\left\{\big{\lVert}\phi(x,a)-\phi(x,a^{\prime})\big{\rVert}_{\mathcal{H}_{t}^{-1}}\right\}.$ (13)
-Intuitively, $\mathcal{H}_{t}$ captures the directions explored so far, the selection criterion essentially seeks the tuple whose feature difference is poorly covered by the past observations.
-Policy Output.
-After $T$ rounds, we define the reward function as $\widetilde{r}_{T}(x,a)=\frac{1}{T}\sum_{t=1}^{T}\phi(x,a)^{\top}\widetilde{\theta}_{t}$ as the average of all the past estimations.
-Then, the final policy is given by
-$\pi_{T}(x)=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{r}_{T}(x,a).$ (14)
+5.2 Online RLHF with Active Data Collection
+As established in Theorem 1, the sub-optimality gap depends on the concentrability coefficient, which quantifies the distributional mismatch between the optimal policy and the collected data.
+In this subsection, we propose an active data collection method that removes this dependency.
+Active Data Collection.
+At each iteration, we obtain a new preference data $(x_{t},a_{t},a_{t}^{\prime},y_{t})$.
+We then update the reward model using our one-pass reward modeling method as in Eq.
+(5).
+To guide data acquisition, we actively select the query with the highest uncertainty under the current reward model.
+Specifically, we select the next query by solving:
+$(x_{t+1},a_{t+1},a_{t+1}^{\prime})=\operatorname*{arg\,max}_{x,a,a^{\prime}\in\mathcal{X}\times\mathcal{A}\times\mathcal{A}}\left\{\big{\lVert}\phi(x,a)-\phi(x,a^{\prime})\big{\rVert}_{\mathcal{H}_{t+1}^{-1}}\right\}.$ (7)
+Policy Optimization.
+After $T$ rounds, we define the reward as the average of all the past estimations $\widetilde{r}_{T+1}(x,a)=\frac{1}{T+1}\sum_{t=1}^{T+1}\phi(x,a)^{\top}\widetilde{\theta}_{t}$.
+The policy is given by $\pi_{T+1}(x)=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{r}_{T+1}(x,a)$.
 The detailed procedure is present in Algorithm 3. We show it enjoys the following guarantee.
 Theorem 2.
-Let $\delta\in(0,1]$, set the parameters as in Lemma 1, with probability at least $1-\delta$, Algorithm 2 ensures the following sub-optimality gap:
-$\operatorname*{\mathsf{SubOpt}}(\pi_{T})\leq\widetilde{\mathcal{O}}\big{(}d\sqrt{{\kappa}/{T}}\big{)}.$
+Set parameters as in Lemma 1, with probability at least $1-\delta$, Algorithm 3 ensures
+$\operatorname*{\mathsf{SubOpt}}(\pi_{T+1})=\mathbb{E}_{x\sim\rho}\left[r(x,\pi^{*}(x))-r(x,\pi_{T+1}(x))\right]\leq\widetilde{\mathcal{O}}\big{(}d\sqrt{{\kappa}/{T}}\big{)},$
+where $\rho$ is the context distribution and $\pi^{*}$ is the optimal policy.
 Remark 4.
-In contrast to the passive data collection setting, the sub-optimality gap in Theorem 2 does not depend on the concentratability coefficient term.
-Instead, it decays at a rate of $T^{-1/2}$.
-This improvement arises because the active setting enables our algorithm to actively select queries with the maximum uncertainty, ensuring the uncertainty across all directions is effectively controlled.
+We attain the same sub-optimality gap as  Das et al.
+(2024), but improve the computational efficiency significantly.
+Our algorithm has an $\mathcal{O}(1)$ time and space complexity per round, while their MLE estimator needs $\mathcal{O}(t\log t)$ time and $\mathcal{O}(t)$ space complexity at iteration $t$.
+5.3 Online RLHF with Deployment-Time Adaptation
+In this section, we consider the deployment-time adaptation setting, where users provide input contexts in an online manner, and the learner generates responses while simultaneously collecting feedback to improve the model.
+In this scenario, the learner faces a dual objective: selecting actions that maximize rewards to ensure a positive user experience, while also choosing actions that yield informative feedback to facilitate continual model improvement.
+To this end, we choose the cumulative regret as the performance measure, defined as
+${\operatorname*{Reg}}_{T}=\sum_{t=1}^{T}\left(r\left(x_{t},\pi^{*}(x_{t})\right)-\frac{1}{2}\left(r\left(x_{t},a_{t}\right)+r\left(x_{t},a^{\prime}_{t}\right)\right)\right),$ (8)
+where $\pi^{*}$ is the optimal policy.
+This measure differs from the one introduced by Ji et al.
+(2024), which evaluates the gap between th e optimal action and a single selected action.
+Our measure is designed to ensure that both actions are sufficiently effective, thereby delivering a high-quality user experience.
+While better aligns with deployment needs, it also poses greater challenges.
+Specifically, to optimize the measure in Ji et al.
+(2024), the learner can exploit one high-reward action while using the other for uniform exploration.
+In contrast, our setting imposes a stricter requirement: both actions must be good, thereby limiting the ability to sacrifice one action purely for exploration.
+Action selection.
+At each iteration, given a new preference data $(x_{t},a_{t},a_{t}^{\prime},y_{t})$, the learner updates the reward model using our one-pass reward modeling method as in Eq.
+(5).
+The learner must select queries that are both informative and with high rewards.
+To address this, we choose the first action $a_{t+1}$ to maximize the estimated reward based on the estimator $\widetilde{\theta}_{t+1}$, i.e.,
+$a_{t+1}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\phi(x_{t+1},a)^{\top}\widetilde{\theta}_{t+1}.$ (9)
+The second action $a_{t+1}^{\prime}$ aims to maximize the reward and the distance between the two actions, i.e.,
+$a_{t+1}^{\prime}=\operatorname*{arg\,max}_{a^{\prime}\in\mathcal{A}}~{}\big{\{}\phi(x_{t+1},a^{\prime})^{\top}\widetilde{\theta}_{t+1}+\widetilde{\beta}_{t+1}\lVert\phi(x_{t+1},a^{\prime})-\phi(x_{t+1},a_{t+1})\rVert_{\mathcal{H}_{t+1}^{-1}}\big{\}}.$ (10)
+The overall algorithm is summarized in Algorithm 4. We show it enjoys the following regret bound.
+Theorem 3.
+For any $\delta\in(0,1]$, set parameters as in Lemma 1, Algorithm 4 ensures with probability at least $1-\delta$, the regret satisfies
+${\operatorname*{Reg}}_{T}\leq\widetilde{\mathcal{O}}\big{(}d\sqrt{{\kappa}{T}}\big{)}.$
 Remark 5.
-It is important to note that our result has a dependence on $\sqrt{\kappa}$.
-Though recent advances in the logistic bandits (Abeille et al., 2021; Zhang and Sugiyama, 2023) have shown that this dependence can be eliminated in the dominant term and shifted to lower-order terms, we believe this term remains necessary in our setting.
-This is due to the fact that the sub-optimality gap in our setting is defined as the difference between the reward $z^{\top}\theta^{*}$ while they focus the probability difference, i.e., $\sigma(z^{\top}\theta^{*})$.
-This difference prevents us from removing the $\sqrt{\kappa}$ term in our setting.
-Comparison with Das et al. (2024).
-Our algorithm attains the same sub-optimality gap as the work of Saha (2021), but with a significantly improved computational efficiency.
-Specifically, our algorithm has a time complexity of $\mathcal{O}(T)$ for the total $T$ round, which is more efficient than the $\mathcal{O}(T^{2}\log T)$ complexity of their MLE estimator.
-Moreover, our algorithm does not require storing the historical samples, resulting in only an $\mathcal{O}(1)$ storage complexity.
-In contrast, their MLE estimator necessitates storing all historical samples, leading to an $\mathcal{O}(T)$ storage complexity.
-5 Deployment Stage
-In the deployment stage, the learner uses the trained model to generate responses to user inputs and simultaneously gathering informative feedback to improve the model.
-This contrasts with the active data collection phase during training, which is focused solely on optimizing the final policy.
-In deployment, however, the learner is faced with a dual objective: selecting actions that maximize rewards to ensure a positive user experience, while also choosing actions that provide informative feedback for further model improvement.
-Performance Measure.
-To this end, we consider the cumulative regret compared to the optimal policy, which is defined as
-${\operatorname*{Reg}}_{T}=\sum_{t=1}^{T}\left(r\left(x_{t},\pi^{*}(x_{t})\right)-\frac{r\left(x_{t},a_{t}\right)+r\left(x_{t},a^{\prime}_{t}\right)}{2}\right),$
-where $\pi^{*}$ is the optimal policy, $a_{t}$ and $a^{\prime}_{t}$ are the actions selected by the learner at round $t$.
-This measure is stronger than the measure defined in Ji et al.
-(2024), which considers the gap between the optimal action and a single selected action.
-In the deployment phase, our proposed measure is more appropriate, as both actions must be sufficiently effective to ensure a positive user experience.
-While our measure is more suitable for deployment, it also introduces significant challenges.
-To optimize the measure of Ji et al.
-(2024), the learner can exploit one action with high reward for exploitation while using the other action for exploration.
-In contrast, our setting requires both actions to be good enough.
-To address this, we propose a novel algorithm that balances exploration and exploitation for deployment.
-At a high level, our algorithm consists of two key components: estimator construction and query selection.
-We describe each component in detail below.
-Estimator Construction.
-The estimator construction is the same as the process used in the training stage.
-For each round $t\in[T]$, given $\left\{\left(x_{i},a_{i},a_{i}^{\prime},y_{i}\right)\right\}_{i=1}^{t-1}$, we construct the estimator $\widetilde{\theta}_{t}$ by OMD in an one-pass manner as in Eq.
-(9).
-Specifically, the estimator is updated by
-$\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\big{\langle}g_{t}(\widetilde{\theta}_{t}),\theta\big{\rangle}+\frac{1}{2\eta}\big{\|}\theta-\widetilde{\theta}_{t}\big{\|}_{\widetilde{\mathcal{H}}_{t}}^{2}\Big{\}}.$ (15)
-Query Selection.
-Unlike in the training stage, during deployment, the learner must balance exploration and exploitation to minimize regret.
-Specifically, the learner must select queries that are both informative and likely to yield high rewards.
-Focusing solely on high-reward queries may cause the learner to miss valuable feedback for model improvement, while prioritizing only informative queries could prevent reward maximization.
-To address this, we propose a novel strategy inspired by Saha (2021), which accounts for both reward maximization and uncertainty maximization.
-Based on the estimator $\theta_{t}$, we first define the promising set $\mathcal{A}_{t}$ that only contains the actions with potential to be the best action.
-Specifically, we define the promising set as
-$\mathcal{A}_{t}=\big{\{}a\in\mathcal{A}\mid(\phi(x_{t},a)-\phi(x_{t},a^{\prime}))^{\top}\widetilde{\theta}_{t}+\widetilde{\beta}_{t}\lVert\phi(x_{t},a)-\phi(x_{t},a^{\prime})\rVert_{\mathcal{H}_{t}^{-1}}\geq 0,\forall a^{\prime}\in\mathcal{A}\big{\}}.$ (16)
-Then, we select the action pairs with the maximum uncertainty in the promising set, i.e.,
-$(a_{t},a_{t}^{\prime})=\operatorname*{arg\,max}_{(a,a^{\prime})\in\mathcal{A}_{t}\times\mathcal{A}_{t}}\left\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\right\|_{\mathcal{H}_{t}^{-1}}.$ (17)
-The overall algorithm is summarized in Algorithm 3. We show it enjoys the following regret bound.
-Algorithm 3 Deployment Stage
-0:  Regularization $\lambda>0$, failure probability $\delta\in(0,1]$
-1:  Initialize $\widetilde{\theta}_{1}=0$, $\mathcal{H}_{1}=\lambda I_{d}$.
-2:  for $t=1,2,\ldots,T$ do
-3:     Observes context $x_{t}$.
-4:     Compute the promising set by Eq. (16)
-5:     Choose duel actions by Eq. (17).
-6:     Observe preference feedback $y_{t}$
-7:     Update ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$
-8:     Compute $\widetilde{\theta}_{t+1}$ by Eq. (15)
-9:     Update $\mathcal{H}_{t+1}=\mathcal{H}_{t}+H_{t}(\widetilde{\theta}_{t+1})$
-10:  end for
-Theorem 3.
-Let $\delta\in(0,1]$, set parameters as in Lemma 1, with probability at least $1-\delta$, Algorithm 3 enjoys the following regret guarantee:
-${\operatorname*{Reg}}_{T}\leq\widetilde{\mathcal{O}}\big{(}d\sqrt{{\kappa}{T}}\big{)}.$
-Comparison with Saha (2021).
-Our result improves upon theirs in both computational and statistical efficiency.
-For statistical efficiency, our algorithm achieves a regret of $\widetilde{\mathcal{O}}\big{(}d\sqrt{{\kappa}{T}}\big{)}$, which is tighter by a factor of $\sqrt{\kappa}$ compared to their $\widetilde{\mathcal{O}}\big{(}d{\kappa}\sqrt{{T}}\big{)}$ result.
-For computational efficiency, our algorithm updates in an online manner, with $\mathcal{O}(1)$ storage and $\mathcal{O}(T)$ time complexity for the total $T$ rounds.
-In contrast, their algorithm requires storing all the data and solving an optimization problem each round, leading to $\mathcal{O}(T)$ storage and $\mathcal{O}(T^{2}\log T)$ time complexity.
-As $T$ is large or even infinite during deployment, our algorithm has a significant advantage in computational efficiency.
-6 Experiment
-In this section, we empirically evaluate the performance of our proposed method in terms of both effectiveness and efficiency.
-Specifically, we empirically validate our method with LLMs, and construct an online RLHF pipeline to evaluate different methods in both the training and deployment stage.
-[^1] Equal contribution.
-Next, we first describe the experimental setup, and then present the empirical results.
-Experimental Setup.
-In our experiments, we employ the Llama-3-8B-Instruct  [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct as the foundation model for reward model.
-We extract features $\phi(x,a)$ using last layer of the model, and the feature space dimension is $d=4096$.
-We use the Ultrafeedback-binarized dataset, a pro-processed version of the original Ultrafeedback dataset (Cui et al., 2023) for evaluation, which is a large-scale, fine-grained, diverse preference dataset, widely used as a benchmark for RLHF.
+Our result improves upon Saha et al.
+(2023) in both computational and statistical efficiency.
+Statistically, Theorem 3 improves their $\widetilde{\mathcal{O}}\big{(}d{\kappa}\sqrt{{T}}\big{)}$ result by a factor of $\sqrt{\kappa}$.
+Computationally, our algorithm has an $\mathcal{O}(1)$ time and space complexity per round, while their MLE estimator needs $\mathcal{O}(t\log t)$ time and $\mathcal{O}(t)$ space complexity at iteration $t$ due to optimization over the historical data.
+6 Practical Implementation
+In this section, we introduce the practical implementation of our proposed algorithm.
+6.1 Computation of Inverse Hessian
+The OMD update in Eq.
+(5) requires computation of the inverse of the Hessian matrix.
+Omitting the projection operation, Eq.
+(5) can be rewritten as $\widetilde{\theta}_{t+1}=\widetilde{\theta}_{t}-\eta\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t})$.
+Computing the full Hessian inverse ~$\widetilde{\mathcal{H}}_{t}^{-1}$ directly incurs a time complexity of $\mathcal{O}(d^{3})$, which is prohibitive for LLMs.
+This cost can be reduced to $\mathcal{O}(d^{2})$ by applying the Sherman-Morrison-Woodbury formula, leveraging the fact that the Hessian is a rank-one update.
+Specifically, for a matrix of the form $A+\mathbf{x}\mathbf{x}^{\top}$ where $A$ is invertible and $\mathbf{x}$ is a vector, the inverse is given by $(A+\mathbf{x}\mathbf{x}^{\top})^{-1}=A^{-1}-\frac{A^{-1}\mathbf{x}\mathbf{x}^{\top}A^{-1}}{1+\mathbf{x}^{\top}A^{-1}\mathbf{x}}$, requiring only $\mathcal{O}(d^{2})$ time.
+Nevertheless, even this reduced complexity is costly for large models.
+To further reduce the computational burden to $\mathcal{O}(d)$, we employ the Hessian-vector product technique combined with conjugate gradient descent (Boyd and Vandenberghe, 2004).
+Instead of explicitly computing ~$\widetilde{\mathcal{H}}_{t}^{-1}$, we solve the linear system ~$\widetilde{\mathcal{H}}_{t}v=g_{t}(\widetilde{\theta}_{t})$ iteratively, where $v$ is the solution we seek.
+The HVP operation ~$\widetilde{\mathcal{H}}_{t}v$ for any vector $v$ can be computed efficiently using twice differentiation: ~$\widetilde{\mathcal{H}}_{t}v=\nabla_{\theta}\big{(}\nabla_{\theta}\mathcal{L}_{t}(\theta)^{\top}v\big{)}\big{|}_{\theta=\widetilde{\theta}_{t}}+\lambda_{t}v$, where $\mathcal{L}_{t}(\theta)$ is the function whose Hessian matrix at $\widetilde{\theta}_{t}$ is ~$\widetilde{\mathcal{H}}_{t}$.
+By leveraging this efficient computation, the conjugate gradient method iteratively refines the solution.
+In practice, finding $\mathcal{L}_{t}(\theta)$ is challenging as ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$ involve complex structure.
+A key insight is ~$\widetilde{\mathcal{H}}_{t}$ can be viewed as the cumulative Hessian of the $\ell_{t}(\theta)$ approximately.
+This motivates the use of a more tractable surrogate $\ell_{t}(\theta)$ in Eq.
+(3) while keeping the update.
+To further approximate the effect of curvature accumulation, we adopt an adaptive damping scheme that approximately preserves historical Hessian information while incorporating new Hessian updates: $\lambda_{t}=\lambda_{0}\cdot\min\{1,f(t/T)\}$, where $f(\cdot)$ is a monotonic increasing function, such as linear functions.
+6.2 Computation of Model Uncertainty
+In both online RLHF with active data collection and deployment-time adaptation, our algorithm utilizes uncertainty-driven query selection strategies.
+While quantifying uncertainty using the local norm induced by the Hessian matrix offers strong theoretical guarantees, it is computationally prohibitive in practice.
+To address this challenge, we adopt a rejection sampling-based approximation, a technique commonly employed for exploration in the literature (Nakano et al., 2021; Gulcehre et al., 2023; Dong et al., 2023, 2024).
+Specifically, given a prompt, we sample $n$ independent responses by the current model, then use the trained reward model to rank the responses.
+Then, we use different strategies to select the response for different settings.
+Specifically, In active data collection, the key insight is to identify and query samples that exhibit the greatest diversity in prompt action features.
+To this end, we select the response with the highest predicted reward and the response with the lowest predicted reward.
+In deployment-time adaptation, the core idea is to select the first action to maximize the estimated reward, while the second is chosen to balance high reward with sufficient divergence from the first.
+Concretely, we select the response with the highest predicted reward and another from the top-$1/q$ percentile of the reward, where $q$ is a hyperparameter.
+7 Experiments
+In this section, we empirically evaluate the performance of our proposed method.  [^1] Equal contribution. We first describe the experimental setup, and then present the empirical results.
+7.1 Experiment setup
+In our experiments, we employ the Llama-3-8B-Instruct 222huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct and Qwen2.5-7B-Instruct 333huggingface.co/Qwen/Qwen2.5-7B-Instruct as the foundation model for reward model.
+We extract features $\phi(x,a)$ using the last layer of the model, and the dimension is $d=4096$.
+We use two datasets for evaluation.
+The first one is Ultrafeedback-binarized dataset 444huggingface.co/datasets/OpenRLHF/preference_dataset_mixture2_and_safe_pku, a pre-processed version of the original Ultrafeedback dataset (Cui et al., 2023), a widely used benchmark for RLHF.
 It collects about $64,000$ prompts from diverse resources, including question answering, summarization, and dialogue generation.
-Each data consists of a context $x$, two responses $a$ and $a^{\prime}$, and a human preference label $y$.
-11footnotetext: Our online RLHF pipeline code is available at https://github.com/ZinYY/RLHF_Pipeline
-Implementation Details.
-All experiments in this section were performed using four Nvidia A800 (80 GB) GPUs with two Intel(R) Xeon(R) Gold 6430 CPUs.
-The learning rate is set to $1\times 10^{-3}$ for all methods.
-We use DeepSpeed (Rasley et al., 2020) with mixed-precision training (BF16).
-The maximum sequence length is set to 1024 tokens, and we use a global batch size of 32.
-For practical efficiency, we implement our method using Hessian-vector product (HVP) and conjugate gradient methods, which efficiently approximate the Hessian matrix without the need for computing $d^{2}$ matrix.
-We use the default partition of the Ultrafeedback-binarized dataset with the standard train_prefs split for training and test_prefs split for evaluation.
-Furthermore, we also enable flash attention for the improved training efficiency.
-6.1 Training Stage
+Each data consists of a context $x$, two responses $a$ and $a^{\prime}$, and a preference label $y$.
+We also employ a mixed dataset, Mixture2 dataset 444huggingface.co/datasets/OpenRLHF/preference_dataset_mixture2_and_safe_pku, which combines a variety of preference datasets, including HH-RLHF, SHP, UltraFeedback, etc.
+11footnotetext: The code is available at https://github.com/ZinYY/Online_RLHF
 (a) training loss
@@
 (d) evaluation accuracy
-Figure 1: For the training stage with passive data collection, we compare our method with MLE (Zhu et al., 2023). We report the average accuracy and loss of the reward model during training.
-In this part, we conduct experiments to evaluate the performance of our method in the training stage, including both passive and active data collection strategies.
-Contenders.
-In the training stage, we compare with:
-  • Random: A baseline that randomly selects queries.
-  • Passive-MLE (Zhu et al., 2023): A method that uses maximum likelihood estimation to learn the reward function and compute the pessimistic policy based on the learned reward model.
-  • Active-MLE (Das et al., 2024): A method using MLE to optimize the reward model and selects queries based on uncertainty, i.e., the maximum difference of two actions.
-We compared our method with other methods in terms of the accuracy of the reward model, as the method for training the final policy based on the learned reward model is the same as contenders.
-Passive Data Collection Results.
-We first evaluate our method in the passive data collection setting.
-We randomly sample $T=30,000$ data points from the Ultrafeedback dataset for training.
-Figure 1 shows the loss and accuracy vs.
-the number of training samples.
-We observe that our method converges faster to a lower loss and achieves a higher evaluation accuracy compared to baselines.
-The improvement is particularly pronounced in the small-sample regime ($T<10,000$), where our method achieves a higher evaluation accuracy with the same amount of samples compared to Passive-MLE which employs conventional stochastic gradient descent (SGD) updates.
-This demonstrates the superior statistical efficiency of our approach, therefore achieving a better performance with fewer training samples.
-Active Data Collection Results.
-We then evaluate our method in the active data collection setting.
-In this scenario, the global batch size is set to 8, and we only allow the method to select 6, 400 samples out of the whole training datasets for training according to different selection strategies.
-Table 2 and Figure 2 demonstrate that our active learning strategy achieves comparable performance to passive learning while using only 21% of the labeled data.
-Specifically, our method achieves an average accuracy of 70.43% after 6,400 actively selected samples, which is comparable with the previous state-of-the-art method, Active-MLE, while our method is 3 times faster because we do not need to recompute all the selected data in the buffer, but we update the model in an online manner.
-The training loss curves in Figure 2(a) also show that our method converges faster compared to other contenders.
-This improved sample efficiency can be attributed to our theoretically-motivated query selection criterion that effectively identifies the most informative samples, which further validates the two key factors in our method: (1) the use of local norm $\mathcal{H}_{t}$ that better captures the uncertainty in different directions, and (2) the computationally efficient one-pass update rule that enables processing large-scale datasets.
+Figure 2: For Llama-3-8B-Instruct on Ultrafeedback-binarized dataset with passive data collection, we report the comparison of MLE and our OMD-based method for the reward model about (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy.
+(a) training loss
+(b) training accuracy
+(c) evaluation loss
+(d) evaluation accuracy
+Figure 3: For Qwen2.5-7B-Instruct on Ultrafeedback-binarized dataset with passive data collection, we report the comparison of MLE and our OMD-based method for the reward model about (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy.
+(a) training loss
+(b) training accuracy
+(c) evaluation loss
+(d) evaluation accuracy
+Figure 4: For Llama-3-8B-Instruct on Mixture2 dataset with passive data collection, we report the comparison of MLE and our OMD-based method for the reward model about (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy.
+7.2 Experimental results
+In this section, we present the experimental results, including passive data collection, active data collection and deployment-time adaptation.
+7.2.1 Passive Data Collection
+We randomly sample $T=30,000$ data points from the dataset for training the reward model.
+We evaluate both the standard MLE-based method, which uses a stochastic gradient descent (SGD) optimizer, and our proposed OMD-based method in terms of reward model loss and accuracy.
+Figures 23, and 4 present the results for three settings: (i) the Llama-3-8B-Instruct model on the Ultrafeedback-binarized dataset, (ii) the Qwen2.5-7B model on the same dataset, and (iii) the Llama-3-8B-Instruct model on the Mixture2 dataset.
+We report the results across four key metrics: (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy.
+Our method demonstrates faster convergence to a lower loss and achieves higher evaluation accuracy compared to the MLE baseline.
+These results highlight the superior statistical efficiency of the OMD-based approach, which achieves improved performance with fewer training samples.
 (a) training loss
 (b) evaluation accuracy
-Figure 2: For the training stage with active data collection, we report (a) the training loss vs. the training iterations, and (b) the evaluation accuracy vs. the training iterations.
 Method ACC (%) Time (s)
@@
 Ours 70.43$\pm$0.3 1489$\pm$36
-Table 2: Comparison of different methods in terms of evaluation accuracy and training time in seconds. Results are averaged over five runs.
-6.2 Deployment Stage
-In this part, we evaluate our method on deployment stage.
-Contenders.
-In the deployment stage, we compare with:
-  • Random: A baseline strategy that randomly select responses from responses generated by the current policy model.
-  • Best-Two: A baseline strategy that selects the best two responses from responses generated by the current policy model.
-  • Iterative-RLHF (Dong et al., 2024): A method that takes the best and the worst responses from the current policy model to construct a new preference pair for RLHF iteratively.
-Deployment Stage Analysis.
-For the scenario of deployment stage, we evaluate on a pre-processed online-variant of the Ultrafeedback-binarized dataset, where we divide the dataset into 20 chunks and process them sequentially to simulate the online deployment scenario.
-As shown in Figure 3, our method consistently outperforms the baselines in terms of both average cumulative rewards and win rates.
-The average cumulative rewards in Figure 3(a) show that our method learns more effectively from interactions, achieving higher rewards over time.
-The win rates in Figure 3(b) demonstrate that our method makes better decisions when selecting between response pairs, outperforming both the baseline methods of Random and Best-Two, and also better than Iterative-RLHF which selects the best and worst response pairs for training.
-The superior performance of our method further validates the effectiveness of our theoretically grounded action selection strategy.
-This success stems from an efficient online update mechanism that enables rapid adaptation to new preference feedback, as well as the design of the promising set, which balances high-reward responses for immediate performance with sufficient exploration for future model improvements.
-These results align with our theoretical analysis in Section 5, validating the practical effectiveness of our proposed algorithm in the deployment stage.
-(a) average cumulative rewards
-(b) win rates
-Figure 3: For the deployment stage, we divide the Ultrafeedback dataset into 20 chunks, and report (a) the average cumulative rewards and the (b) win rates of different methods.
-7 Conclusion and Future Work
-In this work, we present a unified framework of the RLHF pipeline from the view of contextual bandits, leveraging the Bradley-Terry preference model with a linearly parameterized reward function.
-We decompose the RLHF process into two stages: (post-)training and deployment, exploring both passive and active data collection strategies during the training phase.
-We then develop novel algorithms for each stage, applying online mirror descent with a tailored local norm to replace standard maximum likelihood estimation, and designing distinct query criteria for different stages.
-Our theoretical analysis shows that our method improves upon existing methods in both statistical and computational efficiency.
-Finally, we train and deploy Llama-3-8B-Instruct on the Ultrafeedback-binarized dataset using our approach.
-The empirical results validate our theoretical findings and confirm the effectiveness of our proposed method.
-While our work advances the theoretical understanding of RLHF, several important directions remain for future work.
-First, we considered the fixed feature mapping $\phi$ for the reward model, but in practice, this mapping may evolve during the training process.
-Thus, analyzing this dynamic feature mapping is an important future work.
-Second, though we focused on the Bradley-Terry model, there are other preference models for human feedback, such as the Plackett-Luce model (Luce, 1959; Plackett, 1975).
-Extending our results to these preference models is another promising direction.
+(c) evaluation accuracy and training time
+Figure 5: For Llama-3-8B-Instruct on Ultrafeedback-binarized dataset with active data collection, we report the comparison of different methods about (a) training loss, (b) evaluation accuracy and (c) final evaluation accuracy and training time.
+(a) training loss
+(b) evaluation accuracy
+Method ACC (%) Time (s)
+Rand-MLE 67.36 $\pm$ 0.5 4653 $\pm$ 121
+Active-MLE 67.25 $\pm$ 0.4 4701 $\pm$ 103
+Rand-OMD 66.80 $\pm$ 0.5 1312 $\pm$ 47 0
+Ours 67.11 $\pm$ 0.5 1325 $\pm$ 54 0
+(c) evaluation accuracy and training time
+Figure 6: For Qwen2.5-7B-Instruct on Ultrafeedback-binarized dataset with active data collection, we report the comparison of different methods about (a) training loss, (b) evaluation accuracy and (c) final evaluation accuracy and training time.
+7.2.2 Active Data Collection
+In this experimental setup, we constrain the algorithm to select only 6,400 samples from the full training dataset, based on different data selection strategies.
+We evaluate the performance of both the standard MLE-based method and our proposed OMD-based method under this limited data regime.
+To assess the effectiveness of the data selection strategy itself, we compare our approach against a random selection baseline.
+We report the results across four key metrics: (a) training loss curve, (b) evaluation accuracy curve and (c) final evaluation accuracy and training time.
+Figure 5 and Figure 6 demonstrate that our OMD-based method achieves competitive performance with the MLE-based method for both data collection strategies, while improving the training time by significantly.
+Moreover, our data selection strategy outperforms the random selection strategy, demonstrating that our method can effectively select informative data to improve the performance.
+7.2.3 Deployment-Time Adaptation
+To emulate real-world deployment scenarios, we partition the dataset into 20 sequential chunks and process them incrementally.
+This setup allows us to evaluate the adaptability and robustness of different action selection strategies over time.
+We benchmark our proposed selection strategy against three baselines: (i): random selection, (ii): the best and second best actions, and (iii): the best and worst actions.
+We combine the above strategies with MLE-based and OMD-based reward estimators.
+We report three key metrics: (a) the average cumulative reward for MLE-based methods, (b) the average cumulative reward for OMD-based methods, and (c) the win rate of OMD-based methods over MLE-based counterparts under identical selection strategies.
+As illustrated in Figure 7 and Figure 8, our action selection strategy consistently outperforms the baselines across both estimator types.
+These results highlight the strength of our approach in balancing the exploitation of high-reward responses with sufficient exploration to drive continual model improvement.
+Moreover, the win rate comparisons suggest that our OMD-based estimator shows competitive performance relative to its MLE-based counterpart, further validating its practical effectiveness.
+(a) Reward of MLE-based methods
+(b) Reward of OMD-based methods
+(c) Win rates, all methods
+Figure 7: For Llama-3-8B-Instruct on Ultrafeedback-binarized dataset with deployment-time adaptation, we report (a) cumulative reward of MLE-based methods, (b) cumulative reward of OMD-based methods, and (c) win rates between MLE-based and OMD-based estimators with the same action selection strategy.
+(a) Reward of MLE-based methods
+(b) Reward of OMD-based methods
+(c) Win rates, all methods
+Figure 8: For Qwen2.5-7B-Instruct on Ultrafeedback-binarized dataset with deployment-time adaptation, we report (a) cumulative reward of MLE-based methods, (b) cumulative reward of OMD-based methods, and (c) win rates between MLE-based and OMD-based estimators with the same action selection strategy.
+8 Conclusion
+In this work, we study the problem of online RLHF from the perspective of contextual bandits.
+To address the linear growth in computational complexity with respect to the number of iterations in reward modeling, we propose a novel one-pass reward modeling algorithm that achieves constant-time complexity per iteration.
+Our method is built upon the online mirror descent framework and incorporates a carefully designed local norm that captures second-order information.
+We apply our method to a broad range of online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation, and design tailored algorithms for each scenario.
+We provide theoretical guarantees that our approach achieves superior statistical and computational efficiency relative to existing methods.
+Finally, we provide practical implementations of our method and conduct experiments using the Llama-3-8B-Instruct and Qwen2.5-7B-Instruct models on the Ultrafeedback-binarized and Mixture2 datasets, empirically validating its practical effectiveness.
+While our work advances both the statistical and computational understanding of online RLHF, several important directions remain for future exploration.
+First, we assume a fixed feature mapping for the reward model; however, in practice, this mapping may evolve throughout the training process.
+Analyzing the impact of such dynamically changing feature representations presents a compelling direction for future research.
+Second, although our analysis is based on the Bradley-Terry model, extending the framework to other preference models, such as the Plackett-Luce model (Luce, 1959; Plackett, 1975), is another promising avenue that may broaden the applicability of our results.
 Appendix A Useful Lemmas
 Lemma 2.
-For any $t\in[T]$, define the second-order approximation of the loss function $\ell_{t}(\theta)$ at the estimator $\widetilde{\theta}_{t}$ as
-~$\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+\langle\nabla\ell_{t}(\widetilde{\theta}_{t}),\theta-\widetilde{\theta}_{t}\rangle+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}.$
-Then, for the following update rule
+For any $t\in[T]$, define the second-order approximation of the loss function $\ell_{t}(\theta)$ at the estimator $\widetilde{\theta}_{t}$ as ~$\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+\langle\nabla\ell_{t}(\widetilde{\theta}_{t}),\theta-\widetilde{\theta}_{t}\rangle+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}.$ Then, for the following update rule
 $\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\left\{\widetilde{\ell}_{t}(\theta)+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{t}\rVert_{\mathcal{H}_{t}}^{2}\right\},$
@@
 where the last inequality is by $\|\widetilde{\theta}_{1}-\theta^{*}\|_{\mathcal{H}_{1}}^{2}\leq\lambda\|\widetilde{\theta}_{1}-\theta^{*}\|_{2}^{2}\leq 4\lambda B^{2}$.
-Set $\zeta=2\eta$ finishes the proof.
+Set $\zeta=2\eta$ ends the proof.
 ∎
@@
 To address this issue, following the analysis in Zhang and Sugiyama [2023], we decompose the loss gap into two components by introducing an intermediate term.
-Specifically, we define the softmax function as
-$[\sigma_{i}(q)]_{1}=\frac{\exp(q)}{1+\exp(q)}\quad\text{and}\quad[\sigma_{i}(q)]_{0}=\frac{1}{1+\exp(q)},$
-where $[\cdot]_{i}$ denotes the $i$-th element of the vector. Then, the loss function $\ell_{i}(\theta)$ can be rewritten as
+Specifically, we define the softmax function as $[\sigma_{i}(q)]_{1}=\frac{\exp(q)}{1+\exp(q)}$ and $[\sigma_{i}(q)]_{0}=\frac{1}{1+\exp(q)}$, where $[\cdot]_{i}$ denotes the $i$-th element of the vector.
+Then, the loss function $\ell_{i}(\theta)$ can be rewritten as
 $\ell(q_{t},y_{t})=-\mathds{1}_{\{y_{t}=1\}}\cdot\log\left({[\sigma(q_{t})]_{1}}\right)-\mathds{1}_{\{y_{t}=0\}}\cdot\log\left({[\sigma(q_{t})]_{0}}\right).$
-Then, we define the pseudo-inverse function of $\sigma^{-1}(p)$ with
-$[\sigma^{-1}(p)]=\log({q}/({1-q}))\quad\text{and}\quad[\sigma^{-1}(p)]_{0}=\log((1-q)/{q}).$
+Then, we define the pseudo-inverse function of $\sigma^{-1}(p)$ with $[\sigma^{-1}(p)]_{1}=\log({q}/({1-q}))$ and $[\sigma^{-1}(p)]_{0}=\log((1-p)/{p})$.
 Then, we decompose the regret into two terms by introducing an intermediate term.
-$\sum_{i=1}^{t}\ell_{s}\left(\theta^{*}\right)-\sum_{i=1}^{t}\ell_{s}(\widetilde{\theta}_{i+1})=$ $\underbrace{\sum_{i=1}^{t}\ell_{s}\left(\theta^{*}\right)-\sum_{i=1}^{t}\ell_{s}(q_{i},y_{i})}_{\mathtt{term~{}(a)}}+\underbrace{\sum_{i=1}^{t}\ell_{s}(q_{i},y_{i})-\sum_{i=1}^{t}\ell_{s}(\widetilde{\theta}_{i+1})}_{\mathtt{term~{}(b)}}$
+$\sum_{i=1}^{t}\ell_{i}\left(\theta^{*}\right)-\sum_{i=1}^{t}\ell_{i}(\widetilde{\theta}_{i+1})=$ $\underbrace{\sum_{i=1}^{t}\ell_{i}\left(\theta^{*}\right)-\sum_{i=1}^{t}\ell_{i}(q_{i},y_{i})}_{\mathtt{term~{}(a)}}+\underbrace{\sum_{i=1}^{t}\ell_{i}(q_{i},y_{i})-\sum_{i=1}^{t}\ell_{i}(\widetilde{\theta}_{i+1})}_{\mathtt{term~{}(b)}}$
 where $q_{i}$ is an aggregating forecaster for logistic loss defined by $q_{i}=\sigma^{-1}(\mathbb{E}_{\theta\sim P_{i}}[\sigma(\theta^{\top}z_{i})])$ and $P_{i}=\mathcal{N}(\widetilde{\theta}_{i},(1+c\mathcal{H}_{i}^{-1}))$ is the Gaussian distribution with mean $\widetilde{\theta}_{i}$ and covariance $(1+c\mathcal{H}_{i}^{-1})$, where $c>0$ is a constant to be specified later.
@@
 Combing the above two bounds, we have
-$\big{\|}\widetilde{\theta}_{t+1}-\theta^{*}\big{\|}_{\mathcal{H}_{t+1}}^{2}$ $\leq 2\eta\left(3\log(1+2t)+2+BL\right)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right)$
-$+4\eta+2\eta\sqrt{6}cd\log\left(1+\frac{2tL^{2}}{d\lambda}\right)+4\lambda L_{\theta}^{2}+12\sqrt{2}BL^{3}\eta\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{2}^{2}+\left(\frac{\eta}{c}-1\right)\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{\mathcal{H}_{i}}^{2}.$
+$\big{\|}\widetilde{\theta}_{t+1}-\theta^{*}\big{\|}_{\mathcal{H}_{t+1}}^{2}$ $\leq 12\sqrt{2}BL^{3}\eta\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{2}^{2}+\left(\frac{\eta}{c}-1\right)\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{\mathcal{H}_{i}}^{2}+C.$
+where $C=2\eta\left(3\log(1+2t)+2+BL\right)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right)+4\eta+2\eta\sqrt{6}cd\log\left(1+\frac{2tL^{2}}{d\lambda}\right)+4\lambda B^{2}$.
 Setting $c=7\eta/6$ and $\lambda\geq 84\sqrt{2}BL^{3}\eta$, we have
-$12\sqrt{2}BL^{3}\eta\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{2}^{2}+\left(\frac{\eta}{c}-1\right)\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{\mathcal{H}_{i}}^{2}\leq\left(12\sqrt{2}BL^{3}\eta-\frac{\lambda}{7}\right)\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{2}^{2}\leq 0$
+$12\sqrt{2}BL^{3}\eta\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{2}^{2}+\left(\frac{\eta}{c}-1\right)\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{\mathcal{H}_{i}}^{2}$
+[…]
+[…]
 Note that $84\sqrt{2}\left(BL^{3}+dL^{2}\right)\eta\geq\max\left\{2L^{2},72cdL^{2},84\sqrt{2}BL^{3}\eta\right\}$, so we set $\lambda\geq 84\sqrt{2}\left(BL^{3}+\right.$ $\left.dL^{2}\right)\eta$.
@@
 Proof.
-Define $J(\pi)=\mathbb{E}_{x\sim\rho}[r\left(x,\pi(x)\right)]$, $J^{\prime}(\pi)=J(\pi)-\langle\theta^{*},v\rangle$, we have
-$\operatorname*{\mathsf{SubOpt}}\left(\pi_{T}\right)=\left(J^{\prime}\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)\right)+\left(\widetilde{J}\left(\pi^{*}\right)-\widetilde{J}\left(\pi_{T}\right)\right)+\left(\widetilde{J}\left(\pi_{T}\right)-J^{\prime}\left(\pi_{T}\right)\right).$
+Define $J(\pi)=\mathbb{E}_{x\sim\rho}[r\left(x,\pi(x)\right)]$, we have
+$\operatorname*{\mathsf{SubOpt}}\left(\pi_{T}\right)=\left(J\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)\right)+\left(\widetilde{J}\left(\pi^{*}\right)-\widetilde{J}\left(\pi_{T}\right)\right)+\left(\widetilde{J}\left(\pi_{T}\right)-J\left(\pi_{T}\right)\right).$
 Since $\pi_{T}$ is the optimal policy under expected value $\widetilde{J}(\pi)$, i.e., $\widetilde{J}(\pi_{T})=\max_{\pi\in\Pi}\widetilde{J}(\pi)$, we have
@@
 For the third term, we have with probability at least $1-\delta$, it holds that
-$\widetilde{J}\left(\pi_{T}\right)-J^{\prime}\left(\pi_{T}\right)=\min_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\theta^{\top}(\phi(s,\pi_{T}(s))-v)\right]-\mathbb{E}_{x\sim\rho}\left[\theta^{*\top}(\phi(s,\pi_{T}(s))-v)\right]\leq 0,$ (19)
+$\widetilde{J}\left(\pi_{T}\right)-J\left(\pi_{T}\right)=\min_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\theta^{\top}\phi(s,\pi_{T}(s))\right]-\mathbb{E}_{x\sim\rho}\left[\theta^{*\top}\phi(s,\pi_{T}(s))\right]\leq 0,$ (12)
 where the last inequality holds by $\theta^{*}\in\mathcal{C}_{T}$ with probability at least $1-\delta$.
 For the first term, we have with probability at least $1-\delta$, it holds that
-$J^{\prime}\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)$
-=\ = $\mathbb{E}_{x\sim\rho}\left[(\theta^{*})^{\top}(\phi(s,\pi^{*}(s))-v)\right]-\min_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\theta^{\top}(\phi(s,\pi^{*}(s))-v)\right]$
-=\ = $\sup_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\left(\theta^{*}-\widetilde{\theta}_{T}+\widetilde{\theta}_{T}-\theta\right)^{\top}\left(\phi(x,\pi^{*}(x))-v\right)\right]$
-=\ = $\mathbb{E}_{x\sim\rho}\left[\left(\theta^{*}-\widetilde{\theta}_{T}\right)^{\top}\left(\phi(x,\pi^{*}(x))-v\right)\right]+\sup_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\left(\widetilde{\theta}_{T}-\theta\right)^{\top}\left(\phi(x,\pi^{*}(x))-v\right)\right]$
+$J\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)=$ $\mathbb{E}_{x\sim\rho}\left[(\theta^{*})^{\top}\phi(s,\pi^{*}(s))\right]-\min_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\theta^{\top}\phi(s,\pi^{*}(s))\right]$
+=\ = $\sup_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\left(\theta^{*}-\widetilde{\theta}_{T}+\widetilde{\theta}_{T}-\theta\right)^{\top}\phi(x,\pi^{*}(x))\right]$
+=\ = $\mathbb{E}_{x\sim\rho}\left[\left(\theta^{*}-\widetilde{\theta}_{T}\right)^{\top}\phi(x,\pi^{*}(x))\right]+\sup_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\left(\widetilde{\theta}_{T}-\theta\right)^{\top}\phi(x,\pi^{*}(x))\right]$
 […]
 where the first inequality holds by the Cauchy-Schwarz inequality.
-Since it holds $\widetilde{\theta}_{T}\in\mathcal{C}_{T}$ with probability at least $1-\delta$ by Lemma 1, we have $\lVert\theta^{*}-\widetilde{\theta}_{T}\rVert_{\mathcal{H}_{T}}\leq\widetilde{\beta}_{T}$ and $\sup_{\theta\in\mathcal{C}_{T}}\lVert\theta-\widetilde{\theta}_{T}\rVert_{\mathcal{H}_{T}}\leq\widetilde{\beta}_{T}$.
+Since it holds $\theta^{*}\in\mathcal{C}_{T}$ with probability at least $1-\delta$ by Lemma 1, we have $\lVert\theta^{*}-\widetilde{\theta}_{T}\rVert_{\mathcal{H}_{T}}\leq\widetilde{\beta}_{T}$ and $\sup_{\theta\in\mathcal{C}_{T}}\lVert\theta-\widetilde{\theta}_{T}\rVert_{\mathcal{H}_{T}}\leq\widetilde{\beta}_{T}$.
 Thus, we obtain
-$J^{\prime}\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)\leq$ $2\widetilde{\beta}_{T}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}.$ (20)
+$J\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)\leq$ $2\widetilde{\beta}_{T}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big{\rVert}_{\mathcal{H}_{T}^{-1}}.$ (13)
 Combining Eq. (11), Eq. (12), and Eq. (13) and substituting $\widetilde{\beta}_{T}=\mathcal{O}(\sqrt{d}(\log(T/\delta))^{2})$, we have with probability at least $1-\delta$, it holds that
-$\operatorname*{\mathsf{SubOpt}}\left(\pi_{T}\right)$ $=\left(J^{\prime}\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)\right)+\left(\widetilde{J}\left(\pi^{*}\right)-\widetilde{J}\left(\pi_{T}\right)\right)+\left(\widetilde{J}\left(\pi_{T}\right)-J^{\prime}\left(\pi_{T}\right)\right)$
-$\leq 2\widetilde{\beta}_{T}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}$
-$\leq\mathcal{O}\left(\sqrt{d}\left(\log\frac{T}{\delta}\right)^{2}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}\right).$
+$\operatorname*{\mathsf{SubOpt}}\left(\pi_{T}\right)\leq 2\widetilde{\beta}_{T}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big{\rVert}_{\mathcal{H}_{T}^{-1}}\leq\mathcal{O}\left(\sqrt{d}\left(\log\frac{T}{\delta}\right)^{2}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big{\rVert}_{\mathcal{H}_{T}^{-1}}\right).$
 This completes the proof. ∎
@@
 Furthermore, by the definition of $\mathcal{H}_{t}$, we have
-$\mathcal{H}_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}\dot{\sigma}\left(z_{s}^{\top}\widetilde{\theta}_{s+1}\right)z_{s}z_{s}^{\top}\geq\lambda I_{d}+\frac{1}{\kappa}\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}=\frac{1}{\kappa}\left(\kappa\lambda I_{d}+\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}\right)\triangleq\frac{1}{\kappa}V_{t}.$
+$\mathcal{H}_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}\dot{\sigma}\left(z_{s}^{\top}\widetilde{\theta}_{s+1}\right)z_{s}z_{s}^{\top}\geq\lambda I_{d}+\frac{1}{\kappa}\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}=\frac{1}{\kappa}\left(\kappa\lambda I_{d}+\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}\right)=\frac{1}{\kappa}V_{t}.$
 Thus, we have
@@
 =\ = $\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\theta^{*}+\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}\theta^{*}$
-=\ = $\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\widetilde{\theta}_{t}+\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}(\theta^{*}-\widetilde{\theta}_{t})$
-$+\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}\widetilde{\theta}_{t}+\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}(\theta^{*}-\widetilde{\theta}_{t})$
+=\ = $2\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\theta^{*}+\big{(}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}\theta^{*}$
+=\ = $2\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}(\theta^{*}-\widetilde{\theta}_{t})+2\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\widetilde{\theta}_{t}$
+$+\big{(}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}(\theta^{*}-\widetilde{\theta}_{t})+\big{(}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}\widetilde{\theta}_{t}$
 […]
-$+\widetilde{\beta}_{t}\big{\lVert}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a^{\prime}_{t})\big{\rVert}_{\mathcal{H}_{t}^{-1}}+\big{\lVert}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a^{\prime}_{t})\big{\rVert}_{\mathcal{H}_{t}^{-1}}\big{\lVert}\theta^{*}-\widetilde{\theta}_{t}\big{\rVert}_{\mathcal{H}_{t}}$
+$+\big{\lVert}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{\rVert}_{\mathcal{H}_{t}^{-1}}\big{\lVert}\theta^{*}-\widetilde{\theta}_{t}\big{\rVert}_{\mathcal{H}_{t}}+\big{(}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}\widetilde{\theta}_{t}$
 […]
-$+\widetilde{\beta}_{t}\lVert\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\rVert_{\mathcal{H}_{t}^{-1}}+\widetilde{\beta}_{t}\lVert\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\rVert_{\mathcal{H}_{t}^{-1}}$
+$+\widetilde{\beta}_{t}\big{\lVert}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{\rVert}_{\mathcal{H}_{t}^{-1}}$
 […]
-Here, the first inequality holds since both $a_{t},a^{\prime}_{t}\in\mathcal{A}_{t}$ and by the definition of $\mathcal{A}_{t}$, it implies that
-$\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\widetilde{\theta}_{t}$ $\leq\widetilde{\beta}_{t}\big{\lVert}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{\rVert}_{\mathcal{H}_{t}^{-1}},$
-$\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a^{\prime}_{t})\big{)}^{\top}\widetilde{\theta}_{t}$ $\leq\widetilde{\beta}_{t}\big{\lVert}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big{\rVert}_{\mathcal{H}_{t}^{-1}}.$
-The second inequality holds by selection strategy of $(a_{t},a_{t}^{\prime})$, it holds that
-$\lVert\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\rVert_{\mathcal{H}_{t}^{-1}}$ $\leq\lVert\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\rVert_{\mathcal{H}_{t}^{-1}},$
-$\lVert\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a^{\prime}_{t})\rVert_{\mathcal{H}_{t}^{-1}}$ $\leq\lVert\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\rVert_{\mathcal{H}_{t}^{-1}}.$
+$+\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}\widetilde{\theta}_{t}+\widetilde{\beta}_{t}\big{\lVert}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{\rVert}_{\mathcal{H}_{t}^{-1}}$
+=\ = $3\widetilde{\beta}_{t}\big{\lVert}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{\rVert}_{\mathcal{H}_{t}^{-1}},$
+where the first inequality holds by the Holder’s inequality and the arm selection strategy of $a_{t}$ such that $\phi(x_{t},\pi^{*}(x_{t}))^{\top}\widetilde{\theta}_{t}\leq\phi(x_{t},a_{t})^{\top}\widetilde{\theta}_{t}$, the second inequality holds by $\widetilde{\theta}_{t}\in\mathcal{C}_{t}$ with probability at least $1-\delta$ by Lemma 1, the third inequality holds by arm selection strategy of $a_{t}^{\prime}$ such that $a_{t}^{\prime}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\phi(x_{t},a)^{\top}\widetilde{\theta}_{t}+2\widetilde{\beta}\lVert\phi(x_{t},a)-\phi(x_{t},a_{t})\rVert_{\mathcal{H}_{t}^{-1}}$.
+By the definition of $\mathcal{H}_{t}$, we have
+$\mathcal{H}_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}\dot{\sigma}\left(z_{s}^{\top}\widetilde{\theta}_{s+1}\right)z_{s}z_{s}^{\top}\geq\lambda I_{d}+\frac{1}{\kappa}\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}=\frac{1}{\kappa}\left(\kappa\lambda I_{d}+\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}\right)=\frac{1}{\kappa}V_{t}.$
 Thus, we have
-${\operatorname*{Reg}}_{T}=\frac{1}{2}\Big{(}\big{(}r(x_{t},\pi^{*}(x_{t}))-r(x_{t},a_{t})\big{)}+\big{(}r(x_{t},\pi^{*}(x_{t}))-r(x_{t},a_{t}^{\prime})\big{)}\Big{)}\leq 2\widetilde{\beta}_{T}\sum_{t=1}^{T}\lVert\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\rVert_{\mathcal{H}_{t}^{-1}}.$
-Furthermore, by the definition of $\mathcal{H}_{t}$, we have
-$\mathcal{H}_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}\dot{\sigma}\left(z_{s}^{\top}\widetilde{\theta}_{s+1}\right)z_{s}z_{s}^{\top}\geq\lambda I_{d}+\frac{1}{\kappa}\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}=\frac{1}{\kappa}\left(\kappa\lambda I_{d}+\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}\right)\triangleq\frac{1}{\kappa}V_{t}$
-Thus, we obtain
 $\sum_{t=1}^{T}\left\|z_{t}\right\|_{\mathcal{H}_{t}^{-1}}\leq\sqrt{\kappa}\sum_{t=1}^{T}\left\|z_{t}\right\|_{V_{t}^{-1}}\leq\sqrt{\kappa}\sqrt{T\sum_{t=1}^{T}\left\|z_{t}\right\|^{2}_{V_{t}^{-1}}}\leq\sqrt{2\kappa dT\log\left(1+\frac{4TL^{2}}{\lambda\kappa d}\right)},$
 where the first inequality holds by the fact that $\mathcal{H}_{t}\succeq\frac{1}{\kappa}V_{t}$, the second inequality holds by the Cauchy-Schwarz inequality, and the last inequality holds by the elliptic potential lemma in Lemma 4.
-Thus, we have with probability at least $1-\delta$,
-${\operatorname*{Reg}}_{T}\leq 2{\widetilde{\beta}_{T}}\sqrt{2\kappa dT\log\left(1+\frac{4TL^{2}}{\lambda\kappa d}\right)}\leq\widetilde{\mathcal{O}}\left(d\sqrt{{\kappa}{T}}\right).$
-This finishes the proof. ∎
+Therefore, we have
+${\operatorname*{Reg}}_{T}\leq\frac{3}{2}{\widetilde{\beta}_{T}}\sqrt{2\kappa dT\log\left(1+\frac{4\kappa TL^{2}}{\lambda d}\right)}\leq\widetilde{\mathcal{O}}\big{(}d\sqrt{\kappa T}\big{)}.$
+where the This completes the proof. ∎
 Appendix F Supporting Lemmas
@@
 Let $V_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}x_{s}x_{s}^{\top}$ for $\lambda\geq 0$.
-Then, we have
-$\sum_{s=1}^{t}\left\|z_{s}\right\|_{V_{s}^{-1}}^{2}\leq 2d\log\left(1+\frac{tL^{2}}{\lambda d}\right).$
+Then, we have $\sum_{s=1}^{t}\left\|z_{s}\right\|_{V_{s}^{-1}}^{2}\leq 2d\log\left(1+\frac{tL^{2}}{\lambda d}\right)$.
 Lemma 5 (Campolongo and Orabona [2020, Proposition 4.1]).
@@
 For each prompt, four model completions were generated using various open-source and proprietary language models, with GPT-4 providing comprehensive evaluations across multiple criteria including helpfulness, honesty, and truthfulness.
-The binarized version was constructed by selecting the completion with the highest overall score as the ”chosen” response and randomly selecting one of the remaining completions as the ”rejected” response, creating clear preference pairs suitable for reward modeling and direct preference optimization.
+The binarized version was constructed by selecting the completion with the highest overall score as the "chosen" response and randomly selecting one of the remaining completions as the "rejected" response, creating clear preference pairs suitable for reward modeling and direct preference optimization.
 This dataset structure aligns well with our experimental setup, providing a robust foundation for evaluating different preference learning approaches.
 The dataset’s diverse prompt sources and evaluation criteria make it particularly valuable for training and evaluating reward models in a real-world context.
-To further tailor the dataset to our experimental setup, we organize the dataset as follows for the training and deployment stages:
-  • Training stage with passive data collection: We randomly sample $T=30,000$ data points from the UltraFeedback-binarized dataset’s train_prefs split for training.
+To further tailor the dataset to our experimental setup, we organize the dataset as follows:
+  • Online RLHF with passive data collection: We randomly sample $T=30,000$ data points from the UltraFeedback-binarized dataset’s train_prefs split for training.
 Each data point consists of a prompt and two responses with a binary preference label indicating the preferred response.
 We use the test_prefs split for evaluation.
-  • Training stage with active data collection: We allow the method to actively select 6,400 samples from the train_prefs split according to different selection strategies.
+  • Online RLHF with active data collection: We allow the method to actively select 6,400 samples from the train_prefs split according to different selection strategies.
 The global batch size is set to 8 for training.
 The selection is performed iteratively, where in each iteration, the method selects the most informative samples based on its selection criterion.
-  • Deployment stage: We use a pre-processed online variant of the UltraFeedback-binarized dataset from the test_gen split.
+  • Deployment-time adaption: We use a pre-processed online variant of the UltraFeedback-binarized dataset from the test_gen split.
 The dataset is divided into 20 sequential chunks to simulate an online deployment scenario.
@@
 G.2 Validating the Magnitude of $\kappa$
-We validate the magnitude of $\kappa$ in Eq.
-(2) by computing its value during the training process.
-The results show that $\kappa=171.62\pm 85.49$ during our training process, which is relatively large.
+We validate the magnitude of $\kappa$ by computing its value during the training process. The results show that $\kappa=171.62\pm 85.49$ during our training process, which is relatively large.
 G.3 Combined with Adam Optimizer
-(a) training loss
-(b) training accuracy
-(c) evaluation loss
-(d) evaluation accuracy
-Figure 4: For the training stage with passive data collection, we compare our proposed method and MLE [Zhu et al., 2023] in with passive data collection combined with Adam.
-We report the average accuracy and loss of the reward model during the training process.
 In previous experiments, we used SGD to update model parameters.
@@
 G.4 Comparison with DPO
-Figure 5: Comparison with DPO in the deployment stage.
-We also compare with DPO (Direct Preference Optimization) [Rafailov et al., 2023] in the deployment stage, as a reward-free method, DPO optimizes the policy directly using preference feedback without explicit reward modeling.
+We also compare with DPO [Rafailov et al., 2023] in the deployment stage, as a reward-free method, DPO optimizes the policy directly using preference feedback without explicit reward modeling.
 To ensure a fair comparison, we initialize the policy with 400 samples and use the same dataset settings as PPO to iteratively update the policy model using the DPO algorithm.
 The results are illustrated in Figure 11.
-While DPO outperforms the random baseline (Rand-MLE), it achieves lower cumulative rewards than the PromisingSet-based method.
+While DPO outperforms the random baseline (Rand-MLE), it achieves lower cumulative rewards than the methods using our action selection.
 This result suggests that DPO’s online learning capability remains a challenge.
 In contrast, the reward model learned by our selection strategy effectively learned streaming data and continuously updates the policy as new data arrive, indicating that in our deployment stage, a reward model with PPO may be a more suitable choice for sequentially learning from new data.
+(a) training loss
+(b) training accuracy
+(c) evaluation loss
+(d) evaluation accuracy
+Figure 9: For online RLHF with passive data collection, we compare our proposed method and MLE [Zhu et al., 2023] in with passive data collection combined with Adam.
+We report the average accuracy and loss of the reward model during the training process.
+Figure 10: Comparison of DPO and our method in deployment-time adaptation.
+Figure 11: Comparison of different methods for full update in deployment-time adaptation.
+G.5 Full Update of Reward Model
+Figure 11 shows deployment-time adaptation results using the meta-llama/Llama-3.2-1B model, where we update all parameters of the reward model instead of only the final layer.
+Both our method and MLE use the same action selection strategy.
+Our approach achieves comparable performance with MLE, indicating that our OMD-based method is still compatible with full-model updates.

### v3

@@
 Reinforcement Learning from Human Feedback (RLHF) has shown remarkable success in aligning Large Language Models (LLMs) with human preferences.
-Traditional RLHF approaches rely on a fixed dataset, which often suffers from limited coverage.
-To this end, online RLHF has emerged as a promising direction, enabling iterative data collection and model improvement.
-Despite its potential, this paradigm faces a key bottleneck: the requirement to continuously integrate new data into the historical dataset and re-optimize the model from scratch at each iteration, resulting in computational and storage costs that grow linearly with the number of iterations.
-In this work, we address this challenge by proposing a one-pass reward modeling method that does not require storing the historical data and can be computed in constant time.
-Specifically, we first formalize RLHF as a contextual preference bandit problem and design an online mirror descent algorithm with a tailored local norm to replace the standard maximum likelihood estimation for reward modeling.
-We then apply our method to various online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation.
-We provide theoretical guarantees showing that our method improves both statistical and computational efficiency.
-Finally, we provide practical algorithms and conduct experiments using Llama-3-8B-Instruct and Qwen2.5-7B-Instruct models on the Ultrafeedback-binarized and Mixture2 datasets, validating the effectiveness of our proposed method.
+Traditional RLHF methods rely on a fixed dataset, which often suffers from limited coverage.
+To this end, online RLHF has emerged as a promising direction, enabling iterative data collection and refinement.
+Despite its potential, this paradigm faces a key bottleneck: the requirement to continuously integrate new data into the dataset and re-optimize the model from scratch at each iteration, resulting in computational and storage costs that grow linearly with the number of iterations.
+In this work, we address this challenge by proposing a one-pass reward modeling method that eliminates the need to store historical data and achieves constant-time updates per iteration.
+Specifically, we first formalize RLHF as a contextual preference bandit and develop a new algorithm based on online mirror descent with a tailored local norm, replacing the standard maximum likelihood estimation for reward modeling.
+We then apply it to various online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation.
+We provide theoretical guarantees showing that our method enhances both statistical and computational efficiency.
+Finally, we design practical algorithms for LLMs and conduct experiments with the Llama-3-8B-Instruct and Qwen2.5-7B-Instruct models on Ultrafeedback and Mixture2 datasets, validating the effectiveness of our approach.
 1 Introduction
@@
 Traditional RLHF methods primarily rely on fixed preference datasets, which typically suffer from limited coverage.
-As a result, the learned reward models struggle to generalize to out-of-distribution (OOD) samples, constraining the effectiveness of the aligned models.
+As a result, the learned reward models struggle to generalize to out-of-distribution samples, constraining the effectiveness of the aligned models.
 To address this, online RLHF has emerged as a promising paradigm, enabling iterative data collection and model improvement.
-The general process can be described as (i) collect the preference data; (ii) update the model using the collected preference data.
+The general process can be described as (i) collect the preference data; (ii) update the model using the collected data.
 The above two steps are repeated for several iterations to boost model performance.
-In practice, the Claude (Bai et al., 2022) and LLaMA-2 (Touvron et al., 2023) projects have demonstrated that online RLHF can significantly enhance model performance.
-Despite its empirical success, online RLHF introduces significant computational challenges.
-Specifically, the typical process of online RLHF involves continuously integrating newly collected data into the dataset and re-optimizing the model from scratch over the expanded dataset.
+In practice, Claude (Bai et al., 2022) and LLaMA-2 (Touvron et al., 2023) have demonstrated that online RLHF can significantly enhance model performance (Dong et al., 2024).
+Theoretically, recent works (Xie et al., 2025; Cen et al., 2025) indicate that online exploration can improve the statistical efficiency of RLHF.
+Beyond performance gains, online RLHF serves as a crucial step toward agentic applications, where models can continuously integrate environmental feedback to enable real-time interaction, adaptive reasoning, and autonomous decision-making (Silver and Sutton, 2025).
+Despite its empirical success, online RLHF may introduces significant computational challenges.
+Specifically, the typical process of online RLHF involves continuously integrating newly collected data into the historical dataset and re-optimizing the model from scratch over the expanded dataset.
 While this strategy is statistically efficient, its computational and storage costs scale linearly with the number of iterations, which becomes prohibitive in long-term iterations, especially on edge devices where computation and memory resources are inherently limited.
@@
 Can we design online RLHF algorithms that are both statistically and computationally efficient?
-[…]
+Table 1: Comparison between previous works and our work in terms of the statistical and computational efficiency across different online RLHF settings.
+The column “Context” and “Action” represent the context and action are determined by the environment () or the algorithm ().
+For the computational efficiency (time and storage), we highlight the dependence on the $t$ at iteration $t$.
+Here, $d$ is the feature dimension, $T$ is the total number of iterations, $\kappa$ is the non-linearity coefficient, $\Phi=\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]$ is the concentrability vector, $V_{T}$ and $\mathcal{H}_{T}$ are two local norms satisfying $\lVert\Phi\rVert_{\mathcal{H}_{T}^{-1}}\leq\sqrt{\kappa}\lVert\Phi\rVert_{V_{T}^{-1}}$ (*: amortized complexity over $T$).
 Setting Context Action Gap/Regret Time Storage Reference
-[…]
+Passive $\widetilde{\mathcal{O}}\big(\sqrt{d}\cdot\kappa\lVert\Phi\rVert_{V_{T}^{-1}}\big)$ $\mathcal{O}(\log T)^{*}$ $\mathcal{O}(t)$ Zhu et al. (2023)
 $\widetilde{\mathcal{O}}\big(\sqrt{d}\cdot\lVert\Phi\rVert_{\mathcal{H}_{T}^{-1}}\big)$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ Ours (Theorem 1)
-[…]
+Active $\widetilde{\mathcal{O}}\big(d\sqrt{\kappa/T}\big)$ $\mathcal{O}(t\log t)$ $\mathcal{O}(t)$ Das et al. (2025)
 $\widetilde{\mathcal{O}}\big(d\sqrt{\kappa/T}\big)$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ Ours (Theorem 2)
-[…]
+Deploy $\widetilde{\mathcal{O}}\big(d\kappa\sqrt{T}\big)$ $\mathcal{O}(t\log t)$ $\mathcal{O}(t)$ Saha et al. (2023)
 $\widetilde{\mathcal{O}}\big(d\sqrt{\kappa T}\big)$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ Ours (Theorem 3)
 In this work, we provide an affirmative answer to this question in the setting of contextual preference bandits with linearly parameterized reward functions.
-Specifically, building on recent theoretical advancements in RLHF (Zhu et al., 2023; Das et al., 2024; Ji et al., 2024), we formulate the RLHF problem as a contextual dueling bandit problem (Yue et al., 2012; Saha, 2021).
+Specifically, building on recent theoretical advancements in RLHF (Zhu et al., 2023; Das et al., 2025; Ji et al., 2025), we formulate the RLHF problem as a contextual dueling bandit problem (Yue et al., 2012; Saha, 2021).
 While prior work has explored this formulation, most existing methods focus on statistical efficiency and overlook the growing computational burden.
-To bridge this gap, we introduce a novel one-pass reward modeling algorithm which does not require storing the historical data and can be computed in constant time.
-Our method is based on the online mirror descent framework with a tailored local norm that captures second-order information.
+To bridge this gap, inspired by recent progress in bandit learning (Zhang and Sugiyama, 2023; Li et al., 2024), we introduce a novel one-pass reward modeling method based on the online mirror descent framework with a tailored local norm that captures second-order information.
+Unlike traditional approaches, our method eliminates the need to store historical data and achieves constant-time updates per iteration, i.e., the computational cost remains invariant with respect to the cumulative number of iterations.
 We then apply our method to several online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation.
 We establish theoretical guarantees showing that our method improves both statistical and computational efficiency.
-Table 1 summarizes the comparison of our method with the existing methods.
+Table 1 summarizes the comparison of our method with the existing works.
 To enable usage in LLMs, we develop practical variants of our method.
@@
 Additionally, we employ rejection sampling to approximate model uncertainty in a computationally efficient manner.
-With the above techniques, we conduct experiments using the LLaMA-3-8B-Instruct (Llama Team, 2023) and Qwen2.5-7B-Instruct (Qwen Team, 2024) models on the Ultrafeedback-binarized (Cui et al., 2023) and Mixture2 (Dong et al., 2024) datasets, validating the effectiveness of our method.
+With the above techniques, we conduct experiments using the LLaMA-3-8B-Instruct (Llama Team, 2023) and Qwen2.5-7B-Instruct (Qwen Team, 2024) models on the Ultrafeedback (Cui et al., 2024) and Mixture2 (Dong et al., 2024) datasets.
+Experimental results validate the effectiveness of our method.
 To summarize, our contributions are as follows:
@@
   • We apply our method to a broad range of online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation.
-For each setting, we design tailored algorithms and establish corresponding theoretical guarantees, demonstrating that our approach achieves improved statistical and computational efficiency than existing methods.
+For each setting, we design tailored algorithms and establish corresponding theoretical guarantees, demonstrating that our approach achieves improved statistical and computational efficiency compared to existing methods.
   • We develop practical algorithms by approximating the update using Hessian-Vector Products combined with conjugate gradient descent, and estimating uncertainty via rejection sampling.
-Based on the above techniques, we conduct empirical evaluations using LLaMA-3-8B-Instruct and Qwen2.5-7B-Instruct models on Ultrafeedback-binarized and Mixture2 datasets, showing our method achieves better performance in both statistical and computational efficiency.
+Based on the above techniques, we conduct empirical evaluations using the LLaMA-3-8B-Instruct and Qwen2.5-7B-Instruct models on the Ultrafeedback and Mixture2 datasets, showing that our method improves both statistical and computational efficiency compared to existing methods.
 Organization.
-The rest of the paper is organized as follows: Section 3 introduces the problem setup.
-Section 4 presents our proposed one-pass reward modeling method.
-Section 5 applies our method to different online RLHF settings.
-Section 6 provides the practical implementations.
+Section 2 reviews the related work.
+Section 3 introduces the problem setup.
+Section 4 presents our proposed one-pass reward modeling method and section 5 applies it to various online RLHF settings.
+Section 6 provides practical versions of our method.
 Section 7 presents experimental results.
 Section 8 concludes the paper.
-We defer the proofs, experiment details, and more experimental results to the appendix.
+The proofs and experiment details are deferred to the appendix.
 2 Related Work
-In this section, we review the works most closely related to ours, including offline RLHF, online RLHF, contextual dueling bandits, and active learning.
-Offline RLHF.
-Learning from human preferences with deep learning models dates back to Christiano et al.
-(2017) and has been recently popularized by the success of large language models (Bai et al., 2022; Touvron et al., 2023; Achiam et al., 2023).
-Existing results generally fall into two categories: reward-based and reward-free.
-Reward-based methods typically consist of two steps: reward modeling and reward maximization (Ouyang et al., 2022).
-Given preference data, the reward model is trained to predict preference labels, and the LLM is fine-tuned based on the reward model using RL algorithms such as PPO (Schulman et al., 2017).
-In contrast, reward-free methods directly optimize the LLM using human preference feedback, without relying on explicit reward modeling (Rafailov et al., 2023; Azar et al., 2024).
-Among these works, Direct Preference Optimization (DPO) (Rafailov et al., 2023) is one of the most popular reward-free methods, which treats generative models directly as reward models and optimizes them using human preference feedback.
-Given the empirical success of RLHF, recent efforts have been devoted to developing a deeper theoretical understanding of this approach.
-Zhu et al.
-(2023) proposed to formulate RLHF as a contextual preference bandit problem and proved the maximum likelihood estimator converges under both the Bradley-Terry (BT) model and Plackett-Luce (PL) model with linear reward functions.
+In this section, we review the works most closely related to ours, including online RLHF, contextual dueling bandits, and active learning.
 Online RLHF.
-Offline RLHF methods rely on fixed preference datasets, which often suffer from limited data coverage.
+Traditional RLHF methods predominantly rely on fixed datasets, which often suffer from limited data coverage.
 Consequently, the resulting reward models struggle to generalize to out-of-distribution samples, thereby limiting the effectiveness of the aligned models.
@@
 (2024) identified key challenges in offline RLHF and theoretically demonstrated the potential benefits of online exploration.
-Recent work has incorporated optimism-driven bonus terms into the DPO loss to encourage exploration in online RLHF (Xie et al., 2025; Cen et al., 2025; Zhang et al., 2025).
-These studies primarily focus on the sample efficiency, but do not consider the accompanying increase in computational complexity.
-In this paper, we propose algorithms that achieve both statistical and computational efficiency.
+Recent work has incorporated optimism-driven bonus terms into the objective to encourage exploration in online RLHF (Xie et al., 2025; Cen et al., 2025; Zhang et al., 2025; Zhao et al., 2025).
+These approaches primarily focus on the sample efficiency, but do not consider the accompanying increase in computational complexity.
+To improve computational efficiency, Foster et al.
+(2025) tackled the challenge of enumerating an exponentially large response space.
+Differently, our work focuses on alleviating the computational burden that scales linearly with the number of iterations in online RLHF.
 Contextual Dueling Bandits and RL.
-Dueling bandits are a variant of the multi-armed bandit problem in which the learner sequentially selects a pair of arms and receives noisy binary feedback (Yue et al., 2012).
+Dueling bandits are a variant of the multi-armed bandit problem in which the learner sequentially selects a pair of arms and receives binary feedback (Yue et al., 2012).
 The contextual dueling bandit framework extends this setting by incorporating contextual information (Dudík et al., 2015; Saha, 2021; Bengs et al., 2022).
@@
 Additionally, Sekhari et al.
-(2024) investigated the contextual dueling bandit problem under an active learning paradigm, where the learner adaptively queries for feedback, aiming to minimize both regret and the number of queries.
+(2023) investigated the contextual dueling bandit problem under an active learning paradigm, where the learner adaptively queries to minimize both regret and the number of queries.
 To move beyond linear reward functions, Verma et al.
-(2025) introduced the neural dueling bandit problem, modeling the reward function using neural networks.
+(2025a) introduced the neural dueling bandit problem, modeling the reward function using neural networks.
 These prior works commonly rely on maximum likelihood estimation to learn the reward function, leading to computational complexity that grows linearly with the number of iterations.
-In contrast, we propose algorithms that maintain constant per-iteration computational complexity, while preserving statistical efficiency.
+In contrast, we propose algorithms that maintain constant per-iteration computational complexity.
 Active Learning.
@@
 Within the context of RLHF, Das et al.
-(2024) studied pool-based active learning, while Ji et al.
+(2025) and Verma et al.
+(2025b) studied pool-based active learning, while Ji et al.
 (2025) focused on the stream-based setting.
@@
 3 Problem Setup
-Following recent advancements in RLHF (Zhu et al., 2023; Das et al., 2024; Xiong et al., 2024), we formulate RLHF as a contextual bandit problem.
+Following recent advancements in RLHF (Zhu et al., 2023; Das et al., 2025; Xiong et al., 2024), we formulate RLHF as a contextual bandit problem.
 Specifically, we have a set of contexts $\mathcal{X}$ and a set of possible actions $\mathcal{A}$ per context.
@@
 Definition 1 (Bradley-Terry Model).
-Given a context $x\in\mathcal{X}$ and two actions $a,a^{\prime}\in\mathcal{A}$, the probability of the human preferring action $a$ over action $a^{\prime}$ is given by
-$\mathbb{P}\left[y=1\mid x,a,a^{\prime}\right]=\frac{\exp\left(r(x,a)\right)}{\exp\left(r(x,a)\right)+\exp\left(r\left(x,a^{\prime}\right)\right)},$ (1)
-where $r:\mathcal{X}\times\mathcal{A}\to\mathbb{R}$ is a latent reward function.
-We consider the widely used linear reward in the literature (Zhu et al., 2023; Cen et al., 2025).
+Given a context $x\in\mathcal{X}$ and two actions $a,a^{\prime}\in\mathcal{A}$, the probability of the human preferring action $a$ over action $a^{\prime}$ is given by $\mathbb{P}\left[y=1\mid x,a,a^{\prime}\right]=\frac{\exp\left(r(x,a)\right)}{\exp\left(r(x,a)\right)+\exp\left(r\left(x,a^{\prime}\right)\right)}$, where $r:\mathcal{X}\times\mathcal{A}\to\mathbb{R}$ is a latent reward function.
+To facilitate theoretical analysis, following prior works (Zhu et al., 2023; Cen et al., 2025), we consider the linear realizable setting, where the reward function is parameterized by a linear model.
 Assumption 1.
@@
 Remark 1.
-The feature mapping $\phi$ can be constructed by removing the last layer of a pre-trained large language model, and $\theta^{*}$ corresponds to the weights of the last layer.
+While this setting is a simplification of the real-world problem, it serves as a useful and analytically tractable starting point.
+Specifically, the feature mapping $\phi$ can be obtained by removing the final layer of a pre-trained large language model, with $\theta^{*}$ corresponding to the weights of that layer.
+Moreover, this assumption can be further relaxed by allowing model misspecification (Jin et al., 2020) and neural function approximation (Du et al., 2024; Verma et al., 2025b).
 Then, we can rewrite the probability as $\mathbb{P}\left[y=1\mid x,a,a^{\prime}\right]=\sigma(\phi(x,a)^{\top}\theta^{*}-\phi(x,a^{\prime})^{\top}\theta^{*})$, where $\sigma(w)=\frac{1}{1+\exp(-w)}$.
@@
 Definition 2.
-The non-linearity coefficient $\kappa$ is defined as
-$\kappa=\max_{x\in\mathcal{X},a,a^{\prime}\in\mathcal{A}}\max_{\theta\in\Theta}\frac{1}{\dot{\sigma}\left(\phi(x,a)^{\top}\theta-\phi(x,a^{\prime})^{\top}\theta\right)},$ (2)
-where $\dot{\sigma}(w)=\sigma(w)(1-\sigma(w))$ is the derivative function.
-The quantity $\kappa$ captures the learning complexity that satisfies $\kappa\leq 2+\exp(2BL)+\exp(-2BL)$.
-It may be exceedingly large with exponential dependence on the scale of features and parameters.
+Let $\dot{\sigma}(w)=\sigma(w)(1-\sigma(w))$ be the derivative function of $\sigma$, the non-linearity coefficient $\kappa$ is defined as $\kappa=\max_{x\in\mathcal{X},a,a^{\prime}\in\mathcal{A},\theta\in\Theta}\frac{1}{\dot{\sigma}(\phi(x,a)^{\top}\theta-\phi(x,a^{\prime})^{\top}\theta)}$.
+Intuitively, the quantity $\kappa$, defined as the inverse of the derivative, characterizes the learning difficulty of the reward function.
+In particular, a smaller derivative leads to a larger $\kappa$, implying that the model output changes less for the same input variation and thus the function is harder to learn.
+By direct calculation, we have $\kappa\leq 3+\exp(2BL)$.
+Therefore, $\kappa$ can be exceedingly large, exhibiting an exponential dependence on the magnitude of the features and the model parameters.
 4 Our Framework
@@
 $\bar{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big\{\ell_{t}(\theta)+\frac{1}{2\eta}\left\|\theta-\bar{\theta}_{t}\right\|_{\bar{\mathcal{H}}_{t}}^{2}\Big\},$ (2)
-where ¯$\bar{\mathcal{H}}_{t}=\sum_{i=1}^{t-1}H_{i}\left(\bar{\theta}_{i+1}\right)+\lambda I$ is the local norm, and $\eta$ is the step size.
+where $\bar{\mathcal{H}}_{t}=\sum_{i=1}^{t-1}H_{i}(\bar{\theta}_{i+1})+\lambda I$ is the local norm, and $\eta$ is the step size.
 The optimization problem can be decomposed into two terms.
@@
 The second consists of a quadratic proxy for the past losses constructed through the sequence $\{\bar{\theta}_{i}\}_{i\leq t}$.
-A key component is the design of the local norm ¯$\bar{\mathcal{H}}_{t}$, which approximates the Hessian matrix by $H_{i}\left(\bar{\theta}_{i+1}\right)$ at a lookahead point $\bar{\theta}_{i+1}$.
+A key component is the design of the local norm $\bar{\mathcal{H}}_{t}$, which approximates the Hessian matrix by $H_{i}(\bar{\theta}_{i+1})$ at a lookahead point $\bar{\theta}_{i+1}$.
 Such a Hessian matrix effectively captures local information and is crucial for ensuring statistical efficiency.
@@
 To this end, we propose to approximate the current loss $\ell_{t}(\theta)$ using a second-order Taylor expansion, drawing inspiration from Zhang and Sugiyama (2023).
-Define the second-order approximation of $\ell_{t}(\theta)$ as ~$\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+g_{t}(\widetilde{\theta}_{t})^{\top}(\theta-\widetilde{\theta}_{t})+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}$.
+Define the second-order approximation of $\ell_{t}(\theta)$ as $\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+g_{t}(\widetilde{\theta}_{t})^{\top}(\theta-\widetilde{\theta}_{t})+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}$.
 Then, we replace the loss $\ell_{t}(\theta)$ in Eq.
-(4) with the approximation ~$\widetilde{\ell}_{t}(\theta)$, leading to the update rule:
+(2) with the approximation $\widetilde{\ell}_{t}(\theta)$, leading to the update rule:
 $\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big\{\big\langle g_{t}(\widetilde{\theta}_{t}),\theta\big\rangle+\frac{1}{2\eta}\big\|\theta-\widetilde{\theta}_{t}\big\|_{\widetilde{\mathcal{H}}_{t}}^{2}\Big\},$ (3)
-where $\eta$ is the step size and ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$ is the local norm with $\mathcal{H}_{t}\triangleq\sum_{i=1}^{t-1}H_{i}(\widetilde{\theta}_{i+1})+\lambda I$.
+where $\eta$ is the step size and $\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$ is the local norm with $\mathcal{H}_{t}\triangleq\sum_{i=1}^{t-1}H_{i}(\widetilde{\theta}_{i+1})+\lambda I$.
 Eq.
@@
 Thus, the estimator $\widetilde{\theta}_{t+1}$ provides a closed-form solution, leading to a $\mathcal{O}(1)$ computational complexity per iteration.
-Since the estimator processes the samples in a one-pass manner, it mitigates the memory burden associated with computing the gradient of the full dataset.
+Furthermore, since the estimator processes the samples in a one-pass manner, it mitigates the memory burden associated with computing the gradient of the full dataset.
 These properties make the method particularly suitable for edge devices, where both memory and computational resources are severely constrained.
@@
 1: Define the loss function $\ell_{t}(\theta)$ as Eq. (1)
-2:  Update ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$
+2: Update $\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$
 3: Compute $\widetilde{\theta}_{t+1}^{\prime}=\widetilde{\theta}_{t}-\eta\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t})$
@@
 5: Update $\mathcal{H}_{t+1}=\mathcal{H}_{t}+H_{t}(\widetilde{\theta}_{t+1})$
-5:  $\widetilde{\theta}_{t+1}$
+[…]
 Algorithm 2 Passive Data Collection
-0:  Regularization $\lambda$, step size $\eta$
-1:  Initialize $\widetilde{\theta}_{1}=\bm{0}$ and ~$\widetilde{\mathcal{H}}_{1}=\lambda I$
+0: Regularization parameter $\lambda$, step size $\eta$
+1: Initialize $\widetilde{\theta}_{1}=\bm{0}$ and $\widetilde{\mathcal{H}}_{1}=\lambda I$
 2:for $t=1,2,\ldots,T$ do
@@
 6: Construct $\widetilde{J}_{T+1}(\pi)$ as in Eq. (4)
-6:  $\pi_{T+1}=\operatorname*{arg\,max}_{\pi\in\Pi}\widetilde{J}_{T+1}(\pi)$
+[…]
 4.3 Theoretical guarantee
 Note that the update rule in Eq. (3) is a special case of online mirror descent, specifically:
-$\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\big{\langle}g_{t}(\widetilde{\theta}_{t}),\theta\big{\rangle}+\frac{1}{\eta}\mathcal{D}_{\psi_{t}}(\theta,\widetilde{\theta}_{t})\Big{\}},$
+$\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big\{\eta\big\langle g_{t}(\widetilde{\theta}_{t}),\theta\big\rangle+\mathcal{D}_{\psi_{t}}(\theta,\widetilde{\theta}_{t})\Big\},$
 where $\psi_{t}(\theta)=\frac{1}{2}\lVert\theta\rVert_{\widetilde{\mathcal{H}}_{t}}^{2}$ is the regularizer and $\mathcal{D}_{\psi_{t}}(\theta,\widetilde{\theta}_{t})=\psi_{t}(\theta)-\psi_{t}(\widetilde{\theta}_{t})-\langle\nabla\psi_{t}(\widetilde{\theta}_{t}),\theta-\widetilde{\theta}_{t}\rangle$ is Bregman divergence.
-Based on the analysis of online mirror descent, we have the following lemma.
+Leveraging the modern analysis of online mirror descent (Zhao et al., 2024; Zhang and Sugiyama, 2023), we derive the following estimation error bound.
 Lemma 1.
-Let $\delta\in(0,1]$, set $\eta=(1/2)\log 2+\left(BL+1\right)$ and $\lambda=84\sqrt{2}\eta(dL^{2}+BL^{3})$, define $\mathcal{C}_{t}=\{\theta\in\Theta\mid\|\theta-\widetilde{\theta}_{t}\|_{\mathcal{H}_{t}}\leq\widetilde{\beta}_{t}\triangleq\mathcal{O}\big{(}\sqrt{d}(\log(t/\delta))^{2}\big{)}\}$.
+Let $\delta\in(0,1]$, set $\eta=(1/2)\log 2+\left(BL+1\right)$ and $\lambda=84\sqrt{2}\eta(dL^{2}+BL^{3})$, define $\mathcal{C}_{t}=\{\theta\in\Theta\mid\|\theta-\widetilde{\theta}_{t}\|_{\mathcal{H}_{t}}\leq\widetilde{\beta}_{t}\triangleq\mathcal{O}\big(\sqrt{d}\log(t/\delta)\big)\}$.
 Then, we have $\operatorname{Pr}\left[\forall t\geqslant 1,\theta^{*}\in\mathcal{C}_{t}\right]\geqslant 1-\delta$.
@@
 For the MLE estimator in Eq.
-(3), prior works (Zhu et al., 2023; Das et al., 2024; Ji et al., 2024) have shown $\|\theta-\widetilde{\theta}_{t}\|_{V_{t}}\leq\widetilde{\mathcal{O}}(\kappa\sqrt{d})$, where $V_{t}=\sum_{i=1}^{t-1}z_{i}z_{i}^{\top}+\lambda I$.
-By the definition of $\mathcal{H}_{t}$, it holds that $\mathcal{H}_{t}\succeq\kappa^{-1}V_{t}$, Lemma 1 implies $\|\theta-\widetilde{\theta}_{t}\|_{V_{t}}\leq\sqrt{\kappa}\|\theta-\widetilde{\theta}_{t}\|_{\mathcal{H}_{t}}\leq\widetilde{\mathcal{O}}(\sqrt{\kappa d})$.
+(1), prior works (Zhu et al., 2023; Das et al., 2025; Ji et al., 2025) have shown $\|\theta-\widetilde{\theta}_{t}\|_{V_{t}}\leq\widetilde{\mathcal{O}}(\kappa\sqrt{d})$, where $V_{t}=\sum_{i=1}^{t-1}z_{i}z_{i}^{\top}+\lambda I$.
+By the definition of $\mathcal{H}_{t}$, it holds that $\mathcal{H}_{t}\succeq\kappa^{-1}V_{t}$, Lemma 1 implies that $\|\theta-\widetilde{\theta}_{t}\|_{V_{t}}\leq\sqrt{\kappa}\|\theta-\widetilde{\theta}_{t}\|_{\mathcal{H}_{t}}\leq\widetilde{\mathcal{O}}(\sqrt{\kappa d})$.
 This result shows that Lemma 1 improves upon previous bounds by at least a factor of $\sqrt{\kappa}$.
@@
 In this section, we apply our framework to three distinct RLHF scenarios, including online RLHF with passive data collection, active data collection, and deployment-time adaptation.
-(a) Passive Data Collection
-(b) Active Data Collection
-(c) Deployment-Time Adaptation
-[…]
 5.1 Online RLHF with passive data collection
@@
 The policy $\pi_{T+1}$ is selected as $\pi_{T+1}=\operatorname*{arg\,max}_{\pi\in\Pi}\widetilde{J}_{T+1}(\pi)$.
-The detailed procedure is present in Algorithm 2, and we show it enjoys the following guarantee:
+The detailed procedure is present in Algorithm 2, and we show it enjoys the following guarantee.
 Theorem 1.
@@
 Remark 2.
-The term $\big{\lVert}\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\big{\rVert}_{\mathcal{H}_{T+1}^{-1}}$ is referred to “concentrability coefficient” in the literature.
+The term $\lVert\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\rVert_{\mathcal{H}_{T+1}^{-1}}$ is usually referred to “concentrability coefficient” in the literature.
 It measures the distribution shift between the optimal policy and the collected data.
 Remark 3.
-Our result improves upon Zhu et al.
-(2023) in both statistical and computational efficiency.
-Specifically, since $\mathcal{H}_{t}\succeq\kappa^{-1}V_{t}$, Theorem 2 directly implies a $\widetilde{\mathcal{O}}(\sqrt{d\kappa}\cdot\lVert\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\rVert_{V_{T+1}^{-1}})$ guarantee, improving their $\widetilde{\mathcal{O}}(\sqrt{d}\kappa\cdot\lVert\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\rVert_{V_{T+1}^{-1}})$ result by a factor of $\sqrt{\kappa}$.
+For statistical efficiency, since $\mathcal{H}_{t}\succeq\kappa^{-1}V_{t}$, Theorem 1 improves the $\widetilde{\mathcal{O}}(\sqrt{d}\kappa\cdot\lVert\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\rVert_{V_{T+1}^{-1}})$ result of  Zhu et al.
+(2023) at least by a factor of $\sqrt{\kappa}$.
 Regarding computational efficiency, their algorithm has a total storage complexity of $\mathcal{O}(T)$ and a time complexity of $\mathcal{O}(T\log T)$, leading to an amortized per-iteration cost of $\mathcal{O}(\log T)$.
 In contrast, our algorithm maintains a strict $\mathcal{O}(1)$ complexity per iteration, offering a substantial computational advantage.
-Algorithm 3 Active Data Collection
-0:  Regularization $\lambda$, step size $\eta$
-1:  Initialize $\widetilde{\theta}_{1}=\bm{0}$ and $\mathcal{H}_{1}=\lambda I$
-2:  for $t=1,2,\ldots,T$ do
-3:     Choose $(x_{t},a_{t},a_{t}^{\prime})$ as Eq. (7), observe $y_{t}$
-[…]
-5:  end for
-6:  Set $\widetilde{r}_{T+1}(x,a)=\frac{1}{T+1}\sum_{t=1}^{T+1}\phi(x,a)^{\top}\widetilde{\theta}_{t}$
-6:  $\pi_{T+1}(x)=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{r}_{T+1}(x,a)$
-Algorithm 4 Deployment-Time Adaptation
-0:  Regularization $\lambda$, step size $\eta$
-1:  Initialize $\widetilde{\theta}_{1}=0$ and $\mathcal{H}_{1}=\lambda I$.
-2:  for $t=1,2,\ldots,T$ do
-3:     Observes the context $x_{t}$.
-4:     Selects $a_{t}$ and $a_{t}^{\prime}$ as Eq. (9) and Eq. (10)
-5:     Observe the preference feedback $y_{t}$
-[…]
-7:  end for
+(a) Passive Data Collection
+(b) Active Data Collection
+(c) Deployment-Time Adaptation
+Figure 1: Different settings of online RLHF. Contexts and actions selected by the environment () are shown in grey, while those selected by the algorithm () are highlighted in color.
 5.2 Online RLHF with active data collection
@@
 Active Data Collection.
-At each iteration, we obtain a new preference data $(x_{t},a_{t},a_{t}^{\prime},y_{t})$.
-We then update the reward model using our one-pass reward modeling method as in Eq.
+At each iteration, we select a triplet $(x_{t},a_{t},a^{\prime}_{t})$ to query for human feedback $y_{t}$, and then update the reward model using our one-pass method as defined in Eq.
 (3).
-To guide data acquisition, we actively select the query with the highest uncertainty under the current reward model.
-Specifically, we select the next query by solving:
+To guide data acquisition, we adopt an active selection strategy that queries the sample with the highest predictive uncertainty under the current reward model.
+Specifically, the next query is chosen by solving:
 $(x_{t+1},a_{t+1},a_{t+1}^{\prime})=\operatorname*{arg\,max}_{x,a,a^{\prime}\in\mathcal{X}\times\mathcal{A}\times\mathcal{A}}\left\{\big\lVert\phi(x,a)-\phi(x,a^{\prime})\big\rVert_{\mathcal{H}_{t+1}^{-1}}\right\}.$ (5)
@@
 In this scenario, the learner faces a dual objective: selecting actions that maximize rewards to ensure a positive user experience, while also choosing actions that yield informative feedback to facilitate continual model improvement.
-To this end, we choose the cumulative regret as the performance measure, defined as
-${\operatorname*{Reg}}_{T}=\sum_{t=1}^{T}\left(r\left(x_{t},\pi^{*}(x_{t})\right)-\frac{1}{2}\left(r\left(x_{t},a_{t}\right)+r\left(x_{t},a^{\prime}_{t}\right)\right)\right),$ (8)
-where $\pi^{*}$ is the optimal policy.
-This measure differs from the one introduced by Ji et al.
-(2024), which evaluates the gap between th e optimal action and a single selected action.
-Our measure is designed to ensure that both actions are sufficiently effective, thereby delivering a high-quality user experience.
-While better aligns with deployment needs, it also poses greater challenges.
-Specifically, to optimize the measure in Ji et al.
-(2024), the learner can exploit one high-reward action while using the other for uniform exploration.
-In contrast, our setting imposes a stricter requirement: both actions must be good, thereby limiting the ability to sacrifice one action purely for exploration.
+To this end, we consider the measure: ${\operatorname*{Reg}}_{T}=\sum_{t=1}^{T}\left(r\left(x_{t},\pi^{*}(x_{t})\right)-\frac{1}{2}\left(r\left(x_{t},a_{t}\right)+r\left(x_{t},a^{\prime}_{t}\right)\right)\right)$, where $\pi^{*}$ is the optimal policy.
 Action selection.
-At each iteration, given a new preference data $(x_{t},a_{t},a_{t}^{\prime},y_{t})$, the learner updates the reward model using our one-pass reward modeling method as in Eq.
-(5).
-The learner must select queries that are both informative and with high rewards.
-To address this, we choose the first action $a_{t+1}$ to maximize the estimated reward based on the estimator $\widetilde{\theta}_{t+1}$, i.e.,
+At each iteration, given a prompt $x_{t}$ from the user, the learner selects two actions $a_{t}$ and $a_{t}^{\prime}$ and obtain the feedback $y_{t}$.
+The learner must select actions that are both informative and with high rewards.
+To this end, we choose the first action $a_{t+1}$ to maximize the estimated reward, i.e.,
 $a_{t+1}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\phi(x_{t+1},a)^{\top}\widetilde{\theta}_{t+1}.$ (6)
@@
 Theorem 3.
-For any $\delta\in(0,1]$, set parameters as in Lemma 1, Algorithm 4 ensures with probability at least $1-\delta$, the regret satisfies
-${\operatorname*{Reg}}_{T}\leq\widetilde{\mathcal{O}}\big{(}d\sqrt{{\kappa}{T}}\big{)}.$
+For any $\delta\in(0,1]$, set parameters as in Lemma 1, Algorithm 4 ensures the regret satisfies ${\operatorname*{Reg}}_{T}\leq\widetilde{\mathcal{O}}\big(d\sqrt{{\kappa}{T}}\big)$ with probability at least $1-\delta$.
 Remark 5.
@@
 Computationally, our algorithm has an $\mathcal{O}(1)$ time and space complexity per round, while their MLE estimator needs $\mathcal{O}(t\log t)$ time and $\mathcal{O}(t)$ space complexity at iteration $t$ due to optimization over the historical data.
+Algorithm 3 Active Data Collection
+0: Regularization parameter $\lambda$, step size $\eta$
+1: Initialize $\widetilde{\theta}_{1}=\bm{0}$ and $\mathcal{H}_{1}=\lambda I$
+2:for $t=1,2,\ldots,T$ do
+3:  Choose $(x_{t},a_{t},a_{t}^{\prime})$ as Eq. (5), observe $y_{t}$
+[…]
+5:end for
+6: Set $\widetilde{r}_{T+1}(x,a)=\frac{1}{T+1}\sum_{t=1}^{T+1}\phi(x,a)^{\top}\widetilde{\theta}_{t}$
+[…]
+Algorithm 4 Deployment-Time Adaptation
+0: Regularization parameter $\lambda$, step size $\eta$
+1: Initialize $\widetilde{\theta}_{1}=0$ and $\mathcal{H}_{1}=\lambda I$.
+2:for $t=1,2,\ldots,T$ do
+3:  Observes the context $x_{t}$.
+4:  Selects $a_{t}$ and $a_{t}^{\prime}$ as Eq. (6) and Eq. (7)
+5:  Observe the preference feedback $y_{t}$
+[…]
+7:end for
 6 Practical Implementation
-In this section, we introduce the practical implementation of our proposed algorithm.
+While the proposed one-pass algorithm completely removes the need to store historical data and achieves constant-time updates per iteration, its computational cost still exhibits an implicit dependence on the feature dimension $d$, which can become non-negligible in large-scale model optimization.
+To further alleviate this issue, we introduce in this section several empirical approximation techniques designed to reduce the effective dependence on dimensionality and enhance practical efficiency.
 6.1 Computation of inverse Hessian
-The OMD update in Eq.
-(5) requires computation of the inverse of the Hessian matrix.
-Omitting the projection operation, Eq.
-(5) can be rewritten as $\widetilde{\theta}_{t+1}=\widetilde{\theta}_{t}-\eta\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t})$.
-Computing the full Hessian inverse ~$\widetilde{\mathcal{H}}_{t}^{-1}$ directly incurs a time complexity of $\mathcal{O}(d^{3})$, which is prohibitive for LLMs.
+Although the OMD update in Eq.
+(3) enjoys one-pass property, it requires the computation of matrix inversion.
+Specifically, by omitting the projection operation, Eq.
+(3) can be rewritten as $\widetilde{\theta}_{t+1}=\widetilde{\theta}_{t}-\eta\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t})$ where $\widetilde{\mathcal{H}}_{t}=\sum_{i=1}^{t-1}H_{i}(\widetilde{\theta}_{i+1})+\eta H_{t}(\widetilde{\theta}_{t})+\lambda I$.
+Computing the full $\widetilde{\mathcal{H}}_{t}^{-1}$ directly incurs a time complexity of $\mathcal{O}(d^{3})$, which is prohibitive for LLMs as $d$ is typically large.
 This cost can be reduced to $\mathcal{O}(d^{2})$ by applying the Sherman-Morrison-Woodbury formula, leveraging the fact that the Hessian is a rank-one update.
 Specifically, for a matrix of the form $A+\mathbf{x}\mathbf{x}^{\top}$ where $A$ is invertible and $\mathbf{x}$ is a vector, the inverse is given by $(A+\mathbf{x}\mathbf{x}^{\top})^{-1}=A^{-1}-\frac{A^{-1}\mathbf{x}\mathbf{x}^{\top}A^{-1}}{1+\mathbf{x}^{\top}A^{-1}\mathbf{x}}$, requiring only $\mathcal{O}(d^{2})$ time.
-Nevertheless, even this reduced complexity is costly for large models.
+Nevertheless, even this reduced complexity remains costly for large models.
 To further reduce the computational burden to $\mathcal{O}(d)$, we employ the Hessian-vector product technique combined with conjugate gradient descent (Boyd and Vandenberghe, 2004).
-Instead of explicitly computing ~$\widetilde{\mathcal{H}}_{t}^{-1}$, we solve the linear system ~$\widetilde{\mathcal{H}}_{t}v=g_{t}(\widetilde{\theta}_{t})$ iteratively, where $v$ is the solution we seek.
-The HVP operation ~$\widetilde{\mathcal{H}}_{t}v$ for any vector $v$ can be computed efficiently using twice differentiation: ~$\widetilde{\mathcal{H}}_{t}v=\nabla_{\theta}\big{(}\nabla_{\theta}\mathcal{L}_{t}(\theta)^{\top}v\big{)}\big{|}_{\theta=\widetilde{\theta}_{t}}+\lambda_{t}v$, where $\mathcal{L}_{t}(\theta)$ is the function whose Hessian matrix at $\widetilde{\theta}_{t}$ is ~$\widetilde{\mathcal{H}}_{t}$.
-By leveraging this efficient computation, the conjugate gradient method iteratively refines the solution.
-In practice, finding $\mathcal{L}_{t}(\theta)$ is challenging as ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$ involve complex structure.
-A key insight is ~$\widetilde{\mathcal{H}}_{t}$ can be viewed as the cumulative Hessian of the $\ell_{t}(\theta)$ approximately.
-This motivates the use of a more tractable surrogate $\ell_{t}(\theta)$ in Eq.
-(3) while keeping the update.
-To further approximate the effect of curvature accumulation, we adopt an adaptive damping scheme that approximately preserves historical Hessian information while incorporating new Hessian updates: $\lambda_{t}=\lambda_{0}\cdot\min\{1,f(t/T)\}$, where $f(\cdot)$ is a monotonic increasing function, such as linear functions.
+Instead of explicitly computing $\widetilde{\mathcal{H}}_{t}^{-1}$, we define $v_{t}=\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t})$ and solve the linear system $\widetilde{\mathcal{H}}_{t}v_{t}=g_{t}(\widetilde{\theta}_{t})$ using the conjugate gradient method.
+The required matrix–vector product decomposes as $\widetilde{\mathcal{H}}_{t}v_{t}={\sum\nolimits_{i=1}^{t-1}H_{i}(\widetilde{\theta}_{i+1})v_{t}}+{\lambda v_{t}}+{\eta H_{t}(\widetilde{\theta}_{t})v_{t}}$.
+For the first term, materializing and storing all past Hessians $H_{i}(\widetilde{\theta}_{i+1})$ is infeasible.
+We therefore absorb their effect into the second term by replacing $\lambda$ with $\lambda_{t}=\lambda_{0}\cdot\min\{1,f(t/T)\}$, where $f(\cdot)$ is a monotonic increasing function, such as a linear or logarithmic function.
+The last term can be computed via the Pearlmutter trick as $H_{t}(\widetilde{\theta}_{t})v_{t}=\nabla_{\theta}\big(\nabla_{\theta}\ell_{t}(\theta)^{\top}v_{t}\big)\big|_{\theta=\widetilde{\theta}_{t}}$.
+Each iteration therefore requires only HVPs and vector operations, yielding an overall $\mathcal{O}(d)$ per-iteration cost with a small fixed number of iterations.
 6.2 Computation of model uncertainty
 In both online RLHF with active data collection and deployment-time adaptation, our algorithm utilizes uncertainty-driven query selection strategies.
-While quantifying uncertainty using the local norm induced by the Hessian matrix offers strong theoretical guarantees, it is computationally prohibitive in practice.
-To address this challenge, we adopt a rejection sampling-based approximation, a technique commonly employed for exploration in the literature (Nakano et al., 2021; Gulcehre et al., 2023; Dong et al., 2023, 2024).
-Specifically, given a prompt, we sample $n$ independent responses by the current model, then use the trained reward model to rank the responses.
+While quantifying uncertainty using the local norm induced by the inverse Hessian matrix offers strong theoretical guarantees, it is computationally prohibitive in practice.
+To address this challenge, we adopt a rejection sampling-based approximation, a technique commonly employed for exploration in the RLHF literature (Nakano et al., 2021; Gulcehre et al., 2023; Dong et al., 2023, 2024).
+Specifically, given a prompt, we sample $n$ independent responses by the current model, then use the trained reward function to rank the responses.
 Then, we use different strategies to select the response for different settings.
 Specifically, In active data collection, the key insight is to identify and query samples that exhibit the greatest diversity in prompt action features.
-To this end, we select the response with the highest predicted reward and the response with the lowest predicted reward.
-In deployment-time adaptation, the core idea is to select the first action to maximize the estimated reward, while the second is chosen to balance high reward with sufficient divergence from the first.
-Concretely, we select the response with the highest predicted reward and another from the top-$1/q$ percentile of the reward, where $q$ is a hyperparameter.
+To this end, we select the response with the highest predicted reward and the one with the lowest predicted reward.
+In deployment-time adaptation, the core idea is to select the first arm to maximize the estimated reward, while the second is chosen to balance high reward with sufficient divergence from the first.
+Concretely, we select the response with the highest predicted reward and another from the top-$1/q$ percentile of the reward to ensure diversity, where $q$ is a hyperparameter.
 7 Experiments
-In this section, we empirically evaluate the performance of our proposed method.  [^1] Equal contribution. We first describe the experimental setup, and then present the empirical results.
-7.1 Experiment setup
-In our experiments, we employ the Llama-3-8B-Instruct 222huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct and Qwen2.5-7B-Instruct 333huggingface.co/Qwen/Qwen2.5-7B-Instruct as the foundation model for reward model.
-We extract features $\phi(x,a)$ using the last layer of the model, and the dimension is $d=4096$.
-We use two datasets for evaluation.
-The first one is Ultrafeedback-binarized dataset 444huggingface.co/datasets/OpenRLHF/preference_dataset_mixture2_and_safe_pku, a pre-processed version of the original Ultrafeedback dataset (Cui et al., 2023), a widely used benchmark for RLHF.
-It collects about $64,000$ prompts from diverse resources, including question answering, summarization, and dialogue generation.
-Each data consists of a context $x$, two responses $a$ and $a^{\prime}$, and a preference label $y$.
-We also employ a mixed dataset, Mixture2 dataset 444huggingface.co/datasets/OpenRLHF/preference_dataset_mixture2_and_safe_pku, which combines a variety of preference datasets, including HH-RLHF, SHP, UltraFeedback, etc.
-11footnotetext: The code is available at https://github.com/ZinYY/Online_RLHF
 (a) training loss
@@
 (d) evaluation accuracy
-Figure 2: For Llama-3-8B-Instruct on Ultrafeedback-binarized dataset with passive data collection, we report the comparison of MLE and our OMD-based method for the reward model about (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy.
-(a) training loss
-(b) training accuracy
-(c) evaluation loss
-(d) evaluation accuracy
-Figure 3: For Qwen2.5-7B-Instruct on Ultrafeedback-binarized dataset with passive data collection, we report the comparison of MLE and our OMD-based method for the reward model about (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy.
-(a) training loss
-(b) training accuracy
-(c) evaluation loss
-(d) evaluation accuracy
-Figure 4: For Llama-3-8B-Instruct on Mixture2 dataset with passive data collection, we report the comparison of MLE and our OMD-based method for the reward model about (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy.
-7.2 Experimental results
-In this section, we present the experimental results, including passive data collection, active data collection and deployment-time adaptation.
-7.2.1 Passive Data Collection
-We randomly sample $T=30,000$ data points from the dataset for training the reward model.
-We evaluate both the standard MLE-based method, which uses a stochastic gradient descent (SGD) optimizer, and our proposed OMD-based method in terms of reward model loss and accuracy.
-Figures 23, and 4 present the results for three settings: (i) the Llama-3-8B-Instruct model on the Ultrafeedback-binarized dataset, (ii) the Qwen2.5-7B model on the same dataset, and (iii) the Llama-3-8B-Instruct model on the Mixture2 dataset.
-We report the results across four key metrics: (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy.
-Our method demonstrates faster convergence to a lower loss and achieves higher evaluation accuracy compared to the MLE baseline.
-These results highlight the superior statistical efficiency of the OMD-based approach, which achieves improved performance with fewer training samples.
+Figure 2: For online RLHF with passive data collection, we report the comparison of MLE and our method about (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy.
 (a) training loss
@@
 Method ACC (%) Time (s)
-Rand-MLE 69.51$\pm$0.5 4876$\pm$47
-Active-MLE 69.82$\pm$0.4 4982$\pm$52
-Rand-OMD 68.97$\pm$0.6 1456$\pm$31
-Ours 70.43$\pm$0.3 1489$\pm$36
+[…]
+[…]
+[…]
+[…]
 (c) evaluation accuracy and training time
-Figure 5: For Llama-3-8B-Instruct on Ultrafeedback-binarized dataset with active data collection, we report the comparison of different methods about (a) training loss, (b) evaluation accuracy and (c) final evaluation accuracy and training time.
-(a) training loss
-(b) evaluation accuracy
-Method ACC (%) Time (s)
-Rand-MLE 67.36 $\pm$ 0.5 4653 $\pm$ 121
-Active-MLE 67.25 $\pm$ 0.4 4701 $\pm$ 103
-Rand-OMD 66.80 $\pm$ 0.5 1312 $\pm$ 47 0
-Ours 67.11 $\pm$ 0.5 1325 $\pm$ 54 0
-(c) evaluation accuracy and training time
-Figure 6: For Qwen2.5-7B-Instruct on Ultrafeedback-binarized dataset with active data collection, we report the comparison of different methods about (a) training loss, (b) evaluation accuracy and (c) final evaluation accuracy and training time.
-7.2.2 Active Data Collection
-In this experimental setup, we constrain the algorithm to select only 6,400 samples from the full training dataset, based on different data selection strategies.
-We evaluate the performance of both the standard MLE-based method and our proposed OMD-based method under this limited data regime.
-To assess the effectiveness of the data selection strategy itself, we compare our approach against a random selection baseline.
-We report the results across four key metrics: (a) training loss curve, (b) evaluation accuracy curve and (c) final evaluation accuracy and training time.
-Figure 5 and Figure 6 demonstrate that our OMD-based method achieves competitive performance with the MLE-based method for both data collection strategies, while improving the training time by significantly.
-Moreover, our data selection strategy outperforms the random selection strategy, demonstrating that our method can effectively select informative data to improve the performance.
-7.2.3 Deployment-Time Adaptation
-To emulate real-world deployment scenarios, we partition the dataset into 20 sequential chunks and process them incrementally.
-This setup allows us to evaluate the adaptability and robustness of different action selection strategies over time.
-We benchmark our proposed selection strategy against three baselines: (i): random selection, (ii): the best and second best actions, and (iii): the best and worst actions.
-We combine the above strategies with MLE-based and OMD-based reward estimators.
-We report three key metrics: (a) the average cumulative reward for MLE-based methods, (b) the average cumulative reward for OMD-based methods, and (c) the win rate of OMD-based methods over MLE-based counterparts under identical selection strategies.
-As illustrated in Figure 7 and Figure 8, our action selection strategy consistently outperforms the baselines across both estimator types.
-These results highlight the strength of our approach in balancing the exploitation of high-reward responses with sufficient exploration to drive continual model improvement.
-Moreover, the win rate comparisons suggest that our OMD-based estimator shows competitive performance relative to its MLE-based counterpart, further validating its practical effectiveness.
-(a) Reward of MLE-based methods
-(b) Reward of OMD-based methods
+Figure 3: For online RLHF with active data collection, we report the comparison of different methods about (a) training loss, (b) evaluation accuracy and (c) final evaluation accuracy and training time.
+In this section, we empirically evaluate the performance of our proposed method.
+[^1] The code is available at https://github.com/ZinYY/Online_RLHF We first describe the experimental setup, and then present the empirical results.
+7.1 Experiment setup
+In our experiments, we employ the Llama-3-8B-Instruct and Qwen2.5-7B-Instruct as the base model for reward model.
+We extract features $\phi(x,a)$ using the last layer of the model, and the dimension is $d=4096$.
+We use two datasets for evaluation.
+The first one is Ultrafeedback-binarized dataset, a pre-processed version of the original Ultrafeedback dataset (Cui et al., 2024), a widely used benchmark for RLHF.
+It collects about $64,000$ prompts from diverse resources, including question answering, summarization, and dialogue generation.
+Each data consists of a context $x$, two responses $a$ and $a^{\prime}$, and a preference label $y$.
+We also employ a mixed dataset, Mixture2 dataset (Dong et al., 2024), which combines a variety of preference datasets, including HH-RLHF, SHP, UltraFeedback, Capybara, etc.
+The dataset follows the same format as the UltraFeedback-binarized dataset.
+7.2 Experimental results
+We present the experimental results for Llama-3-8B-Instruct on the Ultrafeedback dataset.
+Due to page limits, more detailed results including comparisons with Adam, DPO, full model updates, additional models of Qwen2.5-7B-Instruct, and Mixture2 dataset are deferred to appendix.
+Passive data collection.
+We evaluate the performance of our proposed method in terms of the loss and accuracy of the reward model.
+We compare our OMD-based method with the MLE-based method.
+We randomly sample $T=30,000$ data points from the Ultrafeedback dataset for training.
+Figure 2 shows the loss and accuracy vs.
+the number of training samples.
+We observe that our method converges faster to a lower loss and achieves a higher evaluation accuracy compared to baselines.
+The improvement is particularly pronounced in the small-sample regime ($T<10,000$), where our method achieves a higher evaluation accuracy with the same amount of samples compared to MLE which employs conventional stochastic gradient descent (SGD) updates.
+This shows the superior statistical efficiency of our approach, achieving a better performance with fewer training samples.
+Active data collection.
+In this setting, we only allow the algorithm to select $6,400$ samples out of the whole training datasets for training according to different selection strategies.
+To evaluate the effectiveness of the data selection strategy, we compare our method with the random selection strategy.
+We evaluate the performance of the MLE-based method and our proposed OMD-based method.
+Figure 3 demonstrates that our OMD-based method achieves comparable performance with the MLE-based method for both data collection strategies, while improving the training time by approximately three times.
+Moreover, our data selection strategy outperforms the random selection strategy, showing that our method can effectively select informative data to improve the performance.
+Deployment-time adaptation.
+We divide the dataset into 20 chunks and process them sequentially to simulate the deployment scenario.
+We compare our action selection strategy with (i): random selection, (ii): select the best and second best actions, and (iii): select the best and worst actions.
+We combine the above strategies with MLE-based and OMD-based methods.
+We report both the average cumulative rewards and win rates of each method, where the win rate is defined as the proportion of pairwise comparisons in which a method outperforms all others.
+As shown in Figure 4, our action selection strategy outperforms the baselines for both MLE-based and OMD-based methods.
+This validates the effectiveness of our selection strategy that balances the exploitation of high-reward responses with sufficient exploration to facilitate model improvement.
+Besides, the win rates show that our OMD-based method achieves competitive performance with the MLE-based method.
+(a) Rewards of MLE methods
+(b) Rewards of OMD methods
 (c) Win rates, all methods
-Figure 7: For Llama-3-8B-Instruct on Ultrafeedback-binarized dataset with deployment-time adaptation, we report (a) cumulative reward of MLE-based methods, (b) cumulative reward of OMD-based methods, and (c) win rates between MLE-based and OMD-based estimators with the same action selection strategy.
-(a) Reward of MLE-based methods
-(b) Reward of OMD-based methods
-(c) Win rates, all methods
-Figure 8: For Qwen2.5-7B-Instruct on Ultrafeedback-binarized dataset with deployment-time adaptation, we report (a) cumulative reward of MLE-based methods, (b) cumulative reward of OMD-based methods, and (c) win rates between MLE-based and OMD-based estimators with the same action selection strategy.
+Figure 4: For online RLHF with deployment-time adaptation, we report (a) cumulative rewards of MLE-based methods, (b) cumulative rewards of OMD-based methods, and (c) win rates.
 8 Conclusion
-In this work, we study the problem of online RLHF from the perspective of contextual bandits.
-To address the linear growth in computational complexity with respect to the number of iterations in reward modeling, we propose a novel one-pass reward modeling algorithm that achieves constant-time complexity per iteration.
-Our method is built upon the online mirror descent framework and incorporates a carefully designed local norm that captures second-order information.
-We apply our method to a broad range of online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation, and design tailored algorithms for each scenario.
-We provide theoretical guarantees that our approach achieves superior statistical and computational efficiency relative to existing methods.
-Finally, we provide practical implementations of our method and conduct experiments using the Llama-3-8B-Instruct and Qwen2.5-7B-Instruct models on the Ultrafeedback-binarized and Mixture2 datasets, empirically validating its practical effectiveness.
+In this work, we address a key challenge in online RLHF, where the computational complexity typically grows linearly with the number of iterations.
+To overcome this limitation, we propose a novel one-pass algorithm that eliminates the need to store historical data and achieves constant-time complexity per iteration.
+Our approach is built upon the online mirror descent framework with a carefully designed local norm.
+We apply our method to three online RLHF settings and design tailored algorithms for each scenario.
+We provide both theoretical guarantees and efficient implementations, demonstrating that our approach improves statistical and computational efficiency over existing methods.
+Finally, we validate the effectiveness of our method through extensive experiments.
 While our work advances both the statistical and computational understanding of online RLHF, several important directions remain for future exploration.
@@
 Second, although our analysis is based on the Bradley-Terry model, extending the framework to other preference models, such as the Plackett-Luce model (Luce, 1959; Plackett, 1975), is another promising avenue that may broaden the applicability of our results.
+Acknowledgments
+This research was supported by National Science and Technology Major Project (2022ZD0114800) and NSFC (U23A20382, 62206125). Peng Zhao was supported in part by the Xiaomi Foundation.
 Appendix A Useful Lemmas
 Lemma 2.
-For any $t\in[T]$, define the second-order approximation of the loss function $\ell_{t}(\theta)$ at the estimator $\widetilde{\theta}_{t}$ as ~$\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+\langle\nabla\ell_{t}(\widetilde{\theta}_{t}),\theta-\widetilde{\theta}_{t}\rangle+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}.$ Then, for the following update rule
+For any $t\in[T]$, define the second-order approximation of the loss function $\ell_{t}(\theta)$ at the estimator $\widetilde{\theta}_{t}$ as
+$\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+\langle\nabla\ell_{t}(\widetilde{\theta}_{t}),\theta-\widetilde{\theta}_{t}\rangle+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}.$
+Then, for the following update rule
 $\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\left\{\widetilde{\ell}_{t}(\theta)+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{t}\rVert_{\mathcal{H}_{t}}^{2}\right\},$
@@
 $\lVert\widetilde{\theta}_{t+1}-\theta^{*}\rVert_{\mathcal{H}_{t+1}}^{2}$
-[…]
+$\leq{}$ $2\eta\left(\sum_{i=1}^{t}\ell_{i}(\theta^{*})-\sum_{i=1}^{t}\ell_{i}(\widetilde{\theta}_{i+1})\right)+4\lambda B^{2}+12\sqrt{2}BL^{3}\eta\sum_{i=1}^{t}\lVert\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\rVert_{2}^{2}-\sum_{i=1}^{t}\lVert\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\rVert_{\mathcal{H}_{i}}^{2}.$
 Proof.
 Based on the analysis of (implicit) OMD update (see Lemma 5), for any $i\in[T]$, we have
-~$\big{\langle}\nabla\widetilde{\ell}_{i}(\widetilde{\theta}_{i+1}),\widetilde{\theta}_{i+1}-\theta^{*}\big{\rangle}\leqslant\frac{1}{2\eta}\left(\|\widetilde{\theta}_{i}-\theta^{*}\|_{\mathcal{H}_{i}}^{2}-\|\widetilde{\theta}_{i+1}-\theta^{*}\|_{\mathcal{H}_{i}}^{2}-\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\|_{\mathcal{H}_{i}}^{2}\right)$
+$\big\langle\nabla\widetilde{\ell}_{i}(\widetilde{\theta}_{i+1}),\widetilde{\theta}_{i+1}-\theta^{*}\big\rangle\leqslant\frac{1}{2\eta}\left(\|\widetilde{\theta}_{i}-\theta^{*}\|_{\mathcal{H}_{i}}^{2}-\|\widetilde{\theta}_{i+1}-\theta^{*}\|_{\mathcal{H}_{i}}^{2}-\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\|_{\mathcal{H}_{i}}^{2}\right)$
 According to Lemma 6, we have
@@
 We can further bound the first term of the right-hand side as:
-$\big{\langle}\nabla\ell_{i}(\widetilde{\theta}_{i+1})-\nabla\widetilde{\ell}_{i}(\widetilde{\theta}_{i+1}),\widetilde{\theta}_{i+1}-\theta^{*}\big{\rangle}=$ $\big{\langle}\nabla\ell_{i}(\widetilde{\theta}_{i+1})-\nabla\ell_{i}(\widetilde{\theta}_{i})-\nabla^{2}\ell_{i}(\widetilde{\theta}_{i})(\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}),\widetilde{\theta}_{i+1}-\theta^{*}\big{\rangle}$
-=\ = $\big{\langle}D^{3}\ell_{i}({\xi}_{i+1})[\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}](\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}),\widetilde{\theta}_{i+1}-\theta^{*}\big{\rangle}$
+$\big\langle\nabla\ell_{i}(\widetilde{\theta}_{i+1})-\nabla\widetilde{\ell}_{i}(\widetilde{\theta}_{i+1}),\widetilde{\theta}_{i+1}-\theta^{*}\big\rangle=\$$\big\langle\nabla\ell_{i}(\widetilde{\theta}_{i+1})-\nabla\ell_{i}(\widetilde{\theta}_{i})-\nabla^{2}\ell_{i}(\widetilde{\theta}_{i})(\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}),\widetilde{\theta}_{i+1}-\theta^{*}\big\rangle$
 […]
-[…]
-[…]
+$\leqslant\$$3\sqrt{2}L\big\|\widetilde{\theta}_{i+1}-\theta^{*}\big\|_{2}\big\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\big\|_{\nabla^{2}\ell_{i}({\xi}_{i+1})}^{2}$
+$\leqslant\$$6\sqrt{2}BL\big\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\big\|_{\nabla^{2}\ell_{i}({\xi}_{i+1})}^{2}$
+$\leqslant\$$6\sqrt{2}BL^{3}\big\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\big\|_{2}^{2},$
 where the second equality holds by the mean value theorem, the first inequality holds by the self-concordant-like property of $\ell_{i}(\cdot)$ in Lemma 3, and the last inequality holds by $\widetilde{\theta}_{i+1}$ and $\theta^{*}$ belong to $\Theta=\{\theta\in\mathbb{R}^{d},\lVert\theta\rVert_{2}\leq B\}$, and ${\nabla^{2}\ell_{i}({\xi}_{i+1})}\preceq L^{2}I_{d}$.
@@
 $\lVert\widetilde{\theta}_{t+1}-\theta^{*}\rVert_{\mathcal{H}_{t+1}}^{2}$
-[…]
+$\leq{}$ $2\eta\left(\sum_{i=1}^{t}\ell_{i}(\theta^{*})-\sum_{i=1}^{t}\ell_{i}(\widetilde{\theta}_{i+1})\right)+4\lambda B^{2}+12\sqrt{2}BL^{3}\eta\sum_{i=1}^{t}\lVert\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\rVert_{2}^{2}-\sum_{i=1}^{t}\lVert\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\rVert_{\mathcal{H}_{i}}^{2}.$
 It remains to bound the right-hand side of the above inequality in the following.
 The most challenging part is to bound the term $\sum_{i=1}^{t}\ell_{i}(\theta^{*})-\sum_{i=1}^{t}\ell_{i}(\widetilde{\theta}_{i+1})$.
-This term might seem straightforward to control, as it can be observed that $\theta^{*}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\bar{\ell}(\theta)\triangleq\mathbb{E}_{y_{i}}[\ell_{i}(\theta)]$, where $\ell_{i}(\theta)$ serves as an empirical observation of ¯$\bar{\ell}(\theta)$.
+This term might seem straightforward to control, as it can be observed that $\theta^{*}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\bar{\ell}(\theta)\triangleq\mathbb{E}_{y_{i}}[\ell_{i}(\theta)]$, where $\ell_{i}(\theta)$ serves as an empirical observation of $\bar{\ell}(\theta)$.
 Consequently, the loss gap term seemingly can be bounded using appropriate concentration results.
@@
 To address this issue, following the analysis in Zhang and Sugiyama [2023], we decompose the loss gap into two components by introducing an intermediate term.
-Specifically, we define the softmax function as $[\sigma_{i}(q)]_{1}=\frac{\exp(q)}{1+\exp(q)}$ and $[\sigma_{i}(q)]_{0}=\frac{1}{1+\exp(q)}$, where $[\cdot]_{i}$ denotes the $i$-th element of the vector.
-Then, the loss function $\ell_{i}(\theta)$ can be rewritten as
+Specifically, we define the softmax function as
+$[\sigma_{i}(q)]_{1}=\frac{\exp(q)}{1+\exp(q)}\quad\text{and}\quad[\sigma_{i}(q)]_{0}=\frac{1}{1+\exp(q)},$
+where $\cdot]_{i}$ denotes the $i$-th element of the vector. Then, the loss function $\ell_{i}(\theta)$ can be rewritten as
 $\ell(q_{t},y_{t})=-\mathds{1}_{\{y_{t}=1\}}\cdot\log\left({[\sigma(q_{t})]_{1}}\right)-\mathds{1}_{\{y_{t}=0\}}\cdot\log\left({[\sigma(q_{t})]_{0}}\right).$
-Then, we define the pseudo-inverse function of $\sigma^{-1}(p)$ with $[\sigma^{-1}(p)]_{1}=\log({q}/({1-q}))$ and $[\sigma^{-1}(p)]_{0}=\log((1-p)/{p})$.
+Then, we define the pseudo-inverse function of $\sigma^{-1}(p)$ with
+$[\sigma^{-1}(p)]_{1}=\log({q}/({1-q}))\quad\text{and}\quad[\sigma^{-1}(p)]_{0}=\log((1-p)/{p}).$
 Then, we decompose the regret into two terms by introducing an intermediate term.
@@
 For $\mathtt{term~(a)}$, let $\delta\in(0,1)$ and $\lambda\geq 1$. With probability at least $1-\delta$, for all $t\in[T]$, we have
-$\mathtt{term~{}(a)}\leq\left(3\log(1+2t)+2+LB\right)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right)+2.$
+$\mathtt{term~(a)}\leq 11\cdot(3\log(1+2t)+2+LB)\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+2.$
 For $\mathtt{term~(b)}$, let $\lambda\geq\max\{2,72cd\}$. Then, for all $t\in[T]$, we have
@@
 $\big\|\widetilde{\theta}_{t+1}-\theta^{*}\big\|_{\mathcal{H}_{t+1}}^{2}$ $\leq 12\sqrt{2}BL^{3}\eta\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{2}^{2}+\left(\frac{\eta}{c}-1\right)\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{\mathcal{H}_{i}}^{2}+C.$
-where $C=2\eta\left(3\log(1+2t)+2+BL\right)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right)+4\eta+2\eta\sqrt{6}cd\log\left(1+\frac{2tL^{2}}{d\lambda}\right)+4\lambda B^{2}$.
+where $C=22\eta(3\log(1+2t)+2+LB)\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+4\eta+2\eta\sqrt{6}cd\log\left(1+\frac{2tL^{2}}{d\lambda}\right)+4\lambda B^{2}$.
 Setting $c=7\eta/6$ and $\lambda\geq 84\sqrt{2}BL^{3}\eta$, we have
 $12\sqrt{2}BL^{3}\eta\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{2}^{2}+\left(\frac{\eta}{c}-1\right)\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{\mathcal{H}_{i}}^{2}$
-[…]
-[…]
-Note that $84\sqrt{2}\left(BL^{3}+dL^{2}\right)\eta\geq\max\left\{2L^{2},72cdL^{2},84\sqrt{2}BL^{3}\eta\right\}$, so we set $\lambda\geq 84\sqrt{2}\left(BL^{3}+\right.$ $\left.dL^{2}\right)\eta$.
+$\leq\$$\left(12\sqrt{2}BL^{3}\eta-\frac{\lambda}{7}\right)\sum_{i=1}^{t}\left\|\widetilde{\theta}_{i+1}-\widetilde{\theta}_{i}\right\|_{2}^{2}$
+$\leq\$$0.$
+Note that $84\sqrt{2}\left(BL^{3}+dL^{2}\right)\eta\geq\max\left\{2L^{2},72cdL^{2},84\sqrt{2}BL^{3}\eta\right\}$, so we set $\lambda\geq 84\sqrt{2}\left(BL^{3}+\right$.
+$\left.dL^{2}\right)\eta$.
 As we have $\eta=(1/2)\log 2+\left(BL+1\right)$, we have
-$\big{\|}\widetilde{\theta}_{t+1}-\theta^{*}\big{\|}_{\mathcal{H}_{t+1}}\leq\mathcal{O}\Big{(}\sqrt{d}(\log(t/\delta))^{2}\Big{)}.$
+$\big\|\widetilde{\theta}_{t+1}-\theta^{*}\big\|_{\mathcal{H}_{t+1}}\leq\mathcal{O}\Big(\sqrt{d}\log(t/\delta)\Big).$
 This finishes the proof. ∎
@@
 For the first term, we have with probability at least $1-\delta$, it holds that
-$J\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)=$ $\mathbb{E}_{x\sim\rho}\left[(\theta^{*})^{\top}\phi(s,\pi^{*}(s))\right]-\min_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\theta^{\top}\phi(s,\pi^{*}(s))\right]$
-=\ = $\sup_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\left(\theta^{*}-\widetilde{\theta}_{T}+\widetilde{\theta}_{T}-\theta\right)^{\top}\phi(x,\pi^{*}(x))\right]$
-=\ = $\mathbb{E}_{x\sim\rho}\left[\left(\theta^{*}-\widetilde{\theta}_{T}\right)^{\top}\phi(x,\pi^{*}(x))\right]+\sup_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[\left(\widetilde{\theta}_{T}-\theta\right)^{\top}\phi(x,\pi^{*}(x))\right]$
+$J\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)$
 […]
+[…]
+[…]
+$\leq\$$\Big(\lVert\theta^{*}-\widetilde{\theta}_{T}\rVert_{\mathcal{H}_{T}}+\sup_{\theta\in\mathcal{C}_{T}}\lVert\theta-\widetilde{\theta}_{T}\rVert_{\mathcal{H}_{T}}\Big)\cdot\big\lVert\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big\rVert_{\mathcal{H}_{T}^{-1}},$
 where the first inequality holds by the Cauchy-Schwarz inequality.
@@
 Thus, we obtain
-$J\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)\leq$ $2\widetilde{\beta}_{T}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big{\rVert}_{\mathcal{H}_{T}^{-1}}.$ (13)
+$J\left(\pi^{*}\right)-\widetilde{J}\left(\pi^{*}\right)\leq\$$2\widetilde{\beta}_{T}\cdot\big\lVert\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big\rVert_{\mathcal{H}_{T}^{-1}}.$ (10)
 Combining Eq. (8), Eq. (9), and Eq. (10) and substituting $\widetilde{\beta}_{T}=\mathcal{O}(\sqrt{d}(\log(T/\delta))^{2})$, we have with probability at least $1-\delta$, it holds that
-$\operatorname*{\mathsf{SubOpt}}\left(\pi_{T}\right)\leq 2\widetilde{\beta}_{T}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big{\rVert}_{\mathcal{H}_{T}^{-1}}\leq\mathcal{O}\left(\sqrt{d}\left(\log\frac{T}{\delta}\right)^{2}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big{\rVert}_{\mathcal{H}_{T}^{-1}}\right).$
+$\operatorname*{\mathsf{SubOpt}}\left(\pi_{T}\right)$ $\leq 2\widetilde{\beta}_{T}\cdot\big\lVert\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big\rVert_{\mathcal{H}_{T}^{-1}}$
+$\leq\mathcal{O}\left(\sqrt{d}\left(\log\frac{T}{\delta}\right)^{2}\cdot\big\lVert\mathbb{E}_{x\sim\rho}[\phi(x,\pi^{*}(x))]\big\rVert_{\mathcal{H}_{T}^{-1}}\right).$
 This completes the proof. ∎
@@
 $\big(r(x_{t},\pi^{*}(x_{t}))-r(x_{t},a_{t})\big)+\big(r(x_{t},\pi^{*}(x_{t}))-r(x_{t},a_{t}^{\prime})\big)$
-=\ = $\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\theta^{*}+\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}\theta^{*}$
-=\ = $2\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\theta^{*}+\big{(}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}\theta^{*}$
-=\ = $2\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}(\theta^{*}-\widetilde{\theta}_{t})+2\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\widetilde{\theta}_{t}$
+$=\$$\big(\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big)^{\top}\theta^{*}+\big(\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big)^{\top}\theta^{*}$
+$=\$$2\big(\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big)^{\top}\theta^{*}+\big(\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big)^{\top}\theta^{*}$
+$=\$$2\big(\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big)^{\top}(\theta^{*}-\widetilde{\theta}_{t})+2\big(\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big)^{\top}\widetilde{\theta}_{t}$
 $+\big(\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big)^{\top}(\theta^{*}-\widetilde{\theta}_{t})+\big(\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big)^{\top}\widetilde{\theta}_{t}$
-[…]
+$\leq\$$2\big\lVert\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big\rVert_{\mathcal{H}_{t}^{-1}}\big\lVert\theta^{*}-\widetilde{\theta}_{t}\big\rVert_{\mathcal{H}_{t}}+\big(\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big)^{\top}\widetilde{\theta}_{t}$
 $+\big\lVert\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big\rVert_{\mathcal{H}_{t}^{-1}}\big\lVert\theta^{*}-\widetilde{\theta}_{t}\big\rVert_{\mathcal{H}_{t}}+\big(\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big)^{\top}\widetilde{\theta}_{t}$
-[…]
+$\leq\$$2\widetilde{\beta}_{t}\big\lVert\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big\rVert_{\mathcal{H}_{t}^{-1}}+\big(\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big)^{\top}\widetilde{\theta}_{t}$
 $+\widetilde{\beta}_{t}\big\lVert\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big\rVert_{\mathcal{H}_{t}^{-1}}$
-[…]
+$\leq\$$2\widetilde{\beta}_{t}\big\lVert\phi(x_{t},a_{t}^{\prime})-\phi(x_{t},a_{t})\big\rVert_{\mathcal{H}_{t}^{-1}}+(\phi(x_{t},a_{t}^{\prime})-\phi(x_{t},\pi^{*}(x_{t})))^{\top}\widetilde{\theta}_{t}$
 $+\big(\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big)^{\top}\widetilde{\theta}_{t}+\widetilde{\beta}_{t}\big\lVert\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big\rVert_{\mathcal{H}_{t}^{-1}}$
-=\ = $3\widetilde{\beta}_{t}\big{\lVert}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{\rVert}_{\mathcal{H}_{t}^{-1}},$
+$=\$$3\widetilde{\beta}_{t}\big\lVert\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big\rVert_{\mathcal{H}_{t}^{-1}},$
 where the first inequality holds by the Holder’s inequality and the arm selection strategy of $a_{t}$ such that $\phi(x_{t},\pi^{*}(x_{t}))^{\top}\widetilde{\theta}_{t}\leq\phi(x_{t},a_{t})^{\top}\widetilde{\theta}_{t}$, the second inequality holds by $\widetilde{\theta}_{t}\in\mathcal{C}_{t}$ with probability at least $1-\delta$ by Lemma 1, the third inequality holds by arm selection strategy of $a_{t}^{\prime}$ such that $a_{t}^{\prime}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\phi(x_{t},a)^{\top}\widetilde{\theta}_{t}+2\widetilde{\beta}\lVert\phi(x_{t},a)-\phi(x_{t},a_{t})\rVert_{\mathcal{H}_{t}^{-1}}$.
+Thus, we have
+${\operatorname*{Reg}}_{T}$ $=\frac{1}{2}\sum_{t=1}^{T}\Big(\big(r(x_{t},\pi^{*}(x_{t}))-r(x_{t},a_{t})\big)+\big(r(x_{t},\pi^{*}(x_{t}))-r(x_{t},a_{t}^{\prime})\big)\Big)$
+$\leq\frac{3}{2}\widetilde{\beta}_{T}\sum_{t=1}^{T}\big\lVert\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big\rVert_{\mathcal{H}_{t}^{-1}}.$
 By the definition of $\mathcal{H}_{t}$, we have
@@
 Let $V_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}x_{s}x_{s}^{\top}$ for $\lambda\geq 0$.
-Then, we have $\sum_{s=1}^{t}\left\|z_{s}\right\|_{V_{s}^{-1}}^{2}\leq 2d\log\left(1+\frac{tL^{2}}{\lambda d}\right)$.
+Then, we have
+$\sum_{s=1}^{t}\left\|z_{s}\right\|_{V_{s}^{-1}}^{2}\leq 2d\log\left(1+\frac{tL^{2}}{\lambda d}\right).$
 Lemma 5 (Campolongo and Orabona [2020, Proposition 4.1]).
@@
 $\mathbf{w}_{t+1}=\operatorname*{arg\,min}_{\mathbf{w}\in\mathcal{V}}\big\{\eta\ell_{t}(\mathbf{w})+\mathcal{D}_{\psi}\left(\mathbf{w},\mathbf{w}_{t}\right)\big\},$
-[…]
+where $\mathcal{V}\subseteq\mathcal{W}\subseteq\mathbb{R}^{d}$ is a non-empty convex set.
+Further supposing $\psi(\mathbf{w})$ is 1 -strongly convex w.r.t.
+a certain norm $\|\cdot\|$ in $\mathcal{W}$, then there exists a $\mathbf{g}_{t}^{\prime}\in\partial\ell_{t}\left(\mathbf{w}_{t+1}\right)$ such that
 $\left\langle\eta_{t}\mathbf{g}_{t}^{\prime},\mathbf{w}_{t+1}-\mathbf{u}\right\rangle\leq\left\langle\nabla\psi\left(\mathbf{w}_{t}\right)-\nabla\psi\left(\mathbf{w}_{t+1}\right),\mathbf{w}_{t+1}-\mathbf{u}\right\rangle$
@@
 To further tailor the dataset to our experimental setup, we organize the dataset as follows:
-  • Online RLHF with passive data collection: We randomly sample $T=30,000$ data points from the UltraFeedback-binarized dataset’s train_prefs split for training.
-Each data point consists of a prompt and two responses with a binary preference label indicating the preferred response.
+  • Passive data collection: We randomly choose $30,000$ samples from the UltraFeedback-binarized dataset’s train_prefs split for training.
+Each sample consists of a prompt and two responses with a label indicating the preferred response.
 We use the test_prefs split for evaluation.
-  • Online RLHF with active data collection: We allow the method to actively select 6,400 samples from the train_prefs split according to different selection strategies.
+  • Active data collection: We allow the method to actively select 6,400 samples from the train_prefs split according to different selection strategies.
 The global batch size is set to 8 for training.
@@
 Algorithm 5 Efficient Update using Hessian-Vector Product with Conjugate Gradient
-0:  Current parameter $\widetilde{\theta}_{t}$, gradient $g_{t}(\widetilde{\theta}_{t})$, learning rate $\eta$, max CG steps $K$, base damping $\lambda_{0}$, error tolerance $\epsilon$
+0: Current parameter $\widetilde{\theta}_{t}$, gradient $g_{t}(\widetilde{\theta}_{t})$, learning rate $\eta$, max CG steps $K$, parameter $\lambda_{0}$, $\epsilon$
 1: Initialize $v_{0}=0$, $r_{0}=g_{t}(\widetilde{\theta}_{t})$, $p_{0}=r_{0}$
@@
 3:for $k=0,1,\ldots,K-1$ do
-4:     Compute HVP: ~$\widetilde{\mathcal{H}}_{t}p_{k}=\nabla_{\theta}(\nabla_{\theta}\mathcal{L}(\theta)^{\top}p_{k})|_{\theta=\widetilde{\theta}_{t}}+\lambda_{t}p_{k}$
+4:  Compute HVP: $\widetilde{\mathcal{H}}_{t}p_{k}=\nabla_{\theta}(\nabla_{\theta}\mathcal{L}(\theta)^{\top}p_{k})|_{\theta=\widetilde{\theta}_{t}}+\lambda_{t}p_{k}$
 5:  $\alpha_{k}=\frac{r_{k}^{\top}r_{k}}{p_{k}^{\top}\widetilde{\mathcal{H}}_{t}p_{k}}$, $v_{k+1}=v_{k}+\alpha_{k}p_{k}$, $r_{k+1}=r_{k}-\alpha_{k}\widetilde{\mathcal{H}}_{t}p_{k}$,
@@
 11: Updated parameter $\widetilde{\theta}_{t+1}$
-Practical Implementation.
-To efficiently implement the OMD update in Eq.
-(5) without the costly computing and storing the full Hessian matrix, we utilize the Hessian-vector product (HVP) method combined with conjugate gradient descent.
-The key insight is that the update can be reformulated as solving a linear system.
-For the OMD update:
-$\widetilde{\theta}_{t+1}$ $=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\big{\langle}g_{t}(\widetilde{\theta}_{t}),\theta\big{\rangle}+\frac{1}{2\eta}\big{\|}\theta-\widetilde{\theta}_{t}\big{\|}_{\widetilde{\mathcal{H}}_{t}}^{2}\Big{\}}\approx\widetilde{\theta}_{t}-\eta\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t}),$
-where we ignore the projection operation.
-The key challenge is computing ~$\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t})$ efficiently.
-Instead of explicitly computing ~$\widetilde{\mathcal{H}}_{t}^{-1}$, we solve the linear system ~$\widetilde{\mathcal{H}}_{t}v=g_{t}(\widetilde{\theta}_{t})$
-~$\widetilde{\mathcal{H}}_{t}v=g_{t}(\widetilde{\theta}_{t}),$
-where $v$ is the solution we seek.
-We use the conjugate gradient method with a damped Hessian to solve this system iteratively.
-The HVP operation ~$\widetilde{\mathcal{H}}_{t}v$ for any vector $v$ can be computed efficiently using automatic differentiation ~$\widetilde{\mathcal{H}}_{t}v=\nabla_{\theta}\big{(}\nabla_{\theta}\mathcal{L}(\theta)^{\top}v\big{)}\big{|}_{\theta=\widetilde{\theta}_{t}}+\lambda v$, where $\lambda$ is a damping coefficient that ensures numerical stability and $\mathcal{L}(\theta)$ is the loss function.
+Update details.
+As described in Section 6.1, we can implement the OMD update using the HVP with conjugate gradient descent.
 The full algorithm is summarized in Algorithm 5.
-To ensure stability and convergence, we employ an adaptive damping strategy as $\lambda_{t}=\lambda_{0}\cdot\min\{1,f(t/T)\}$, where $f(\cdot)$ is a damping growth function that can be linear, logarithmic, or cosine.
-In practice, we set $K=3$ and $\lambda_{0}=0.8$ for all experiments.
+In our experiments, we set $K=3$ and $\lambda_{0}=0.8$ and choose the linear function $f(t/T)=t/T$ as the damping function.
 G.2 Validating the Magnitude of $\kappa$
@@
 With the momentum term, our method remains superior to the MLE-based method; however, the performance gap is reduced.
-We think that this is because the Adam optimizer incorporates second-order information for optimization, diminishing the advantage of our method compared to the SGD cases.
+This may be because the Adam optimizer incorporates second-order information for optimization, diminishing the advantage of our method compared to the SGD cases.
 G.4 Comparison with DPO
-We also compare with DPO [Rafailov et al., 2023] in the deployment stage, as a reward-free method, DPO optimizes the policy directly using preference feedback without explicit reward modeling.
+We also compare with DPO [Rafailov et al., 2023] in the deployment stage.
+As a reward-free method, DPO optimizes the policy directly using preference feedback without explicit reward modeling.
 To ensure a fair comparison, we initialize the policy with 400 samples and use the same dataset settings as PPO to iteratively update the policy model using the DPO algorithm.
@@
 This result suggests that DPO’s online learning capability remains a challenge.
-In contrast, the reward model learned by our selection strategy effectively learned streaming data and continuously updates the policy as new data arrive, indicating that in our deployment stage, a reward model with PPO may be a more suitable choice for sequentially learning from new data.
+In contrast, the reward model learned by our selection strategy effectively learned streaming data and continuously updates the policy as new data arrive, indicating that a reward model with PPO may be a more suitable choice for sequentially learning from new data.
 (a) training loss
@@
 G.5 Full Update of Reward Model
-Figure 11 shows deployment-time adaptation results using the meta-llama/Llama-3.2-1B model, where we update all parameters of the reward model instead of only the final layer.
+Figure 7 shows deployment-time adaptation results using the Llama-3.2-1B model, where we update all parameters of the reward model instead of only the final layer.
 Both our method and MLE use the same action selection strategy.
-Our approach achieves comparable performance with MLE, indicating that our OMD-based method is still compatible with full-model updates.
+Our approach achieves comparable performance with MLE, indicating that our OMD-based update method is still compatible with full-model updates.
+G.6 More Foundation Models and Datasets
+In this section, we provide more experimental results about other foundation models and datasets.
+Figure 8 shows the training and evaluation curves for reward model learning under passive data collection using the Qwen2.5-7B-Instruct model.
+We compare our method with MLE and report the loss and accuracy over training.
+Our method consistently shows stable training dynamics and competitive evaluation performance compared to MLE, suggesting its effectiveness in offline settings.
+Figure 10 present results for online RLHF with active data collection using the same Qwen model.
+Figure 9(a) shows training loss curves, while Figure 9(b) reports evaluation accuracy over training iterations.
+Table 9(c) further compares various methods (Rand-MLE, Active-MLE, Rand-OMD, and our approach) in terms of final accuracy and training time.
+While Active-MLE achieves slightly higher accuracy, our method provides significant speedup in training time with comparable performance, highlighting the efficiency of our approach.
+Figure 10 illustrates the deployment-time performance of various methods on the Ultrafeedback dataset.
+We split the dataset into 20 chunks and measure cumulative rewards across these chunks.
+Our method demonstrates robust adaptation capabilities, achieving competitive reward accumulation.
+(a) training loss
+(b) training accuracy
+(c) evaluation loss
+(d) evaluation accuracy
+Figure 8: For Qwen2.5-7B-Instruct model with passive data collection, we compare our method with MLE. We report average accuracy and loss curve of the reward model.
+(a) training loss
+(b) evaluation accuracy
+Method ACC (%) Time (s)
+Rand-MLE 67.36 $\pm$ 0.5 4653 $\pm$ 121
+Active-MLE 67.25 $\pm$ 0.4 4701 $\pm$ 103
+Rand-OMD 66.80 $\pm$ 0.5 1312 $\pm$ 47 0
+Ours 67.11 $\pm$ 0.5 1325 $\pm$ 54 0
+(c) training time
+Figure 9: For Qwen2.5-7B-Instruct with active data collection, we report the comparison of different methods about (a) training loss, (b) evaluation accuracy and (c) evaluation accuracy and training time.
+Figure 10: Results of deployment-time adaptation for Qwen2.5-7B-Instruct.
+(a) training loss
+(b) training accuracy
+(c) evaluation loss
+(d) evaluation accuracy
+Figure 11: For online RLHF with passive data collection on the Llama-3-8B-Instruct model on the Mixture2 dataset, we compare our method with MLE.
+We report average accuracy and loss curve of the reward model.
+Finally, Figure 11 shows results on the Llama-3-8B-Instruct model trained on the Mixture2 dataset in a passive data collection setup.
+Similar to earlier observations, our method achieves competitive or superior performance compared to MLE, both in terms of training and evaluation loss/accuracy, demonstrating its generality across different model and dataset combinations.
+Appendix H Broader Impact
+Our work advances the efficiency of RLHF, a central technique in aligning large language models with human values and preferences.
+By proposing a new one-pass reward modeling method that eliminates the need to store historical data and re-train from scratch, we reduce the computational and environmental costs commonly associated with online RLHF pipelines.
+This could enable the development and deployment of aligned language models by institutions with limited resources.
+However, the broader deployment of RLHF, particularly in an online and adaptive setting, raises important ethical and societal considerations.
+On the positive side, it can enable more responsive and value-aligned AI systems, with potential applications in education, healthcare, and accessibility.
+Yet, the ability to iteratively adapt to user feedback in deployment may also increase the risk of reinforcing harmful biases or being gamed by adversarial users, especially in high-stakes or open-ended domains.